A Complete Bibliography of ACM Transactions on Database Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 September 2024 Version 2.95

Title word cross-reference

(β, γ, δ) [250]. ^{+} [602]. ϵ [1013]. F_{0} [1069]. K [633, 986, 586, 800, 548, 843, 896, 760, 874, 678, 580, 648, 898, 702]. N [226]. pq [749]. R [1025].

-constraints [586]. -gram [749]. -Nearest [874]. -nn [633]. -Regret [986]. -Relational [226]. -tree [602]. -trees [1025].

1 [94, 93, 140, 112, 129, 92]. 1.1 [968]. 13 [403]. 16 [388]. 1975 [1088]. 1976&ndash1978 [110].

2007 [713, 710, 709, 716]. 2008 [739].

’86 [1092].

9 [244]. 92a [388]. 9th [1090].

= [344].

Abstract [353, 274]. Abstraction [9, 550]. Abstractions [67, 28]. Accelerating [574, 1027]. Acceleration [1040, 582]. Access [4, 902, 894, 40, 1055, 1063, 80, 57, 321, 386, 308, 187, 25, 500, 757, 664, 538, 498]. accessed [681]. Accesses [239, 1]. accessibility [581]. accessible [580]. Account [125]. accuracy [729]. Accurate [1085, 959, 765]. Achieving [191, 321]. ACM [1091, 1094, 403, 388, 253, 679, 244, 2, 710]. acquisitional [596]. across [1075, 630, 873]. ACTA [446]. Action [946]. actions [553]. Active [452, 464, 462, 743, 527]. Activity [468]. Activity-Driven [468]. Ad [117, 884, 957, 1076]. Ad-Hoc [884]. Adaptable [207]. Adaptive [636, 992, 1017, 338, 652, 558, 247, 490, 685, 549, 602, 484]. Addendum [244]. Additive [1028]. Addressable [151]. Addressing [1086]. Adjustment [291]. Admission [248]. Advanced [595]. adversaries [753]. Against [69]. Aggregate [979, 390, 420, 1028, 366, 305, 606, 1080, 799, 844, 702]. Aggregates [939, 715, 701]. Aggregation [1062, 968, 996, 28, 1031, 637, 726, 678]. AGM [332]. Agreement [262]. Ahead [397]. aim [2]. Algebra [473, 390, 981, 366, 305, 396, 322, 1023, 403, 650, 677, 534]. Algebraic [847, 527]. Algorithm [230, 167, 391, 153, 368, 24, 356, 10, 927, 557, 484, 753]. algorithmic [814]. Algorithms [363, 215, 195, 190, 11, 1003, 320, 1017, 1014, 1084, 457, 888, 1000, 428, 357, 415, 1041, 18, 953, 874, 1034, 234, 445, 1087, 386, 694, 312, 460, 827, 963, 42, 211, 1010, 71, 786, 770, 604, 631, 522, 558, 751, 813]. Alignment [829]. Allocation [417, 318, 124, 149, 6]. allow [642]. Almost [1078, 848]. Almost-linear [848]. Alternating [458]. Alternative [520, 830, 564]. Alternatives [307, 485]. Altruistic [438]. Amalgamating [442]. Ambiguity [461, 354]. Among [441]. Analyses [747]. Analysis [452, 974, 984, 902, 231, 244, 290, 136, 261, 883, 348, 143, 170, 418, 99, 38, 1087, 228, 386, 841, 210, 302, 565, 241, 25, 743, 527, 501, 515, 718, 535, 792, 638, 750, 858, 780, 719, 654, 517, 840, 618, 784]. Analytics [995, 828, 953]. annotation [862]. Annotations [382]. Anonymization [728, 729, 706, 776, 753]. Anonymization-based [728]. anonymized [812]. Answer [918, 885]. Answering [983, 1055, 960, 540, 178, 799, 699, 723, 819, 810].

Answers [1063, 419, 410, 923, 993]. anticipated [628]. Antijoin [913]. any [574]. Apologizing [353]. Application [461, 193, 198, 15, 125, 1076, 33, 793, 583]. Application-Level [1076]. Applications [284, 262, 100, 946, 915, 1076, 356, 501, 478, 5, 711]. Applying [483, 533]. Approach [412, 8, 139, 266, 355, 1053, 20, 934, 462, 1065, 468, 198, 431, 405, 1012, 993, 359, 74, 192, 315, 165, 147, 578, 527, 687, 649, 582, 839, 630, 664, 528, 787, 632, 682, 784, 490]. Approximate [698, 912, 1027, 956, 1010, 908, 668, 554, 726, 551, 626, 712, 671, 813, 688, 723]. Approximating [1]. Approximation [1053, 965, 631, 1033, 963]. Arbitrary [875, 1029, 167, 637]. Architectural [38]. Architecture [133, 155, 939, 769, 796]. architecture-sensitive [796]. Architectures [258, 189, 261]. Archiving [572]. area [636]. ARIES [475, 397]. arithmetic [823]. array [685]. Articles [887]. Artifact [823, 1006]. Asking [353]. Aspect [63]. Aspects [447]. Assignment [968, 402]. assisted [732]. association [537]. Associative [189, 132, 50, 30, 286, 5]. Assuming [137, 203]. Assumption [160, 219]. Assumptions [212]. Asymmetric [849]. Asymptotically [786]. Asynchronous [1017]. Atlantic [1094]. Atomic [938, 942]. Atomicity [196, 546]. ATSQL [536]. attacks [728]. Attention [991]. Attribute [95, 804, 25]. Attribute-Based [95]. Attributes [305, 804]. Authentication [891]. Author [433, 491]. Authority [690]. Authority-based [690]. Authorization [59, 14, 373, 562]. Autocompletion [927]. automata [591]. Automated [440, 1084]. Automatic [444, 630, 411, 339, 155, 409, 752, 7]. autonomous [634]. Autonomously [340]. Availability [337, 416, 236]. available [612]. Average [222]. Aware [946, 763, 652, 725, 706]. axiomatic [480, 814]. Axiomatization [383].

B [23, 942, 761, 818, 802, 602, 624, 415, 142, 121]. B-Tree [415, 761, 802, 624]. B-Trees [142, 121, 942, 23, 818]. Back [261, 742]. Back-End [261]. Backup [380, 32, 223, 445]. Bag [798, 1042]. bags [557]. Balanced [89, 250]. Balancing [1045]. Bancilhon [303]. Base [414, 520]. Based [370, 151, 426, 398, 400, 894, 40, 950, 365, 355, 472, 901, 131, 95, 177, 1067, 1085, 428, 70, 219, 308, 1083, 122, 315, 25, 967, 1030, 691, 483, 720, 837, 636, 589, 1003, 582, 1073, 690, 664, 602, 758, 475, 842, 791, 670, 856, 1048, 728, 705, 618, 657, 806, 784, 490]. Bases [1092, 1088, 1089, 9, 1090, 442, 539, 1089]. Batch [277, 32, 908]. Batched [348, 242, 251, 16]. Bayesian [976]. BCNF [389]. be [474]. Behavior [452, 213, 468, 429, 535]. Behind [946]. Being [930]. Belief [993]. Benchmark [394]. Benchmarking [953]. benchmarks [535]. Benefit [302]. Benefits [76]. Best [904]. Better [1079, 821]. Between [77, 749, 623, 809, 944, 754]. BEVA [927]. Beyond [398, 833, 459, 945, 821]. Bi [1068]. Bi-objective [1068]. Big [1073]. Binary [108]. bindings [540]. bitmap [675, 622, 747]. Blazes [974]. blind [660]. Block [935, 1]. Blocking [37]. Blocks [468]. Bloom [1001, 992]. Board [897, 922, 952, 977, 1011]. BonXai [966]. Boolean [866, 841, 443]. Boosting [912]. Borealis [692]. both [734]. Bound [1058]. Bounded [983, 941, 982, 456, 449, 1041, 182, 326, 843, 1039]. Bounds [1084, 899, 771, 866, 601, 705]. Bouquets [934]. bridge [711]. Bringing [742]. Broadcast [506]. browsing [511]. Bucket [222]. bucketing [592]. Buffer [229, 41, 341, 281]. Buffers [31]. Building [468, 980, 944]. Bulk [1025, 578]. Bulk-loading [1025]. business [488]. Byzantine [262].

Cache [942, 227, 697, 532, 485, 643]. Cache-Oblivious [942, 697]. Caching [361, 410]. Cactis [338]. CAD [254]. Calculating [104]. Calculus [62, 366, 305, 322, 403, 377, 352, 531]. callback [490]. callback-based [490]. CALM [918]. CALM-Conjecture [918]. candidate [601]. Capabilities [61]. Capacity [360, 754]. Capture [84, 1051]. Capturing [799, 933, 614, 865, 695]. Cardinal [809]. Cardinality [328, 1085, 387, 913, 807]. Carlo [792]. Cartesian [149]. CASDAL [47]. Case [432, 945, 827, 1078, 750, 542, 1036, 420]. CASE-DB [420]. Casper* [744]. CAS}SM [47]. Catching [1020]. categorical [870]. Category [1084]. Causes [1077]. Cautious [407, 248]. Centrality [1083]. Centralized [257]. centric [680]. Certain [212, 923, 810]. Chaining [143]. change [834]. Changes [988, 125, 673, 523]. Characteristics [177]. Characterization [187]. Characterizations [1060, 814]. Characterizing [797, 576, 466]. Charter [434, 451]. Chase [335]. Checking [456, 955, 512, 502]. Checkpointing [32]. Chopping [460]. ChronicleDB [1007]. circuits [787]. Citations [931]. Cite [931]. Citizens [474]. City [1094]. Class [85, 975, 474, 174, 522]. classes [526]. Classification [862, 588]. Classifier [987]. clause [371]. Cleaning [928, 947, 1065, 1040, 638]. CLIDE [681]. client [485, 490]. client-server [485, 490]. clique [672]. cliques [795]. close [792]. Closed [972, 788, 767, 672]. closest [765]. Closing [905]. Closure [363, 428, 367, 610]. Cloud [1068, 838]. Cluster [238]. Clustered [232, 167, 60]. Clustering [365, 896, 368, 959, 24, 871, 247, 659, 755].

CNN [1027]. Coalesced [231, 244]. CODASYL [213, 146]. Codd [202]. Coding [76, 1064, 308]. Coefficient [365]. Coherency [423]. coherent [672]. Collaborative [800, 852]. Color [358]. column [769, 835]. column-store [769]. Combine [885]. Combining [56, 966]. Comment [303]. Comments [91, 760, 129, 251]. commerce [504]. Commercial [258]. Commission [422]. Commit [422, 301, 625]. common [488]. Communication [468]. Commutativity [398, 718]. Compaction [215]. Comparing [811]. Comparison [158, 302, 141]. Comparisons [896]. Compile [362]. Compile-Time [362]. Compiling [711]. complements [561]. Complete [383]. Completeness [437, 344, 772, 1037]. Complex [1046, 49, 1085, 468, 630, 651, 854, 623, 809, 508]. Complexity [936, 1035, 937, 876, 264, 893, 372, 199, 988, 714, 621, 651, 857, 640, 736]. component [747]. Components [39, 1082]. Composing [619]. Composite [336, 758]. Composition [663, 793]. compound [523]. Comprehensive [1003, 949, 962]. COMPRESS [962]. Compressed [1054, 747, 581, 816]. Compression [1003, 962, 367, 1041, 844, 821, 622]. Compromise [306]. Compromising [179, 744]. Computable [340]. Computation [455, 268, 900, 561, 594, 812]. Computation-Tuple [268]. Computational [66, 955]. Computations [994, 1052, 1027, 1013]. Computer [61, 175]. Computers [210]. Computing [875, 226, 1015, 760, 648, 701]. Concatenation [975]. Concept [386]. Concepts [61, 254, 317, 365, 295]. Conceptual [249, 404, 501, 909, 147, 839, 647, 517]. Concise [598, 756]. Concurrency [259, 307, 398, 94, 93, 195, 230, 358, 316, 292, 345, 237, 401, 353, 447, 449, 416, 123, 196, 224, 129, 209, 423, 52, 74, 624]. Concurrent [338, 108, 142, 312]. condensation [691]. condensation-based [691]. Conditional [695, 617]. Conditions [406, 364, 467, 575]. Conference [1094, 1092, 1088, 1089, 1090, 679, 253]. Confidence [705]. Confidential [306]. configuration [752]. Conflict [393, 218]. Conflict-free [218]. Conflicts [181]. Conjecture [918]. Conjunctive [1035, 1055, 1063, 459, 825, 960, 1051, 1060, 862, 810]. Connected [78, 1082, 813]. Consecutive [232]. Consensus [448, 625, 74]. Consequences [137, 203]. Consistency [426, 277, 225, 357, 392, 159, 485, 499]. Consistent [325, 960, 187, 855, 872]. Constant [1032, 389, 399]. Constant-Delay [1032]. Constant-Time [399]. Constant-Time-Maintainable [389]. Constrained [420, 1038, 566]. Constraint [157, 665, 493]. Constraints [1005, 456, 947, 300, 104, 960, 1040, 586, 500, 738, 729, 614, 842, 731, 489, 754]. construct [517]. Construction [1084, 101, 770, 631, 708]. Constructs [35]. Containing [431]. Containment [459, 1042, 862, 558]. Content [151, 582]. Contention [401]. Context [1066]. Continuous [762, 783, 1036, 799, 733, 576, 586, 698, 835, 724, 782, 694]. continuously [633]. Control [259, 307, 398, 94, 93, 195, 230, 40, 358, 316, 401, 353, 132, 248, 447, 416, 123, 196, 224, 129, 392, 209, 423, 52, 74, 500, 538, 624]. Conversion [155]. Converting [396]. Coordination [974, 97, 822]. Coping [785]. coprocessing [741]. Copy [380, 74]. Core [1047, 597, 796, 672]. Correct [162, 187, 780]. Correctness [426, 94, 201, 196, 502]. Correlated [703]. correlation [630]. Correlations [875]. Correspondence [911, 204, 203]. Corresponds [144]. Corrigenda [496, 109].

Corrigendum [188, 388]. Cost [26, 1003, 1073, 416, 874, 386, 246, 302, 552, 909, 25, 524, 613, 664, 504]. Cost-based [1003, 1073, 664]. Cost-Benefit [302]. Cost-Effective [909]. Costs [267]. Counting [951, 975, 888, 1069, 356]. Counts [48]. Coupled [62]. Courts [208]. Cover [365]. Cover-Coefficient-Based [365]. Coverage [901]. Covers [298]. crawlers [567]. Creating [319]. Criteria [426, 21]. Criterion [196]. Crowdsourcing [1002]. CSP [894]. Cubes [885]. Currency [826]. Current [415]. Curves [1025, 871]. Cyclotron [801].

d [144]. DAPLEX [120]. Dark [931]. Data [991, 412, 361, 318, 889, 282, 903, 997, 832, 317, 894, 186, 62, 393, 1079, 3, 461, 1053, 43, 53, 316, 597, 1066, 826, 537, 1094, 330, 49, 291, 353, 1073, 114, 87, 1067, 1092, 1088, 1018, 995, 51, 416, 424, 81, 50, 878, 1071, 895, 256, 953, 1089, 67, 86, 885, 37, 418, 1012, 907, 9, 925, 274, 884, 929, 990, 206, 1083, 1090, 35, 957, 120, 1040, 520, 125, 302, 963, 944, 395, 1037, 388, 375, 976, 799, 691, 797, 576, 742, 786, 586, 524, 636, 572, 650, 779, 589, 726, 525, 823, 820, 757, 785, 695, 844]. data [555, 658, 746, 729, 626, 822, 814, 788, 763, 454, 835, 792, 715, 564, 638, 852, 532, 724, 475, 782, 689, 476, 612, 760, 805, 21, 648, 593, 854, 516, 870, 541, 646, 632, 629, 855, 644, 682, 29, 675, 732, 545, 674, 775, 484, 728, 735, 609, 47, 801]. Data-completeness [1037]. data-driven [822]. data-partitioning [632]. Data-Swapping [206]. Database [1091, 304, 328, 452, 283, 318, 8, 65, 61, 139, 258, 98, 400, 266, 91, 243, 332, 190, 124, 62, 31, 324, 208, 152, 465, 189, 1081, 376, 244, 916, 930, 119, 212, 84, 144, 150, 127, 225, 68, 96, 320, 20, 191, 213, 229, 316, 227, 288, 313, 154, 184, 262, 271, 474, 14, 471, 57, 261, 130, 389, 172, 2, 1067, 338, 357, 410, 205, 219, 41, 107, 507, 67, 27, 273, 196, 235, 342, 267, 216, 276, 349, 387, 100, 99, 234, 274, 38, 346, 239, 430, 373, 423, 314, 34, 915, 52, 439]. Database [238, 281, 246, 55, 382, 263, 270, 339, 19, 63, 155, 210, 236, 28, 54, 7, 102, 183, 297, 1093, 319, 488, 403, 436, 872, 159, 233, 192, 211, 1037, 356, 352, 611, 388, 165, 10, 1, 25, 115, 145, 147, 166, 1095, 1030, 134, 527, 106, 556, 649, 838, 515, 502, 535, 624, 792, 528, 750, 487, 850, 5, 594, 719, 845, 634, 477, 579, 751, 752, 285, 793, 647]. Database-Intensive [100]. Databases [169, 426, 1078, 413, 171, 168, 103, 470, 983, 94, 93, 140, 230, 82, 365, 201, 453, 458, 1068, 48, 185, 358, 406, 109, 105, 69, 421, 337, 36, 131, 926, 313, 462, 323, 892, 886, 457, 180, 112, 22, 362, 351, 116, 128, 78, 350, 101, 176, 235, 431, 133, 223, 405, 331, 97, 392, 435, 306, 469, 322, 92, 156, 296, 187, 118, 72, 15, 378, 88, 463, 125, 255, 403, 257, 42, 74, 188, 252, 315, 450, 301, 364, 179, 79, 578, 548, 740, 549, 762, 722, 536, 482, 506, 534, 767, 699, 759, 777]. databases [697, 511, 690, 652, 673, 835, 703, 708, 748, 731, 547, 580, 645, 711, 587, 606, 791, 661, 493, 113, 734, 812, 584, 481, 672]. Dataflow [998, 332]. datalog [494, 973, 894, 486, 943]. datasets [706, 635, 754]. DB [420]. DBMS [954, 1077, 949, 613]. DBO [712]. DBSCAN [965, 970]. DBTG [42]. Deadlock [173, 153]. decaying [1001]. Decentralized [939]. Deciding [906, 1059]. Decision [20, 359, 302]. Decision-Theoretic [359]. Declarative [918, 973, 940, 453, 928, 578, 822]. Declustering [472]. Decompiling [146]. Decomposition [314, 13, 483, 787]. Decompositions [111, 1074]. Decoupling [592]. Deduction [411]. Deductive [470, 458, 362, 392, 430, 469, 463, 547, 285]. Deep [1027]. Deferred [241]. Definability [936]. Definable [988]. Define [126]. defined [509, 554, 613, 1080]. Defining [43, 1013]. Definition [462, 147]. definitions [814, 864]. Degradation [172].

Degree [983]. Delay [1032]. Deletion [825, 1070, 869]. delivery [636]. Deltas [474]. Dense [135, 1082, 672]. Densest [1043]. dependence [529]. Dependencies [111, 103, 17, 950, 1081, 53, 36, 999, 1026, 264, 1018, 1019, 167, 1015, 948, 86, 314, 1024, 958, 969, 911, 395, 1037, 550, 823, 619, 695, 814, 161, 371, 663, 790, 583]. Dependency [218, 489]. Depth [978, 1074, 329]. Derivation [409]. Derived [340, 461]. Description [556, 130]. descriptions [598]. Descriptors [200]. Design [169, 259, 363, 464, 66, 266, 1045, 404, 376, 119, 126, 313, 181, 133, 267, 519, 234, 274, 439, 12, 319, 909, 1037, 89, 784, 115, 147, 166, 687, 515, 750, 507, 5, 721, 488, 647, 481]. Designed [984]. Designing [972, 205, 796, 199, 100]. Designs [148, 839]. Desired [46]. Detecting [727, 340, 969]. Detection [393, 1071, 153, 961, 834, 789]. Determinacy [766]. Determination [37, 77]. Determining [826, 161]. Determinism [906]. deterministic [591]. Developing [49, 528]. Development [150, 198]. Dichotomies [926, 1009]. Different [942]. Different-Sized [942]. Differential [169, 15, 790, 958, 911, 908]. Differentially [1068, 1013]. Differentiating [805]. Difficulty [64]. digital [504]. dimensional [524, 537, 765]. dimensionality [549]. Dimensions [1050]. Direct [363, 1063]. Directed [1043, 781]. directions [809]. Disambiguation [1071]. Disaster [380]. DisC [901]. Discovering [889, 1026, 560]. Discovery [997, 1043, 1024, 963, 911, 510, 790]. discrete [835, 632]. discrete-valued [835]. discriminative [618]. Disjunctive [894, 950, 486]. Disk [417, 290, 149, 197, 342]. Disks [245]. Disorder [326]. Dissemination [443, 518]. Dissimilarity [901, 688]. Distance [511, 900, 749, 600, 842, 610]. distances [554]. Distortion [256]. Distributed [974, 318, 94, 93, 140, 230, 201, 833, 225, 421, 191, 154, 184, 271, 877, 330, 112, 272, 985, 878, 6, 97, 276, 153, 990, 52, 92, 378, 545, 872, 159, 301, 692, 589, 698, 532, 476, 612, 634, 682, 812]. Distribution [256]. Distributive [391]. Diversification [876, 901]. Diversity [901, 843]. divide [659]. divide-and-merge [659]. DML [146]. Do [931]. Doable [1052]. Document [1032, 1014, 837]. documents [577, 590, 777, 860]. does [494]. Domain [638, 352, 647]. Domain-independent [638]. DomainNet [1071]. Domains [131, 274, 541]. Dominance [1029]. Domination [875]. double [660]. double-blind [660]. Doubly [364]. Driven [468, 913, 822, 570]. DTD [966]. DTDs [756]. Duplicate [186, 1062, 789]. Duration [447]. Dynamic [379, 144, 80, 291, 357, 321, 224, 331, 632, 122, 988, 520, 1082, 241, 310, 10, 691, 522, 542, 566, 480, 559]. dynamically [599, 582].

eager [528]. EAS [198]. EAS-E [198].

EDBT [686]. Edge [173]. Edit

[900, 600, 842, 849]. edit-distance [842]. Editor [283]. Editorial [879, 904, 897, 931, 922, 952, 977, 1011, 1021, 660, 977, 1011]. Effect [347]. Effective [567, 909, 333, 531]. effectively [778]. Effectiveness [365]. Effects [34]. Efficiency [1025, 134]. Efficient [412, 85, 335, 984, 720, 1047, 910, 1038, 1068, 722, 1053, 456, 820, 1062, 1064, 1014, 604, 1058, 730, 888, 566, 142, 708, 874, 678, 37, 133, 223, 907, 870, 900, 661, 898, 265, 1024, 980, 827, 765, 944, 1031, 245, 1010, 789, 967, 927, 786, 585, 506, 502, 729, 770, 781, 760, 648, 849, 622]. Efficiently [82, 1065]. elastic [838]. ElasTraS [838]. electronic [504]. elements [557, 760, 648]. Elevating [474]. ELFS [280]. Elimination [186, 935, 469]. Emancipating [526]. Embedded [998, 343, 1037, 835, 663]. Embedding [791, 693]. Embedding-based [791]. embeddings [667, 600]. Empirical [1033, 1025, 423, 1076]. EmptyHeaded [971]. Enabling [1070, 822]. Encipherment [4]. Enclaved [1068]. Encoding [171, 456]. encountered [523]. Encryption [127, 757, 1067]. End [261]. ENFrame [925]. Engine [971, 712, 607]. Engineering [987]. Engines [1070, 980]. English [280]. Enhancements [183]. Enhancing [38]. Ensuring [499]. Entangled [822]. Enterprise [944]. Entities [940, 1075]. Entity [3, 381, 405, 638]. Entity-Relationship [3, 381, 405, 638]. entropy [807]. Enumeration [1032, 1035, 1055, 1014]. Environment [288, 525, 689, 647]. Environments [401, 559]. ephemeral [478]. Equi [884]. Equi-Join [884]. Equijoin [378]. equivalence [798]. Error [1001, 432, 1041, 927, 771, 616, 707]. Error-Tolerant [927]. Escrow [278]. Essence [915]. Essential [588]. estimates [705]. Estimating [328, 979, 715, 671, 387, 246]. Estimation [1085, 1069, 913, 46, 661, 807, 704]. estimations [670]. Estimators [390, 1033]. Euclidean [965]. Evaluating [580, 1009, 751]. Evaluation [408, 259, 363, 984, 458, 1038, 212, 833, 288, 892, 95, 407, 1041, 1034, 30, 445, 423, 1039, 71, 720, 774, 548, 585, 569, 574, 850, 21, 545]. Even [1082]. Event [1046, 1065, 1007, 854]. evolution [523, 480]. Evolutions [932]. Exact [936, 951, 813, 956, 1060, 908, 849, 688]. E}xample [603, 937, 346]. Examples [919, 963, 797]. exceptional [727]. Exceptions [260, 514]. exchange [597, 785, 658, 788, 852]. exchanged [742]. Exchanging [593]. Execution [190, 201, 461, 939, 806]. Exemplar [1004]. Expansions [170, 287, 418]. expensive [497]. Experimental [1003, 1041, 329]. Experiments [77]. Explaining [354]. Exploiting [586, 947, 808]. Exponential [182]. export [779]. Exporting [681]. EXPRESS [29]. Expressing [579]. Expression [920, 848, 841]. Expressions [85, 335, 455, 951, 975, 104, 242, 396, 756, 857]. Expressive [989, 832, 1008, 650]. Expressiveness [1045, 714, 640, 966, 717]. Extended [62, 446, 144, 669, 381, 478, 425, 405, 322, 403]. Extendible [80]. Extending [84, 949, 305, 297, 868]. Extensible [317, 400, 850]. Extension [31, 218, 544, 541]. External [882, 126, 1044, 860]. Extra [1078]. Extracting [585]. Extraction [928, 299, 909, 725, 732, 29].

F [403]. Facilitate [441]. Facility

[43, 126, 100, 7]. Fact [194, 955]. Factor [920]. Factorised [899]. Factors [158, 37, 63, 141]. Failures [422]. Fairest [1050]. Fairness [906]. Family [1083]. Fast [991, 96, 320, 227, 80, 551, 1074, 457, 995, 959, 872, 796, 629, 688]. Faster [921]. Fault [692]. Fault-tolerance [692]. FDs [508]. Feature [657, 924]. Feature-based [657]. Features [260, 38, 134]. Federal [208]. Fields [73, 835]. Fifth [235]. File [151, 148, 144, 288, 193, 51, 107, 1086, 287, 326, 418, 207, 336, 1087, 122, 89, 18]. File-Organization [89]. Files [169, 26, 90, 4, 186, 472, 232, 149, 80, 348, 143, 37, 242, 251, 60, 15, 16, 45, 33, 505]. Filling [871, 1025]. filter [856]. Filtering [362, 569]. Filters [1001, 992, 1016]. Find [1082]. Finding [1056, 795, 96, 881, 830, 557]. Fine [918, 1079, 397, 490]. Fine-Grained [918, 1079, 490]. Fine-Granularity [397]. Finite [419, 341]. First [936, 474, 206, 296, 110, 610, 810]. First-Class [474]. First-Order [936, 610, 810]. First-Order-Logic [296]. Five [358]. Five-Color [358]. Fixpoint [458]. Flat [396]. Flexible [260, 1029, 538, 850, 487, 562]. Florence [1090]. Flow [316]. Fly [921]. FO [983]. Forensic [719]. forensics [845]. Foreword [656, 887, 571, 709, 768, 794, 817, 824, 853, 739, 615]. Form [171, 66, 17, 36, 131, 128, 466, 289, 166, 577]. Formal [304, 185, 1046, 447, 385, 354, 450, 147, 483]. Formatted [26]. Forms [918, 406, 334, 489, 583]. formulas [568]. Formulation [840]. Forward [284, 662, 846]. Foster [818]. Foundation [202, 521, 547]. Foundations [217]. Fourth [1091]. FPGA [856]. FPGA-based [856]. fragment [730]. Fragmentation [265]. Fragmented [372]. Fragrant [934]. Frame [133]. Framework [1005, 940, 1046, 962, 987, 964, 925, 1048, 310, 333, 1013, 642, 555, 729, 864, 688, 562]. Framingham [1088]. Free [197, 960, 218]. Freedom [173]. Frequent [1005, 893, 599, 557, 566, 760, 648, 618]. Full [136]. Function [44]. Functional [111, 103, 17, 1081, 1026, 1028, 1018, 1015, 948, 371, 314, 120, 395, 1037, 695, 651, 787, 583]. Functions [370, 1029, 180, 366, 305, 986, 1080, 637, 866, 613, 682]. Further [331]. Future [1034, 628]. Fuzzy [314, 868].

Galileo [249, 283]. Game [1012]. Game-theoretic [1012]. Gap [905]. gathering [503]. General [358, 1031, 616]. Generalization [28, 399]. Generalized [412, 43, 57, 395, 609]. Generalizing [845]. generated [582]. Generating [950, 1019, 780, 764]. Generation [444, 402, 235, 373, 60, 7, 670]. generic [689]. geographical [855]. Geometric [400, 877, 945, 682]. getting [597]. Given [46, 663]. Glass [1076]. Global [181, 553]. Globally [187]. GlOSS [510]. Goal [946]. Goal-Aware [946]. Good [326, 369]. Grained [918, 1079, 490]. gram [749]. Grammar [40]. Grammar-Based [40]. granularities [481]. Granularity [397, 34, 75]. Graph [971, 1078, 832, 1038, 994, 1017, 1026, 1052, 1085, 1065, 428, 883, 917, 239, 1083, 618, 722, 851, 839, 638, 865, 657, 672]. Graph-Structured [832]. Graph-Theoretic [239, 839]. Graphical [283]. graphics [741]. Graphs [937, 316, 999, 1020, 1075, 935, 964, 1082, 781, 646]. GraphZeppelin [1082]. Grid [207, 1087]. Group [910, 705, 735]. group-by [735]. Grouping [1062, 592]. Groups [1051]. Growing [45]. Guaranteed [369]. Guaranteeing [406]. guarantees [707]. Guarded [950]. Guarded-Based [950]. Guide [169].

H [403]. habitat [836]. Handling [260, 240]. Hardness [965, 1084]. Hardware [65, 177, 132]. Hardware-Based [177]. Hash [76, 379, 378, 676]. Hash-Partitioned [378]. Hashing [11, 231, 244, 292, 80, 170, 240, 321, 1086, 182, 336, 200, 221, 122, 241]. Having [366]. hB [369]. hB-Tree [369]. HDFS [944]. Headed [342]. Heavy [1044, 771]. Height [1049, 250]. Height-balanced [250]. Heraclitus [474]. Heterogeneity [440]. Heterogeneous [441, 694]. Heuristic [136]. Heuristics [83]. Hierarchical [53, 1006, 193, 707, 116, 242, 24, 854]. Hierarchies [33]. Hierarchy [81]. High [1050, 227, 401, 275, 829, 854, 834, 35, 1007, 980, 134, 524, 569, 765]. high-dimensional [765]. High-Level [980, 134, 834]. High-Performance [275, 1007, 854, 569]. High-Throughput [829]. Higher [406]. highly [612]. highly-available [612]. histogram [631]. Histograms [432, 551]. Historic [997]. Historical [437]. Histories [268]. History [456]. Hitters [1044, 771]. Hoc [117, 884, 957, 1076]. Homogeneous [323, 431]. Homograph [1071]. Horizontally [263]. Horn [371]. Hosting [243]. hot [599]. Human [158, 63, 141]. Hybrid [944, 856]. Hypertree [1074].

I/O [770, 535, 888, 341, 827, 245]. I/O-Efficient [888, 827]. ICDT [709, 639, 716]. Identifying [1077]. iDistance [602]. IFO [304]. Ignorance [449]. Impact [979]. Implementation [139, 317, 152, 338, 414, 133, 12, 252, 514, 582, 724, 477]. Implementing [65, 470, 57, 87, 528]. Implication [471, 814]. Implications [307, 212, 86]. implicit [717]. Implied [157]. Imprecise [343]. Improved [128]. Improving [676, 515, 643, 764, 63, 920, 134]. In-Situ [916]. INCINERATE [454]. Inclusion [969, 848]. Incomplete [892, 78, 267, 315, 165, 627, 507]. Inconsistencies [928, 1020, 695]. inconsistent [699, 759]. incorporate [541]. Incorporating [738, 167]. Increase [345]. Increasing [449, 236]. Incremental [455, 590, 774, 851, 1052, 1027, 610, 386, 1023, 847, 551]. Indefinite [350]. Independence [399]. Independent [351, 39, 638]. Independently [73]. indeterminacy [492]. Index [433, 491, 64, 83, 570, 143, 182, 287, 341, 329, 443, 880, 762, 708]. Index-driven [570]. Index-Sequential [143]. indexability [641]. Indexed [288]. Indexes [26, 591, 675, 747]. Indexing [513, 1038, 194, 369, 964, 628, 156, 863, 549, 602, 781, 632, 841, 592, 618, 505]. Indices [293, 622]. Inequalities [1028]. Inexpressiveness [1045]. Inference [756, 465, 1027, 220, 755]. Inferential [273, 44]. Inferring [1001, 929]. Infinite [419]. Influence [69, 1058]. Information [361, 260, 284, 928, 299, 693, 330, 70, 1042, 503, 78, 350, 6, 411, 267, 359, 306, 441, 909, 360, 315, 165, 295, 443, 627, 649, 772, 725, 750, 526, 778, 518]. information-theoretic [750]. informative [764]. INGRES [12].

Inner [956]. Input [733]. Input-sensitive [733]. inputs [678]. Insertion [1001, 1048]. Insertion-based [1048]. Insights [1079]. Instance [413]. Instances [917, 526]. Instant [776]. instruction [643]. Integer [470, 1003, 820, 494]. Integrated [259, 266, 198, 760, 648]. Integrating [626, 532]. Integration [915, 608]. Integrity [444, 456, 947, 132, 27, 344, 614, 731]. Intelligent [410]. intensional [593]. Intensive [100]. Interaction [1012]. Interactions [63]. Interactive [249, 1004]. interactively [681]. Interface [324, 49, 245, 603]. Interfaces [413, 630]. Interleaving [975]. intermediates [769]. Internal [213]. International [1094, 275, 1092, 1088, 1089, 1090]. Internet [510]. Interoperability [441, 544]. Interpretations [157]. Interpreter [151]. Interval [33]. Introduction [679, 639, 686, 713, 710, 92, 716, 683]. inverses [700]. Inverted [505, 18]. Inverting [684]. investment [532]. Invited [887, 768, 794, 853]. Irrelevant [340]. Isolation [1059, 740, 605, 731, 546]. Issue [887, 709, 768, 794, 853, 739]. Issues [361, 78]. Issuing [930]. Italy [1090]. items [599]. Itemsets [881]. Iterative [327, 563, 522].

Jobs [873]. Join

[335, 919, 347, 264, 432, 960, 996, 885, 907, 391, 884, 239, 314, 263, 270, 211, 293, 967, 733, 676, 633, 563, 666, 655, 803, 704, 868]. Join-Like [885]. Joined [171]. Joining [391]. Joins [1054, 1078, 1000, 226, 945, 1036, 990, 944, 188, 79, 626, 696, 842, 543, 558, 751, 789]. joint [661]. Journal [931]. Judicious [903].

K-d [144]. KBs [834]. Kent [203]. Kernel [949]. Key [472, 269, 960, 730]. Key-Based [472]. Keys [889, 942, 131]. Keyword [910, 874, 737, 690, 755, 778, 840]. Kinds [382]. know [753]. Knowledge [345, 184, 414, 349, 442]. Korth [403]. KTELO [1013].

Label [1038]. Label-constrained [1038]. labeled [749]. Lake [1071]. Language [370, 151, 249, 972, 317, 260, 158, 474, 49, 414, 205, 1019, 194, 280, 235, 100, 346, 35, 980, 120, 88, 294, 354, 141, 603, 503, 689, 47, 334]. Languages [998, 989, 832, 243, 941, 1008, 152, 437, 323, 157, 255, 252, 450, 811, 717, 651, 782, 493, 608]. Large [186, 935, 1092, 1088, 1089, 27, 223, 9, 1090, 15, 270, 513, 779, 549, 457, 781, 706, 635, 498, 855, 672]. large-scale [706, 498]. Layered [915]. Lazy [930]. Learnability [1060]. Learning [919, 861, 937, 1018, 1034, 993, 1039, 488]. Least [89]. Leaves [144, 418]. LEMP [956]. Level [52, 35, 980, 1076, 134, 535, 635, 499, 834, 747]. level-two [635]. Levels [1059, 9, 550]. lexicographic [821]. LH* [476]. LH*RS [612]. libraries [504]. Lightweight [1003, 985, 891]. Like [885, 164]. Limitations [951, 237, 70, 697]. Limited [461]. Limiting [432]. Line [169, 1041]. lineage [525]. Linear [975, 109, 292, 170, 240, 321, 1086, 221, 72, 1023, 356, 908, 848, 820]. Linear-Time [356]. Linearize [364]. Linearly [45]. Linked [197]. Linking [940, 1075]. List [197]. Listing [888]. Lists [56]. lived [478]. Load [384]. Loading [916, 578, 1025]. localization [699]. Locally [549, 767]. Location [914, 744, 763, 784]. location-based [784]. Lock [890]. Locked [31]. Locking [358, 279, 142, 97, 397, 34, 75, 438, 257, 429, 301, 761]. Locks [173]. log [478]. Logarithmic [1074]. Logging [397, 802, 478]. Logic [936, 355, 150, 923, 296, 364, 536, 787, 553, 610, 810]. Logic-Based [355]. Logical [266, 213, 424, 252, 481, 767, 814, 535, 547]. logics [745, 556]. Logs [1065]. Long [447, 478]. Long-Duration [447]. Looking [1076]. Loss [299]. Lossless [314]. Lower [1059]. LRU [341]. LSM [992, 1070]. LSM-Engines [1070]. LSM-Trees [992].

M. [403]. MA [1088]. Machine

[332, 1034, 38, 211, 752]. Machines [124]. MacroBase [991]. Magic [467]. Main [270, 575]. Maintainability [399]. Maintainable [389]. Maintaining [337, 357, 653, 1022]. Maintenance [444, 633, 15, 847, 589, 551, 610]. Majority [74]. Making [605]. Management [8, 62, 208, 20, 229, 1094, 197, 114, 338, 414, 380, 51, 50, 276, 34, 281, 7, 302, 192, 374, 514, 726, 746, 478, 542, 587, 477, 504, 647]. Manager [31, 890, 19, 475]. Managers [7]. Managing [412, 584, 838, 673]. Manipulating [43]. Manipulation [108]. Many [969]. many-core [796]. map [839, 855, 581]. Mapping [1004, 963, 742, 835, 609, 548]. mappings [797, 861, 619, 684, 700, 711, 663]. MapReduce [828, 873]. March [1091]. Marrying [972]. Massive [1082, 961, 795]. Massive-Scale [961]. Massively [1000]. Match [417, 73, 11, 76, 199, 200]. Matcher [177]. Matching [193, 912, 1024, 920, 851, 630, 835, 865, 791, 629, 754, 868]. Materialization [948, 924]. Materialize [367]. Mathematical [172, 864]. Matrices [1008, 867]. matter [811]. Maximal [893, 176, 795]. Maximization [1058, 807]. Maximizing [882, 825]. Maximum [956]. May [1094]. Maybe [202, 350]. Meaning [84]. Means [65, 362, 179]. Measures [1083, 669, 803]. Measuring [954]. Mechanism [59, 14, 106]. Mechanisms [259, 94, 112, 129]. Meets [1069]. Membership [103, 975]. Memories [51, 270]. Memory [31, 882, 133, 19, 1044, 352, 576, 586, 685, 860, 5, 575]. Memory-Resident [352]. Merge [56, 655, 687, 659]. merge-reduce [687].

Mergeable [859]. Merging [992]. Meta [916]. Meta-Operator [916]. metadata [587, 608]. Meter [953]. Method [379, 80, 829, 369, 397, 278, 386, 409, 436, 533, 502, 602, 842]. Methodology [365, 306, 319, 659]. Methods [951, 1053, 320, 123, 308, 691, 483, 497, 516, 840]. Metric [696, 513, 554, 570]. metrics [616, 694]. Metrology [954]. middleware [664]. Minimal [298, 990]. Minimization [831, 83, 136]. Minimum [26]. Mining [905, 1005, 665, 1053, 867, 893, 691, 650, 585, 537, 630, 703, 566, 775, 806, 672]. Missing [933]. Mixed [470, 835]. MMSNP [894]. mobile [506]. Mobility [1048, 1030]. MobilityDB [1030]. MOD [983]. Model [304, 282, 168, 317, 951, 62, 3, 461, 84, 53, 473, 323, 381, 130, 414, 22, 475, 424, 272, 341, 217, 1069, 373, 120, 302, 315, 443, 25, 500, 524, 454, 860, 447, 523, 21, 541, 480, 285]. Model-Based [315]. Modeling [307, 254, 258, 673, 295, 613, 526, 647, 517, 595]. Models [446, 877, 172, 70, 1018, 874, 239, 395, 585, 782, 552]. modern [796]. Modified [436]. modifiers [530]. Modular [376, 405, 192]. Modularization [464]. Moment [1033]. Moments [997]. Monadic [943]. Monitored [947]. Monitoring [82, 877, 985, 721, 682]. Monotone [1029]. Monotonicity [918]. Monte [792]. most [599, 560, 813]. Moving [840, 762, 521, 633, 628, 584]. MR [388]. Multi [36, 218, 675, 747]. multi-component [747]. multi-level [747]. Multi-resolution [675]. Multi-Valued [36, 218]. Multiattribute [144, 369]. Multibucket [418]. Multicast [388, 375]. multicore [796]. multicore/many [796]. multicore/many-core [796]. Multicores [890]. multidatabase [544]. Multidatabases [440, 499]. Multidimensional [284, 107, 118, 654, 632, 113, 674, 609]. Multikey [207, 308]. Multilevel [196, 374].

Multimedia [414]. Multiple

[901, 149, 81, 885, 907, 873, 209, 382, 311, 74, 286, 89, 550, 669, 538, 661, 694, 481]. Multiple-Disk [149]. Multiple-Query [311]. Multiple-Valued [89]. Multiplicative [986]. Multiprocessing [192]. Multiprocessor [124, 210, 211]. MULTISAFE [192]. multiscale [653]. Multivalued [103]. Multiversion [195, 589, 552]. Multiversion-based [589]. Multiview [498]. Multiway [543, 1072, 135, 869]. MVDs [298].

NAIL [519]. Naïve [892]. NaLIX [689]. names [835]. Native [587]. Natural [49, 205, 194, 836, 689, 334]. Navigational [981]. Near [1050, 789]. near-duplicate [789]. Nearest [874, 783, 602, 606, 641, 765]. nearest-neighbor [783]. Necessary [364]. Negation [926]. Negative [920]. Neighbor [1050, 874, 783, 602, 606, 765, 806]. neighbor-based [806]. neighbors [641]. Nested [455, 448, 164, 1019, 425, 466, 289, 396, 322, 403, 677, 644]. Nested-Transaction [448]. Nesting [226, 968]. Network [990, 255, 734]. Networking [918]. Networks [906, 914, 163, 962, 56, 175, 6, 830, 976, 921, 795, 758, 596, 634]. Next [1034, 373]. Next-Generation [373]. NFQL [334]. NJ [1094]. nn [633]. No [1078, 244, 403, 388]. node [846, 662]. node-selecting [846, 662]. Noisy [896]. Non [358, 632]. non-ordered [632]. Non-Two-Phase [358]. Nondeterministic [1032]. Nonequivalent [1080]. Nonprocedural [158, 141]. nonrecursive [651]. Normal [171, 66, 17, 406, 36, 131, 128, 466, 289, 166, 577, 489, 583]. Normalization [53, 320]. Note [50, 387]. November [1090]. NT [447]. NT/PV [447]. Null [425, 145]. nulls [785]. Number [46, 601, 670]. Numeric [1020]. numerical [759].

O [388, 770, 535, 341, 245]. O-Efficient [888, 827]. Object [139, 394, 268, 338, 414, 435, 382, 489, 310, 1095, 556, 762, 539, 531, 502, 613, 547, 804]. Object- [382]. Object-Oriented [139, 338, 414, 435, 1095, 556, 502, 547, 489]. object-relational [613]. objectbase [480]. objective [1068]. Objects [254, 290, 947, 468, 176, 880, 727, 521, 804, 628, 623, 584, 508]. Oblivious [942, 866, 697]. Observing [836]. obstacles [783]. October [1090]. ODE [732]. ODMG [512]. Offset [1064]. Offset-value [1064]. OLAP [614, 661]. OLTP [643, 872]. On-Line [169]. On-the-Fly [921]. One [321, 828, 1036]. One-Access [321]. One-Pass [828]. One-time [1036]. Online [903, 996, 816, 921, 762, 708]. Only [154]. Ontological [886, 517]. ontologies [647]. Ontology [894, 732, 647]. Ontology-assisted [732]. Ontology-Based [894]. OODBMS [490]. opaque [835]. Open [972, 422, 1031]. Open-source [1031]. Operational [177]. Operations [190, 394, 162, 347, 142, 32, 372, 263, 125]. Operator [916, 885]. Operators [202]. opportunities [697]. Optical [290, 245]. Optimal [417, 73, 1054, 1078, 148, 914, 992, 309, 299, 1000, 226, 432, 402, 1015, 32, 6, 1036, 199, 239, 751, 754, 771, 707]. Optimality [1025, 121, 871]. Optimism [225]. Optimistic [353, 123]. Optimization [85, 335, 400, 620, 355, 509, 1068, 421, 309, 271, 886, 497, 175, 362, 366, 329, 263, 311, 409, 463, 352, 71, 556, 677, 479, 652, 621, 532, 522, 858, 634, 841]. Optimizations [998, 924]. optimize [578]. Optimized [1049, 668, 537]. optimizer [680, 725]. Optimizing [837, 1001, 531, 664, 164, 1036, 378, 622, 908, 579]. Optimum [64, 45]. OQL [512]. Order [936, 1055, 1053, 269, 1067, 250, 226, 296, 1072, 867, 821, 610, 494, 810, 619]. Order-Preserving [269, 1053, 1067, 867].

Orderability [426]. ordered [749, 541, 632]. Orders [1063]. Oregon [1091]. Organization [144, 232, 95, 193, 326, 336, 89, 1, 18]. Organizations [287, 418, 122]. Oriented [282, 139, 338, 414, 273, 435, 382, 1095, 556, 502, 547, 489]. Otherwise [439]. Out-of-core [672]. Outerjoin [479]. Outlier [961]. outlying [727]. Output [1000, 775, 360]. Output-Optimal [1000]. Outsourced [884, 757, 735]. Overcoming [592]. Overflow [222, 143, 240]. Overflow-Handling [240]. Overhead [301, 586]. overlapping [552]. Overview [283, 519]. Owner [62]. Owner-Coupled [62].

packed [685]. packed-memory [685]. Packing [1025]. Padding [1068]. Page [410, 81, 567, 624]. page-server [624]. PageRank [1010]. Pages [197]. Paging [41]. Painless [880]. pair [765]. Paper [904, 931]. Papers [253, 931, 679, 768, 794, 853]. Parallel [998, 190, 189, 916, 1074, 1000, 214, 368, 770]. Parallelism [98]. Parallelizability [1025]. Parallelization [1017]. Parallelizing [994]. Parameter [993]. parsing [18]. Partial [417, 73, 11, 76, 833, 170, 287, 418, 397, 199, 200, 1072]. Partial-Match [73, 11, 76, 199, 200]. Partially [31]. partition [842, 285]. partition-based [842]. Partitioned [472, 225, 337, 291, 279, 331, 263, 378, 236]. Partitioning [234, 238, 58, 632, 592]. Pass [828]. past [628]. PATAXÓ [642]. Path [832, 569, 57, 781, 1009, 1051]. Path-tree [781]. Paths [187, 830, 857]. Pattern [984, 177, 943, 851, 703, 865, 629, 819, 806]. Patterns [937, 798, 601]. Peer [658]. Penalties [1001]. people [808]. Perfect [336]. Perform [239]. Performance [408, 259, 307, 363, 31, 189, 212, 290, 227, 288, 275, 95, 261, 407, 415, 170, 326, 369, 342, 418, 38, 30, 423, 228, 1007, 460, 19, 183, 257, 429, 360, 548, 676, 569, 485, 643, 854]. periodicity [500]. Permission [353]. permissions [562]. Persistent [1070, 1057]. Personalized [1010, 800]. Personalizing [758]. perspective [621]. Perturbation [436, 516]. Perturbations [955]. Phase [358, 279, 429]. Physical [168, 687, 213, 313, 27]. Pipelined [214]. Pipelines [1079]. Pipelining [87]. Placement [974, 990, 388, 375, 532]. plain [870]. Plan [934]. Planning [40, 1048]. Platform [828]. Platforms [1073]. PODS [709, 739, 656, 713, 683]. Pointer [42]. Points [148, 45, 633]. policies [636, 567, 757, 538]. Policy [32, 342, 388, 375]. Polymorphism [465]. Portable [205]. Portland [1091]. positions [628]. Possible [1053]. Possible-world [1053]. PostGIS [1030]. PostgreSQL [1030]. Power [1008, 650]. Practical [169, 978, 206, 439]. practice [863]. Precisely [466]. Predicate [62, 309, 33, 569]. predicates [509, 585, 701]. Predicting [452, 1034]. Prediction [877]. predictive [565]. Preference [568, 850]. preferences [758, 793]. Prefetching [1073, 54, 620, 676]. Prefix [23]. presence [783]. Present [904, 628]. Preserving [269, 884, 691, 1053, 693, 867, 777, 1067, 499, 728]. Pricing [895]. Primary [960]. Principles [1091, 229, 163, 374]. Prioritizing [991]. Privacy [763, 884, 908, 691, 744, 729, 864, 775, 728]. Privacy-aware [763]. Privacy-Preserving [884, 691, 728]. Private [1068, 978, 883, 895, 939, 976, 1013, 784]. PrivBayes [976]. Probabilistic [973, 875, 649, 534, 473, 539, 926, 573, 964, 925, 702, 356, 799, 738, 867, 715, 487, 804]. probabilities [786]. Probability [256, 866]. Probing [240]. Problem [936, 103, 157, 273, 220, 58, 814, 813]. Problem-Oriented [273]. Problems [66, 471, 631]. ProbView [487]. Procedural [158, 141]. Procedure [96]. Procedures [297]. Proceedings [1092, 1091, 1094, 1088]. Process [91, 126, 55]. processes [546]. Processing [971, 400, 140, 910, 916, 833, 934, 237, 184, 591, 420, 272, 878, 907, 925, 346, 917, 445, 898, 265, 270, 734, 957, 29, 1031, 13, 927, 733, 800, 692, 524, 668, 722, 744, 600, 801, 604, 712, 596, 854, 849, 694, 856, 816, 609]. Processor [40, 124, 56, 214, 30]. Processors [189, 50, 238, 741, 796]. Product [149, 956]. Production [444, 535]. Programming [973, 470, 243, 465, 474, 88, 134, 820, 522]. Programs [974, 327, 1023, 125, 364, 477]. Progressive [594]. Project [235]. Projection [328, 984, 387, 310, 837]. Projections [427]. Prolog [235]. Propagation [432, 825]. Properties [160, 325, 81, 871, 727]. property [857]. Proportionality [1066]. Propositional [951]. Protection [69]. Protocol [412, 358, 97]. Protocols [423, 422, 301, 506, 528, 498]. Prototypes [360]. Provenance [831, 1079, 717, 852, 780]. Proximity [803, 1061]. Proxy [582]. Proxy-based [582]. Pseudo [670]. Pseudo-random [670]. publishing [714, 555, 763, 728]. Pufferfish [864]. pull [636]. pull-based [636]. PV [447].

Quadtrees [1054]. Quality [912, 725]. quality-aware [725]. Quantification [457]. Quantile [1033]. quantitative [703]. Quasi [700, 672]. quasi-clique [672]. Quasi-inverses [700]. Queries [417, 832, 983, 919, 910, 1035, 1055, 1063, 932, 117, 914, 1038, 930, 48, 109, 105, 949, 926, 327, 892, 174, 390, 420, 1085, 459, 1022, 146, 1028, 960, 885, 964, 342, 1009, 943, 1036, 907, 385, 986, 427, 1051, 72, 378, 178, 1060, 377, 1010, 1061, 908, 1080, 179, 799, 512, 576, 586, 513, 548, 843, 509, 846, 568, 637, 844, 531, 773, 604, 836, 822, 497, 664, 633, 862, 758, 724, 782, 858, 540, 18, 580, 671, 371, 766, 804, 662, 606, 661, 579, 694, 734, 702, 565, 559, 819, 840, 863]. Query [972, 667, 455, 951, 400, 140, 40, 1008, 152, 982, 201, 355, 158, 458, 419, 979, 437, 833, 954, 876, 421, 934, 309, 334, 323, 271, 877, 886, 777, 968, 205, 175, 1042, 362, 164, 214, 960, 272, 878, 1034, 993, 899, 346, 891, 634, 898, 265, 311, 980, 409, 294, 255, 1039, 252, 220, 354, 141, 352, 450, 13, 955, 71, 46, 927, 733, 556, 524, 620, 603, 717, 677, 585, 649, 668, 722, 699, 479, 801, 630, 741, 652, 680, 712, 621, 651, 503, 532, 522, 596, 645, 721, 849, 629, 493, 723, 545, 810, 816, 735]. query [609, 744]. Query-Answers [993]. Query-preserving [777]. Query-sensitive [667]. Querying [989, 779, 536, 759, 349, 587, 627, 837, 698, 521, 808, 681]. Quintary [107]. Quorum [412, 448, 291]. quorums [564].

R [8, 117]. R* [276]. Radii [901]. RAMP [938]. Random [4, 277, 1055, 105, 935, 996, 436, 921, 516, 670]. Random-Order [1055]. Range [882, 978, 674, 1016, 870, 701]. Rank [1056, 652, 803, 751]. rank-aware [652]. Ranked [1063, 678]. Ranking [875, 649, 702, 784]. ranking-aggregate [702]. Rapid [133]. Rationale [343]. RDBMS [574, 595]. RDF [972, 981, 834]. Reachability [1038, 781]. Reactive [392]. Read [154]. Read-Only [154]. Readings [1093, 1095]. Real [408]. Real-Time [408]. Realization [374]. Reasoning [404, 790, 911, 395, 500, 820, 730, 488]. rebalancing [869]. receivers [533]. Recognition [1046]. reconstruction [767]. Record [186, 70, 37, 223, 368, 247]. Record-Based [70]. Records [290, 46]. Recovery [169, 259, 98, 230, 189, 380, 397, 228, 520, 742, 802, 624]. Recursion [1074, 469]. Recursive [126, 221, 364]. recycling [769]. reduce [586, 687].

Reduced [298]. Reducing [915]. reduction [549]. Redundancy [76, 466, 89, 750]. Redundant [81]. Referees [495]. Reference [213, 535]. referential [553]. refinement [687]. refresh [567]. regions [843, 809]. Register [51]. Regression [306]. Regret [1056, 986, 1039]. Regret-bounded [1039]. Regular [941, 975, 1014, 1009, 943, 1051, 920, 756, 848, 857]. Regular-Expression [920]. regulating [757]. Reinforcement [1039]. Related [66]. Relation [151, 328, 160, 347, 137, 219, 176, 217, 372, 387, 35, 204, 203]. Relation-Based [151]. Relational [971, 85, 8, 413, 65, 171, 455, 138, 66, 103, 266, 202, 190, 324, 82, 453, 437, 84, 406, 162, 53, 473, 949, 36, 131, 313, 327, 323, 174, 14, 57, 741, 390, 226, 157, 383, 146, 1018, 214, 987, 104, 351, 782, 116, 128, 350, 101, 405, 44, 366, 274, 30, 305, 239, 314, 322, 156, 281, 246, 155, 88, 183, 403, 188, 377, 352, 450, 608, 115, 166, 134, 79, 745, 548, 677, 568, 555, 777, 613, 652, 303, 561, 787, 5, 737, 541, 477, 644]. Relations [111, 17, 340, 343, 1075, 459, 425, 101, 466, 391, 289, 39, 378, 145, 507, 635]. Relationship [3, 381, 405, 388, 375, 638, 517]. Relationships [77, 515, 623]. Relative [772]. Release [976]. Reliability [112]. Remote [380, 445]. Removal [1062]. Removing [562]. Reordering [821, 479]. Reorganization [148, 32, 45, 10]. Repair [699, 1065]. Repairability [941]. Repairing [1081, 759, 611]. Repairs [1015]. Repetitive [117]. Replicas [330]. Replicated [412, 230, 393, 337, 357, 449, 416, 388, 375, 731]. Replication [181, 872, 564, 528, 498, 484]. Report [275]. Reporting [1044]. Representation [419, 163, 793]. Representations [998, 899]. Representative [917]. Representatives [1056]. Representing [627, 405, 521]. reprinted [1089]. requests [806]. requirements [576]. rescue [619].

Resident [352]. Resolution [440, 675]. Resolutions [945]. Resolving [181, 220]. Resource [97]. Resources [6]. Respect [46]. Responding [179]. Response [210, 958]. Restart [227, 935]. Restoration [520]. Restrictions [902]. Restructurers [87]. Restructuring [9, 99, 29]. Result [876, 901, 764, 755]. Results [979, 432, 331, 899, 329, 360, 649, 805]. Retrieval [73, 361, 324, 11, 290, 232, 284, 194, 321, 359, 199, 200, 295, 333, 649, 506]. Retrospection [102]. Return [755]. Revealing [178]. Reverse [785, 748, 874]. Reversible [1067]. reviewing [660]. Revisited [1058, 75, 970, 958, 911, 204]. Revisiting [846, 819]. Rewiring [921]. Rewriting [982, 637, 886, 766]. RFID [844, 947]. RFID-data [844]. RFID-Monitored [947]. Rigorous [439]. Ring [1078]. Road [914, 962]. Robust [726, 1064, 934, 878, 881, 384, 730]. Robustness [191, 1059]. role [647]. Rollbacks [397]. Root [1077]. rotating [5]. Roth [403]. Route [1048]. Routing [384]. rows [821]. Rule [370, 400, 409, 588, 477]. Rule-Based [370, 400]. Rules [452, 464, 444, 461, 383, 527, 537, 553].

S [834]. Safe [493, 494]. Safety [339, 377]. Sample [105]. Sampling [1050, 1058, 1085, 668, 705]. Sampling-Based [1085, 705]. Satisfiability [471]. Scalable [938, 712, 890, 995, 828, 993, 939, 733, 838, 476, 612]. Scale [905, 961, 706, 498]. Scaling [360]. SCALLA [828]. Scan [342]. SCANRAW [916]. Scans [341]. Schedulers [248]. Scheduling [408, 1073]. Schema [283, 324, 932, 126, 343, 835, 966, 99, 963, 797, 742, 501, 861, 619, 684, 700, 693, 502, 630, 523, 480, 640]. Schema-Mapping [963]. Schemas [66, 902, 839]. SchemaSQL [544]. Schemata [116, 115, 147, 166, 489]. Scheme [351, 321, 431, 89, 801, 781].

Schemes [266, 389, 849]. Science [1079]. scientific [572, 675]. Scope [434, 2, 451]. score [587]. Scotty [1031]. SDD [94, 93, 140, 112, 129, 92]. SDD-1 [94, 93, 140, 112, 129, 92]. SDM [130]. Search [4, 978, 284, 108, 167, 874, 224, 44, 60, 1072, 312, 956, 513, 773, 783, 570, 690, 602, 796, 689, 748, 18, 764, 755, 805, 737, 808, 869, 688, 674, 765, 778, 657]. Searching [90, 550, 554, 348, 242, 251, 16, 286, 870]. Second [619]. Second-order [619]. Secondary [26]. Secrets [178]. section [656, 639, 686, 713, 710, 716, 683]. Secure [105, 69, 435, 812, 939, 192, 450]. Security [48, 68, 22, 331, 516, 113, 118, 436, 233, 106]. Segmentations [37, 223]. Segmented [27]. Selected [253]. selecting [846, 662]. Selection [26, 64, 1003, 223, 44, 575, 302, 924, 548, 804]. Selective [443]. selectivity [671]. Self [613, 338, 960, 838]. Self-Adaptive [338]. Self-Join-Free [960]. self-managing [838]. Self-tuning [613]. Semantic [304, 440, 355, 345, 184, 381, 163, 130, 78, 441, 409, 125, 395, 483, 556, 620, 989, 858]. semantic-based [483]. Semantics [426, 398, 138, 875, 458, 185, 462, 325, 724, 425, 176, 385, 736, 477, 255, 482, 303, 553]. Semantics-Based [426, 398]. Semijoin [211]. semirings [862]. semistructured [545]. sense [488]. sensitive [733, 667, 796]. sensor [726, 596]. sentences [560]. Separators [321]. September [1088]. Sequence [997, 905, 1005, 829, 579]. Sequenced [949]. Sequences [268, 782]. Sequencing [629]. Sequential [994, 348, 143, 242, 251, 16]. Sequentiality [54]. serial [770]. Serializability [300, 499]. Serializable [740, 605]. Series [959, 549, 791]. server [485, 624, 490]. service [681]. service-accessed [681]. Services [330, 744, 784]. Serving [342]. Set [62, 271, 305, 533, 558]. Set-Valued [305]. Sets [370, 1029, 459, 218, 588, 779, 566, 598].

Several [261]. Shadowed [197]. Shared [37, 223, 1048, 806, 812]. Sharing [1067, 873, 423, 384, 1080, 569, 852, 490]. Shift [51]. Shift-Register [51]. shortcomings [592]. Shortest [830, 610]. Should [970]. Shrink [655]. SIDEREUS [283]. SIFT [518]. SIGACT [1091]. SIGACT-SIGMOD [1091]. SIGMOD [1091, 253, 1094, 710, 709, 739, 656, 679, 904]. SIGMOD/PODS [739, 656]. SIGMOD/PODS/ICDT [709]. Signature [379, 472, 299, 402, 849, 505]. Silberschatz [403]. SilkRoute [555]. Similar [193]. Similar-String [193]. Similarity [1000, 967, 513, 570, 696, 842, 849, 868, 789, 657]. Simple [906, 335, 406, 174, 326, 1009, 557]. Simplicity [966]. Simplification [383, 1041, 479]. Simplified [160]. SimRank [967]. SimRank-Based [967]. simulation [865]. Single [1010, 660]. Single- [660]. Single-Source [1010]. Situ [916]. Sixth [1089]. Size [432, 899, 704]. Sized [942]. Sizes [347, 81]. sketch [670]. sketch-based [670]. Sketches [704]. SkinnerDB [1039]. Skycube [948, 816]. Skyline [948, 907, 827, 786, 720, 748, 594, 654, 734, 816]. Skyline-Join [907]. Skylines [875, 1029, 815]. Sliding [898, 653, 724]. Sloth [930]. Small [48, 735]. Smart [953]. Snapshot [731, 740, 605]. snippets [764]. Social [929, 921]. Soft [1081]. Software [376, 132]. Solid [903]. Solid-State [903]. solution [760, 648]. Solutions [157]. Solving [471, 427]. Some [244, 35]. sort [720, 655]. sort-based [720]. Sort-Merge-Shrink [655]. Sorted [56]. Sorting [1062, 1064]. sound [780]. Source [1010, 510, 1031]. Sources [907, 779, 589, 626, 681]. Space [1078, 1047, 771, 898, 1025, 121, 463, 871, 696, 621, 765]. space-complexity [621]. Space-Efficient [1047, 898]. Space-Filling [871, 1025]. Space-optimal [771]. Space-Optimality [121]. Spaces [295, 524, 513, 554, 570, 632, 734, 688]. Spanners [1032, 1014]. SPARQL [968, 857, 736]. Sparse [1018]. Spatial [910, 1066, 1002, 114, 666, 874, 559, 830, 880, 511, 633, 563, 635, 543, 606, 592, 623, 734, 754, 840]. Spatial-Keyword [874]. spatio [565]. spatio-temporal [565]. Spatiotemporal [929]. special [656, 639, 686, 713, 710, 709, 739, 716, 683]. Specialized [65, 56]. specific [560]. Specification [1004, 514, 811, 755]. Specified [73]. Specifying [468]. Speed [1040]. Splitting [309, 241]. Spyratos [303]. SQL [677, 1068, 836, 814, 1059, 164, 544, 923, 280, 1034, 385, 610, 939, 595]. SQL-like [164]. standard [603]. Star [175]. State [903, 272]. statement [530]. States [216]. Static [743, 452, 984, 362, 858, 691, 527]. Statistical [48, 119, 68, 109, 96, 105, 390, 22, 431, 331, 306, 118, 72, 233, 165, 179, 106, 715, 113]. statistics [653]. Steiner [220]. Steps [206]. Still [970]. Stochastic [421, 792, 815]. Storage [903, 171, 258, 1073, 181, 81, 41, 133, 286, 478]. Store [151, 1007, 769]. Stored [944]. Strategies [124, 201, 194, 223, 645, 374, 548]. Strategy [1003, 44, 13, 806]. stratified [668, 494]. Stream [1040, 1031, 692, 600, 746, 591, 653]. Streaming [878, 631, 860, 607, 806]. Streams [985, 891, 961, 576, 586, 620, 698, 591, 715, 557, 724, 782, 760, 737, 648, 682, 735]. Strength [929]. String [193, 912, 842, 671, 868]. Strings [920]. Strong [865, 583, 771]. Strongly [249, 243, 872]. Strongly-Typed [249, 243]. structural [778]. Structure [57, 883, 107, 224, 207, 312, 476, 612, 618]. Structured [832, 462, 1065, 348, 251, 16, 805, 598]. Structures [43, 126, 405, 24, 443, 552, 657]. Studies [460]. Study [894, 1033, 1076].

Subgraph [1043, 1036]. Subgraphs [893, 1083]. Subkeys [127]. Submatrices [1053, 867]. Suboptimality [1077]. Subrelations [343]. Subsampling [881]. Subsequence [1005, 791]. subsets [598]. subspace [654, 816]. Succinct [1016]. Sufficient [364]. suffix [770]. Sum [882]. Summaries [1057, 859]. summarizability [614]. Summarizing [635]. Summary [431, 346, 863]. Summary-Table-by-Example [346]. Superimposed [308]. Support [117, 949, 20, 309, 133, 382, 210, 409, 506, 538, 587]. Supporting [1079, 492, 746, 397, 500, 844]. Surface [1061]. Survey [932, 1003, 1002, 761, 802, 793]. Swapping [206]. Symmetric [207]. Symposium [1091]. Synchronization [91, 932, 55]. Synonyms [912]. Synopses [621, 573, 616, 707, 646, 723, 735]. syntax [494]. Synthesis [446, 430, 1023]. Synthesizing [17]. System [361, 139, 400, 94, 93, 140, 40, 31, 208, 127, 20, 14, 57, 112, 338, 414, 205, 219, 449, 273, 342, 44, 276, 519, 274, 34, 52, 92, 382, 19, 957, 155, 102, 183, 297, 352, 134, 692, 792, 503, 487, 596, 721, 477, 518, 8, 117, 219, 681, 29]. System/U [219]. Systems [1091, 318, 258, 98, 91, 260, 244, 150, 225, 1006, 149, 191, 213, 227, 262, 271, 275, 448, 471, 1073, 447, 81, 41, 107, 953, 67, 199, 946, 445, 373, 423, 314, 52, 281, 55, 441, 263, 270, 236, 54, 1093, 302, 403, 872, 159, 1031, 360, 388, 165, 147, 1095, 743, 512, 556, 514, 726, 823, 746, 624, 850, 18, 594, 634, 480, 579, 751, 793, 2].

Table [1086, 346, 814]. Tableaux [216, 161]. Tables [431, 969, 737]. Tail [469]. Tailored [1037]. tampering [719]. tasks [680]. Taxonomic [404]. Taxonomy [435]. Technical [911, 204, 203]. Technique [152, 87, 367, 449, 265]. Techniques [452, 464, 132, 346, 228, 308, 501, 761, 802, 497, 666, 706]. Technology [290].

Temporal [530, 456, 949, 323, 411, 508, 500, 536, 534, 565, 481, 701, 294]. Temporally [282]. Tensors [1018]. Term [167]. Terms [77]. Terrain [1061]. Testing [86]. Text [365, 510, 673, 680, 708, 587, 505]. text-centric [680]. text-source [510]. Their [22, 216, 187, 15, 33, 836, 915, 583]. Them [1050]. Theoretic [359, 239, 839, 750, 1012]. Theoretical [1025]. Theory [195, 1042, 895, 529, 329, 641, 188, 863, 79, 767, 773, 504]. Theseus [88]. thinning [855]. Third [17, 128]. Thrashing [429]. Threat [68]. Three [163, 923, 9, 110]. Three-Valued [923]. threshold [682]. Throughput [829]. Thwarting [753]. Tiered [1073]. Tight [601]. Tightened [1058]. Time [408, 185, 975, 222, 954, 1001, 389, 420, 362, 959, 898, 121, 210, 399, 356, 549, 492, 1036, 791]. Time- [898, 121]. Time-Constrained [420]. Time-decaying [1001]. Time-Series [959, 791]. Timely [1070, 1044]. Times [1001]. TinyDB [596]. TODS [495, 110, 904, 931, 709, 768, 794, 739]. token [868]. tolerance [692]. Tolerant [927, 868]. Tolerating [422]. Tool [376, 1086]. Tools [439]. Top [548, 843, 896, 898, 800, 664, 760, 678, 580, 648, 702]. Top- [548, 843, 896, 898, 800, 760, 678, 580, 648, 702]. top-k [664]. Topmost [948]. Topological [623, 635]. topology [865]. TPC [535]. TQUEL [294]. Traces [410]. Tracing [525]. Tracker [68, 96]. tracking [599]. Tractable [1063]. Tradeoffs [393, 416]. Traffic [990]. Trajectories [947]. Trajectory [962, 1041, 961]. Trans [244, 403, 388]. Transaction [446, 237, 184, 275, 448, 248, 447, 276, 397, 445, 460, 339, 374, 384, 625]. Transaction-Processing [445]. Transactional [485, 278, 838, 546]. Transactions [408, 938, 277, 117, 345, 154, 2, 383, 430, 872, 159, 1076, 483, 478]. Transducer [906]. transducers [714].

Transformation [118, 125, 113]. Transformational [40]. Transformations [269, 439]. Transition [272]. Transitive [363, 428, 367, 610]. Translation [162, 327, 366, 377]. Transparent [753]. Transposed [90]. Traversal [428]. Traversals [125]. Tree [412, 941, 937, 144, 1084, 174, 95, 348, 943, 900, 251, 427, 16, 798, 600, 770, 761, 802, 602, 624, 781, 796, 819, 415, 369]. tree-edit [600]. Tree-Structured [348, 251]. Trees [984, 4, 942, 992, 250, 108, 107, 220, 749, 23, 846, 135, 820, 818, 662, 1025, 869, 142, 121]. TriAL [981]. Triangle [888, 1022]. Trie [215, 11, 83, 136, 329]. Tries [1049]. Triggers [644]. Triplestores [981]. TuG [723]. tuning [762, 613]. Tuple [950, 268, 1019]. Tuple-Generating [950]. Tuples [1004, 933]. Twelfth [1092]. twigs [629]. Two [358, 279, 904, 342, 429, 537, 635, 499]. two-dimensional [537]. Two-Headed [342]. two-level [499]. Two-Phase [279, 429]. Type [512, 465, 35, 837, 523]. type-based [837]. Type-checking [512]. Typed [249, 243]. Types [353, 274, 848]. tyranny [526].

U [219]. Uncertain [917, 786, 748, 674]. Uncertainty [1053, 584]. Understanding [1071, 553, 807]. Undoable [1052]. UniAD [957]. Unified [1005, 3, 348, 957, 688, 1048, 664]. Unifying [168]. Unions [1035, 1055]. Unique [1060]. Units [286]. Universal [324, 160, 457, 137, 219, 176, 217, 431, 204, 203]. Universal-Scheme [431]. Unknown [979]. Unknowns [979]. unnesting [645]. Unordered [975]. Until [222]. Update [138, 324, 461, 162, 325, 383, 303, 533, 717, 852, 477]. Updates [903, 983, 340, 453, 897, 922, 952, 977, 1011, 1022, 351, 246, 642, 718, 611]. Updating [413, 340, 32, 993, 315]. Upon [932]. upper [601]. USA [1088]. Use [903, 410, 418, 306, 239, 970]. useful [540]. User [919, 69, 319, 1080, 509, 554, 613, 21]. User-defined [1080, 509, 554, 613]. users [506]. Using [1054, 31, 982, 456, 446, 345, 184, 877, 935, 1018, 173, 447, 828, 948, 341, 397, 986, 336, 382, 441, 1044, 778, 220, 1078, 837, 620, 1055, 290, 637, 316, 531, 600, 626, 161, 553, 1036, 632, 200, 807, 819, 611, 816]. Utility [986]. utilizing [764].

Vadalog [1047]. valid [492]. valid-time [492]. Validated [341]. Validating [860]. Validation [42, 590]. Validity [344]. value [1064, 835]. Valued [36, 923, 305, 89, 835, 218]. Values [933, 441, 42, 145, 651, 529]. Variable [329]. Variable-Depth [329]. Variants [231, 244]. Vector [295, 734]. Verification [1006, 339, 475, 735]. Versioning [382]. Versions [209]. Versus [353, 660, 505]. vertex [813]. Vertical [234, 915]. Very [1092, 1088, 1089, 1090, 969, 779]. via [797, 903, 820, 638, 852, 945, 996, 1039, 976]. View [1045, 3, 381, 351, 156, 589, 525, 542, 561, 161]. ViewCache [386]. Views [138, 982, 932, 162, 325, 825, 766, 319, 310, 811, 847, 642, 779, 637, 746, 303, 780, 644, 812, 819]. violations [665]. Virtual [31, 41, 19, 752]. Virtue [930]. Visibility [938]. visual [603]. visualization [855]. VLDB [1092]. VLSI [254]. Voting [357].

W3QL [503]. W3QS [503]. Waiting [407]. Walk [935]. Walks [996, 921]. Wander [996]. Warehouse [944]. Warehouses [903]. warehousing [525]. Watermarking [1067, 777]. Wavelet [616, 573]. wavelets [669]. Weak [413]. Weaker [918]. Web [989, 567, 582, 630, 503, 858, 580, 808, 587, 681, 1076]. Web-accessible [580]. Weight [402]. Well [984]. Well-Designed [984]. WHAM [829]. where [378]. Who [1050, 753]. Wide [582, 636, 503]. Window [898, 1031, 724]. windows [653, 806]. without [744, 869, 178]. workflow [811, 514, 780]. Workload [706, 913]. Workload-aware [706]. Workload-Driven [913]. Workloads [924, 535, 752]. Workshop [275]. World [875, 1053, 788, 582, 503]. World-Wide [503]. Worlds [972, 904]. Worst [1078, 432, 945, 1036, 827, 750]. Worst-Case [432, 827, 1078, 1036, 750]. Write [397]. Write-Ahead [397].

XDB [996]. XLynx [856]. XML [627, 743, 799, 577, 889, 590, 837, 847, 603, 642, 779, 738, 848, 569, 714, 693, 555, 839, 600, 718, 591, 777, 626, 730, 860, 689, 764, 755, 640, 966, 645, 593, 646, 629, 778, 856, 583, 581]. XPath [774, 833, 604, 574, 617, 607]. XQBE [603]. XQ}uery [603, 745, 651, 856]. XSD [889]. XSKETCH [646]. XSQ [607].

years [110].

References

Yao:1977:ABA

[1] S. B. Yao. Approximating block accesses in database organization. Communications of the ACM, 20(4):260– 261, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in [25].

Hsiao:1976:ATD

[2] David K. Hsiao. ACM Transactions on Database Systems: aim and scope. ACM Transactions on Database Systems, 1(1):1–2, 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1976-1-1/p1-hsiao/p1-hsiao.

pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-
1/p1-hsiao/.

Chen:1976:ERM

[3] Peter Pin-Shan S. Chen. The entityrelationship model: Toward a unified view of data. ACM Transactions on Database Systems, 1(1):9– 36, March 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1976-1-1/p9-chen/p9-chen.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1976-1-1/p9-chen/. Reprinted

in [1093].

Bayer:1976:EST

[4] R. Bayer and J. K. Metzger. On the encipherment of search trees and random access files. ACM Transactions on Database Systems, 1(1):37– 52, March 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1976-1-1/p37-bayer/p37-bayer.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1976-1-1/p37-bayer/

. Also published in [1088, p. 508–510].

Lin:1976:DRA

[5] Chyuan Shiun Lin, Diane C. P. Smith, and John Miles Smith. The design of a rotating associative memory for relational database applications. ACM Transactions on Database Systems, 1 (1):53–65, 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1976-1-1/p53-lin/p53-lin.pdf; http:

//www.acm.org/pubs/citations/journals/
tods/1976-1-1/p53-lin/.

Mahmoud:1976:OAR

[6] Samy Mahmoud and J. S. Riordon. Optimal allocation of resources in distributed information networks. ACM Transactions on Database Systems, 1(1):66–78, March 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1976-1-1/p66-mahmoud/p66-mahmoud.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-
1/p66-mahmoud/.

Stemple:1976:DMF

[7] David W. Stemple. A database management facility for automatic generation of database managers. ACM Transactions on Database Systems, 1(1):79–94, March 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1976-1-1/p79-stemple/p79-stemple.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-

1/p79-stemple/. Also published in [1088, p. 252].

Astrahan:1976:SRR

[8] M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, K. P. Eswaran, J. N. Gray, P. P. Griffiths, W. F. King, R. A. Lorie, P. R. McJones, J. W. Mehl, G. R. Putzolu, I. L. Traiger, B. W. Wade, and V. Watson. System R: a relational approach to database management. ACM Transactions on Database Systems, 1(2):97– 137, June 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1976-1-2/p97-astrahan/p97-
astrahan.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1976-1-2/p97-astrahan/. Also pub-

lished in/as: IBM, San Jose, Research Report. No. RJ-1738, Feb. 1976. Reprinted in [1093].

Navathe:1976:RLD

[9] Shamkant B. Navathe and James P. Fry. Restructuring for large data bases: Three levels of abstraction. ACM Transactions on Database Systems, 1 (2):138–158, March 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1976-1-2/p138-navathe/p138-
navathe.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1976-
1-2/p138-navathe/. Also published in

[1088, p. 174].

Yao:1976:DDR

[10] S. B. Yao, K. S. Das, and T. J. Teorey. A dynamic database reorganization algorithm. ACM Transactions on Database Systems, 1(2):159– 174, June 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1976-1-2/p159-yao/p159-yao.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-

2/p159-yao/. Also published in/as: Purdue Un., TR-168, Nov. 1975.

Burkhard:1976:HTA

[11] Walter A. Burkhard. Hashing and trie algorithms for partial-match retrieval. ACM Transactions on Database Systems, 1(2):175–187, June 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1976-1-2/p175-burkhard/p175-burkhard.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-
2/p175-burkhard/. Also published

in/as: UCSD, Appl. Physics and Inf. Sc, CS TR.2, Jun. 1975.

Stonebraker:1976:DII

[12] Michael Stonebraker, Eugene Wong, Peter Kreps, and Gerald Held. The design and implementation of INGRES. ACM Transactions on Database Systems, 1(3):189–222, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1976-1-3/p189-stonebraker/
p189-stonebraker.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1976-1-3/p189-stonebraker/.

Reprinted in [1093]. Also published in/as: UCB, Elec. Res. Lab, Memo No. ERL-M577, Jan. 1976.

Wong:1976:DSQ

[13] Eugene Wong and Karel Youssefi. Decomposition — A strategy for query processing. ACM Transactions on Database Systems, 1(3):223–241, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/

                                        1976-1-3/p223-wong/p223-wong.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1976-1-3/p223-wong/

. Also published in/as: UCB, Elec. Res. Lab, Memo No. ERL-574, Jan. 1976.

Griffiths:1976:AMR

[14] Patricia P. Griffiths and Bradford W. Wade. An authorization mechanism for a relational database system. ACM Transactions on Database Systems, 1(3):242–255, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1976-1-3/p242-griffiths/p242-griffiths.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1976-1-
                                        3/p242-griffiths/.

Severance:1976:DFT

[15] Dennis G. Severance and Guy M. Lohman. Differential files: Their application to the maintenance of large databases. ACM Transactions on Database Systems, 1(3):256–267, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1976-1-3/p256-severance/p256-severance.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1976-1-
                                        3/p256-severance/.

Shneiderman:1976:BSS

[16] Ben Shneiderman and Victor Goodman. Batched searching of sequential and tree structured files. ACM Transactions on Database Systems, 1 (3):268–275, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1976-1-3/p268-shneiderman/
p268-shneiderman.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1976-1-3/p268-shneiderman/.

See comments in [251]. Also published in/as: Indiana Un., CSD Tech. Ref. 0132.

Bernstein:1976:STN

[17] Philip A. Bernstein. Synthesizing third normal form relations from functional dependencies. ACM Transactions on Database Systems, 1(4):277–298, December 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1976-1-4/p277-bernstein/p277-bernstein.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-
4/p277-bernstein/.

Liu:1976:APS

[18] Jane W. S. Liu. Algorithms for parsing search queries in systems with inverted file organization. ACM Transactions on Database Systems, 1(4): 299–316, 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1976-1-4/p299-liu/p299-liu.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1976-1-
4/p299-liu/.

Sherman:1976:PDM

[19] Stephen W. Sherman and Richard S. Brice. Performance of a database manager in a virtual memory system. ACM Transactions on Database Systems, 1(4):317–343, December 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1976-1-4/p317-sherman/p317-sherman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1976-1-
                                        4/p317-sherman/.

Donovan:1976:DSA

[20] John J. Donovan. Database system approach to management decision support. ACM Transactions on Database Systems, 1(4):344–369, December 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1976-1-4/p344-donovan/p344-
                                        donovan.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1976-
                                        1-4/p344-donovan/.

McGee:1976:UCD

[21] William C. McGee. On user criteria for data model evaluation. ACM Transactions on Database Systems, 1 (4):370–387, 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1976-1-4/p370-mcgee/p370-mcgee.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1976-1-
                                        4/p370-mcgee/.

Kam:1977:MSD

[22] John B. Kam and Jeffrey D. Ullman. A model of statistical databases and their security. ACM Transactions on Database Systems, 2(1):1– 10, March 1977. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1977-2-1/p1-kam/p1-kam.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1977-2-1/p1-kam/.

Bayer:1977:PBT

[23] Rudolf Bayer and Karl Unterauer. Prefix B-trees. ACM Transactions on Database Systems, 2(1):11–26, March 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-1/p11-bayer/p11-bayer.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1977-2-1/p11-bayer/

. Also published in/as: IBM Yorktwon, Technical Report RJ1796, Jun. 1976.

Schkolnick:1977:CAH

[24] Mario Schkolnick. A clustering algorithm for hierarchical structures. ACM Transactions on Database Systems, 2(1):27–44, May 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yao:1977:ABM

[25] S. B. Yao. An attribute based model for database access cost analysis. ACM Transactions on Database Systems, 2(1):45–67, March 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1977-2-1/p45-yao/p45-yao.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1977-2-1/p45-yao/. Also pub-

lished in [1].

Anderson:1977:MCS

[26] Henry D. Anderson and P. Bruce Berra. Minimum cost selection of secondary indexes for formatted files. ACM Transactions on Database Systems, 2(1):68–90, March 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1977-2-1/p68-anderson/p68-anderson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
1/p68-anderson/.

Lorie:1977:PIL

[27] Raymond A. Lorie. Physical integrity in a large segmented database. ACM Transactions on Database Systems, 2 (1):91–104, March 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1977-2-1/p91-lorie/p91-lorie.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
1/p91-lorie/.

Smith:1977:DAA

[28] John Miles Smith and Diane C. P. Smith. Database abstractions: Aggregation and generalization. ACM Transactions on Database Systems, 2(2):105– 133, June 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-2/p105-smith/p105-smith.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
2/p105-smith/.

Shu:1977:EDE

[29] N. C. Shu, B. C. Housel, R. W. Taylor, S. P. Ghosh, and V. Y. Lum. EXPRESS: a data EXtraction, Processing, and Restructuring System. ACM Transactions on Database Systems, 2(2):134–174, June 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1977-2-2/p134-shu/p134-shu.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
2/p134-shu/.

Ozkarahan:1977:PER

[30] E. A. Ozkarahan, S. A. Schuster, and K. C. Sevcik. Performance evaluation of a relational associative processor. ACM Transactions on Database Systems, 2(2):175– 195, June 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-2/p175-ozkarahan/p175-ozkarahan.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
2/p175-ozkarahan/.

Brice:1977:EPD

[31] Richard S. Brice and Stephen W. Sherman. An extension on the performance of a database manager in a virtual memory system using partially locked virtual buffers. ACM Transactions on Database Systems, 2(2):196– 207, June 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-2/p196-brice/p196-brice.

                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1977-2-
                                        2/p196-brice/.

Lohman:1977:OPB

[32] Guy M. Lohman and John A. Muckstadt. Optimal policy for batch operations: Backup, checkpointing, reorganization, and updating. ACM Transactions on Database Systems, 2(3):209–222, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1977-2-3/p209-lohman/p209-lohman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1977-2-
                                        3/p209-lohman/.

Wong:1977:IHT

[33] Kai C. Wong and Murray Edelberg. Interval hierarchies and their application to predicate files. ACM Transactions on Database Systems, 2(3):223–232, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1977-2-3/p223-wong/p223-wong.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1977-2-3/p223-wong/

.

Ries:1977:ELG

[34] Daniel R. Ries and Michael Stonebraker. Effects of locking granularity in a database management system. ACM Transactions on Database Systems, 2(3):233–246, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-3/p233-ries/p233-ries.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1977-2-3/p233-ries/

.

Schmidt:1977:SHL

[35] Joachim W. Schmidt. Some high level language constructs for data of type relation. ACM Transactions on Database Systems, 2(3):247–261, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1977-2-3/p247-schmidt/p247-
schmidt.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1977-
2-3/p247-schmidt/.

Fagin:1977:MVD

[36] Ronald Fagin. Multi-valued dependencies and a new normal form for relational databases. ACM Transactions on Database Systems, 2(3):262–278, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-3/p262-fagin/p262-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
3/p262-fagin/.

March:1977:DER

[37] Salvatore T. March and Dennis G. Severance. The determination of efficient record segmentations and blocking factors for shared data files. ACM Transactions on Database Systems, 2(3):279–296, September 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1977-2-3/p279-march/p279-march.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
3/p279-march/.

Ozkarahan:1977:AAF

[38] E. A. Ozkarahan and K. C. Sevcik. Analysis of architectural features for enhancing the performance of a database machine. ACM Transactions on Database Systems, 2(4):297–316, December 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-4/p297-ozkarahan/p297-ozkarahan.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
4/p297-ozkarahan/.

Rissanen:1977:ICR

[39] Jorma Rissanen. Independent components of relations. ACM Transactions on Database Systems, 2(4):317–325, December 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-4/p317-rissanen/p317-rissanen.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
4/p317-rissanen/.

Bonczek:1977:TGB

[40] Robert H. Bonczek, James I. Cash, and Andrew B. Whinston. A transformational grammar-based query processor for access control in a planning system. ACM Transactions on Database Systems, 2(4):326–338, December 1977. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1977-2-4/p326-bonczek/p326-
bonczek.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1977-
2-4/p326-bonczek/.

Lang:1977:DBP

[41] Tomás Lang, Christopher Wood, and Eduardo B. Fernández. Database buffer paging in virtual storage systems. ACM Transactions on Database Systems, 2(4):339–351, December 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1977-2-4/p339-lang/p339-lang.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1977-2-4/p339-lang/

.

Thomas:1977:VAP

[42] D. A. Thomas, B. Pagurek, and R. J. Buhr. Validation algorithms for pointer values in DBTG databases. ACM Transactions on Database Systems, 2(4):352–369, December 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1977-2-4/p352-thomas/p352-
thomas.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1977-
2-4/p352-thomas/.

Claybrook:1977:FDM

[43] Billy G. Claybrook. A facility for defining and manipulating generalized data structures. ACM Transactions on Database Systems, 2(4):370–406, December 1977. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1977-2-4/p370-claybrook/p370-claybrook.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1977-2-
4/p370-claybrook/.

Minker:1978:SSS

[44] Jack Minker. Search strategy and selection function for an inferential relational system. ACM Transactions on Database Systems, 3(1):1– 31, March 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1978-3-1/p1-minker/p1-minker.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1978-3-1/p1-minker/

.

Tuel:1978:ORP

[45] William G. Tuel, Jr. Optimum reorganization points for linearly growing files. ACM Transactions on Database Systems, 3(1):32–40, March 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1978-3-1/p32-tuel/p32-tuel.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
1/p32-tuel/.

Yu:1978:END

[46] C. T. Yu, W. S. Luk, and M. K. Siu. On the estimation of the number of desired records with respect to a given query. ACM Transactions on Database Systems, 3(1):41– 56, March 1978. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1978-3-1/p41-yu/p41-yu.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1978-3-1/p41-yu/.

Su:1978:CCS

[47] Stanley Y. W. Su and Ahmed Emam. CASDAL: CASSM’s DAta Language. ACM Transactions on Database Systems, 3(1):57–91, March 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/tods/1978-
3-1/p57-su/.

Chin:1978:SSD

[48] Francis Y. Chin. Security in statistical databases for queries with small counts. ACM Transactions on Database Systems, 3(1):92–104, March 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1978-3-1/p92-chin/p92-chin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
1/p92-chin/.

Hendrix:1978:DNL

[49] Gary G. Hendrix, Earl D. Sacerdoti, Daniel Sagalowicz, and Jonathan Slocum. Developing a natural language interface to complex data. ACM Transactions on Database Systems, 3(2):105– 147, June 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1978-3-2/p105-hendrix/p105-

                                        hendrix.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1978-
                                        3-2/p105-hendrix/.

Langdon:1978:NAP

[50] Glen G. Langdon, Jr. A note on associative processors for data management. ACM Transactions on Database Systems, 3(2):148–158, June 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1978-3-2/p148-langdon/p148-
                                        langdon.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1978-
                                        3-2/p148-langdon/.

Kluge:1978:DFM

[51] Werner E. Kluge. Data file management in shift-register memories. ACM Transactions on Database Systems, 3 (2):159–177, June 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1978-3-2/p159-kluge/p159-kluge.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1978-3-
                                        2/p159-kluge/.

Rosenkrantz:1978:SLC

[52] David J. Rosenkrantz, Richard E. Stearns, and Philip M. Lewis, II. System level concurrency control for distributed database systems. ACM Transactions on Database Systems, 3 (2):178–198, June 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1978-3-2/p178-rosenkrantz/
                                        p178-rosenkrantz.pdf;   http://

www.acm.org/pubs/citations/journals/
tods/1978-3-2/p178-rosenkrantz/.

Delobel:1978:NHD

[53] Claude Delobel. Normalization and hierarchical dependencies in the relational data model. ACM Transactions on Database Systems, 3(3):201–222, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1978-3-3/p201-delobel/p201-
delobel.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1978-
3-3/p201-delobel/.

Smith:1978:SPD

[54] Alan Jay Smith. Sequentiality and prefetching in database systems. ACM Transactions on Database Systems, 3(3):223–247, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1978-3-3/p223-smith/p223-smith.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
3/p223-smith/.

Schlageter:1978:PSD

[55] Gunter Schlageter. Process synchronization in database systems. ACM Transactions on Database Systems, 3 (3):248–271, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1978-3-3/p248-schlageter/p248-
schlageter.pdf;   http://www.acm.
org/pubs/citations/journals/tods/

                                        1978-3-3/p248-schlageter/. See er-

rata report in [91].

Hollaar:1978:SMP

[56] Lee A. Hollaar. Specialized merge processor networks for combining sorted lists. ACM Transactions on Database Systems, 3(3):272–284, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1978-3-3/p272-hollaar/p272-
                                        hollaar.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1978-
                                        3-3/p272-hollaar/.

Haerder:1978:IGA

[57] Theo Haerder. Implementing a generalized access path structure for a relational database system. ACM Transactions on Database Systems, 3(3): 285–298, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1978-3-3/p285-haerder/p285-
                                        haerder.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1978-
                                        3-3/p285-haerder/.

Yu:1978:PP

[58] C. T. Yu, M. K. Siu, and K. Lam. On a partitioning problem. ACM Transactions on Database Systems, 3(3): 299–309, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1978-3-3/p299-yu/p299-yu.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tods/1978-3-3/p299-yu/.

Fagin:1978:AM

[59] Ronald Fagin. On an authorization mechanism. ACM Transactions on Database Systems, 3(3):310–319, September 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1978-3-3/p310-fagin/p310-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
3/p310-fagin/.

Salton:1978:GSC

[60] G. Salton and A. Wong. Generation and search of clustered files. ACM Transactions on Database Systems, 3(4):321–346, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1978-3-4/p321-salton/p321-salton.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
4/p321-salton/.

Banerjee:1978:CCD

[61] Jayanta Banerjee, Richard I. Baum, and David K. Hsiao. Concepts and capabilities of a database computer. ACM Transactions on Database Systems, 3(4):347–384, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1978-3-4/p347-banerjee/p347-banerjee.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1978-3-
4/p347-banerjee/.

Bradley:1978:EOC

[62] J. Bradley. An extended ownercoupled set data model and predicate calculus for database management. ACM Transactions on Database Systems, 3(4):385–416, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1978-3-4/p385-bradley/p385-
                                        bradley.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1978-
                                        3-4/p385-bradley/.

Shneiderman:1978:IHF

[63] Ben Shneiderman. Improving the human factors aspect of database interactions. ACM Transactions on Database Systems, 3(4):417–439, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1978-3-4/p417-shneiderman/
                                        p417-shneiderman.pdf;   http://
                                        www.acm.org/pubs/citations/journals/
                                        tods/1978-3-4/p417-shneiderman/.

Comer:1978:DOI

[64] Douglas Comer. The difficulty of optimum index selection. ACM Transactions on Database Systems, 3(4): 440–445, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1978-3-4/p440-comer/p440-comer.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1978-3-
                                        4/p440-comer/.

Babb:1979:IRD

[65] E. Babb. Implementing a relational database by means of specialized hardware. ACM Transactions on Database Systems, 4(1):1– 29, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-1/p1-babb/p1-babb.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1979-4-1/p1-babb/.

Beeri:1979:CPR

[66] Catriel Beeri and Philip A. Bernstein. Computational problems related to the design of normal form relational schemas. ACM Transactions on Database Systems, 4(1):30– 59, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-1/p30-beeri/p30-beeri.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1979-4-1/p30-beeri/

. Also published in/as: errata in ACM Transactions on Database Systems, Vol. 4 No. 3, Sep. 1979, pp. 396.

Lockemann:1979:DAD

[67] Peter C. Lockemann, Heinrich C. Mayr, Wolfgang H. Weil, and Wolfgang H. Wohlleber. Data abstractions for database systems. ACM Transactions on Database Systems, 4(1):60– 75, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-1/p60-lockemann/p60-
lockemann.pdf;   http://www.acm.

org/pubs/citations/journals/tods/
1979-4-1/p60-lockemann/.

Denning:1979:TTS

[68] Dorothy E. Denning, Peter J. Denning, and Mayer D. Schwartz. The tracker: a threat to statistical database security. ACM Transactions on Database Systems, 4(1):76–96, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-1/p76-denning/p76-denning.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
1/p76-denning/.

Dobkin:1979:SDP

[69] David Dobkin, Anita K. Jones, and Richard J. Lipton. Secure databases: Protection against user influence. ACM Transactions on Database Systems, 4(1):97–106, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-1/p97-dobkin/p97-dobkin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
1/p97-dobkin/.

Kent:1979:LRB

[70] William Kent. Limitations of recordbased information models. ACM Transactions on Database Systems, 4 (1):107–131, March 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1979-4-1/p107-kent/p107-kent.
pdf;   http://www.acm.org/pubs/

citations/journals/tods/1979-4-
1/p107-kent/.

Yao:1979:OQE

[71] S. Bing Yao. Optimization of query evaluation algorithms. ACM Transactions on Database Systems, 4(2):133– 155, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-2/p133-yao/p133-yao.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
2/p133-yao/.

Schwartz:1979:LQS

[72] M. D. Schwartz, D. E. Denning, and P. J. Denning. Linear queries in statistical databases. ACM Transactions on Database Systems, 4(2):156– 167, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-2/p156-schwartz/p156-schwartz.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
2/p156-schwartz/.

Aho:1979:OPM

[73] Alfred V. Aho and Jeffrey D. Ullman. Optimal partial-match retrieval when fields are independently specified. ACM Transactions on Database Systems, 4(2):168– 179, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-2/p168-aho/p168-aho.
pdf;   http://www.acm.org/pubs/

citations/journals/tods/1979-4-
2/p168-aho/.

Thomas:1979:MCA

[74] Robert H. Thomas. A majority consensus approach to concurrency control for multiple copy databases. ACM Transactions on Database Systems, 4(2):180– 209, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-2/p180-thomas/p180-
thomas.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-2/p180-thomas/.

Ries:1979:LGR

[75] Daniel R. Ries and Michael R. Stonebraker. Locking granularity revisited. ACM Transactions on Database Systems, 4(2):210–227, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-2/p210-ries/p210-ries.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1979-4-2/p210-ries/

.

Burkhard:1979:PMH

[76] Walter A. Burkhard. Partialmatch hash coding: Benefits of redundancy. ACM Transactions on Database Systems, 4(2):228–239, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-2/p228-burkhard/p228-burkhard.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
2/p228-burkhard/.

Raghavan:1979:EDR

[77] Vijay V. Raghavan and C. T. Yu. Experiments on the determination of the relationships between terms. ACM Transactions on Database Systems, 4(2):240–260, June 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-2/p240-raghavan/p240-raghavan.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
2/p240-raghavan/.

Lipski:1979:SIC

[78] Witold Lipski, Jr. On semantic issues connected with incomplete information databases. ACM Transactions on Database Systems, 4(3):262–296, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-3/p262-lipski/p262-
lipski.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-3/p262-lipski/.

Aho:1979:TJR

[79] A. V. Aho, C. Beeri, and J. D. Ullman. The theory of joins in relational databases. ACM Transactions on Database Systems, 4(3):297–314, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-3/p297-aho/p297-aho.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-

3/p297-aho/. See corrigendum [188].

Fagin:1979:EHF

[80] Ronald Fagin, Jürg Nievergelt, Nicholas Pippenger, and H. Raymond Strong. Extendible hashing — a fast access method for dynamic files. ACM Transactions on Database Systems, 4(3): 315–344, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1979-4-3/p315-fagin/p315-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-

3/p315-fagin/. Also published in/as: IBM Research Report RJ2305, Jul. 1978. See [1087].

Lam:1979:PSH

[81] Chat Yu Lam and Stuart E. Madnick. Properties of storage hierarchy systems with multiple page sizes and redundant data. ACM Transactions on Database Systems, 4(3):345–367, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1979-4-3/p345-lam/.

Buneman:1979:EMR

[82] O. Peter Buneman and Eric K. Clemons. Efficiently monitoring relational databases. ACM Transactions on Database Systems, 4(3):368–382, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-3/p368-buneman/p368-
buneman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-3/p368-buneman/. Also published

in/as: Working paper, 76-10-08, Dep. Decision Sciences, The Wharton School, Un. Penn, PA, Jun. 1977.

Comer:1979:HTI

[83] Douglas Comer. Heuristics for trie index minimization. ACM Transactions on Database Systems, 4(3):383–395, September 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-3/p383-comer/p383-comer.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
3/p383-comer/.

Codd:1979:EDR

[84] E. F. Codd. Extending the database relational model to capture more meaning. ACM Transactions on Database Systems, 4(4):397–434, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-4/p397-codd/p397-codd.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1979-4-4/p397-codd/

. Reprinted in [1093, pp. 457–475].

Aho:1979:EOC

[85] A. V. Aho, Y. Sagiv, and J. D. Ullman. Efficient optimization of a class of relational expressions. ACM Transactions on Database Systems, 4(4):435–454, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-4/p435-aho/p435-aho.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
4/p435-aho/.

Maier:1979:TID

[86] David Maier, Alberto O. Mendelzon, and Yehoshua Sagiv. Testing implications of data dependencies. ACM Transactions on Database Systems, 4(4):455–469, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1979-4-4/p455-maier/p455-maier.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
4/p455-maier/.

Housel:1979:PTI

[87] Barron C. Housel. Pipelining: a technique for implementing data restructurers. ACM Transactions on Database Systems, 4(4):470–492, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-4/p470-housel/p470-
housel.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-4/p470-housel/.

Shopiro:1979:TPL

[88] Jonathan E. Shopiro. Theseus — A programming language for relational databases. ACM Transactions on Database Systems, 4(4):493–517, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-4/p493-shopiro/p493-
shopiro.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-4/p493-shopiro/.

Yamamoto:1979:DBM

[89] Sumiyasu Yamamoto, Shinsei Tazawa, Kazuhiko Ushio, and Hideto Ikeda. Design of a balanced multiple-valued file-organization scheme with the least redundancy. ACM Transactions on Database Systems, 4(4):518–530, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1979-4-4/p518-yamamoto/p518-yamamoto.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1979-4-
4/p518-yamamoto/.

Batory:1979:STF

[90] Don S. Batory. On searching transposed files. ACM Transactions on Database Systems, 4(4):531–544, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1979-4-4/p531-batory/p531-
batory.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1979-
4-4/p531-batory/.

Bernstein:1979:CPS

[91] Philip A. Bernstein, Marco A. Casanova, and Nathan Goodman. Comments on “Process Synchronization in Database Systems”. ACM Transactions on Database Systems, 4 (4):545–546, December 1979. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1979-4-4/p545-bernstein/p545-
bernstein.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1979-4-4/p545-bernstein/. See [55].

Rothnie:1980:ISD

[92] James B. Rothnie, Jr., Philip A. Bernstein, S. Fox, N. Goodman, M. Hammer, T. A. Landers, C. Reeve, David W. Shipman, and E. Wong. Introduction to a system for distributed databases (SDD-1). ACM Transactions on Database Systems, 5(1):1– 17, March 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1980-5-1/p1-rothnie/p1-rothnie.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1980-5-
                                        1/p1-rothnie/.

Bernstein:1980:CCS

[93] Philip A. Bernstein, David W. Shipman, and James B. Rothnie, Jr. Concurrency control in a system for distributed databases (SDD-1). ACM Transactions on Database Systems, 5(1):18–51, March 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1980-5-1/p18-bernstein/p18-bernstein.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1980-5-
                                        1/p18-bernstein/.

Bernstein:1980:CCC

[94] Philip A. Bernstein and David W. Shipman. Correctness of concurrency control mechanisms in a system for distributed databases (SDD-1). ACM Transactions on Database Systems, 5(1):52–68, March 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-1/p52-bernstein/p52-bernstein.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
1/p52-bernstein/.

Gopalakrishna:1980:PEA

[95] V. Gopalakrishna and C. E. Veni Madhavan. Performance evaluation of attribute-based tree organization. ACM Transactions on Database Systems, 5(1):69–87, March 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1980-5-1/p69-gopalakrishna/p69-
gopalakrishna.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1980-5-1/p69-gopalakrishna/

.

Denning:1980:FPF

[96] Dorothy E. Denning and Jan Schlörer. Fast procedure for finding a tracker in a statistical database. ACM Transactions on Database Systems, 5(1):88– 102, March 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-1/p88-denning/p88-denning.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
1/p88-denning/.

Menasce:1980:LPR

[97] Daniel A. Menascé, Gerald J. Popek, and Richard R. Muntz. A locking protocol for resource coordination in distributed databases. ACM Transactions on Database Systems, 5(2):103– 138, June 1980. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1980-5-2/p103-menasce/p103-
                                        menasce.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1980-
                                        5-2/p103-menasce/.

Bayer:1980:PRD

[98] R. Bayer, H. Heller, and A. Reiser. Parallelism and recovery in database systems. ACM Transactions on Database Systems, 5(2):139–156, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1980-5-2/p139-bayer/p139-bayer.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1980-5-
                                        2/p139-bayer/.

Navathe:1980:SAD

[99] Shamkant B. Navathe. Schema analysis for database restructuring. ACM Transactions on Database Systems, 5 (2):157–184, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1980-5-2/p157-navathe/p157-
                                        navathe.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1980-
                                        5-2/p157-navathe/. Also published

in/as: Proceedings of the Third Conference on Very Large Databases, Morgan Kaufman pubs. (Los Altos CA), 1977.

Mylopoulos:1980:LFD

[100] John Mylopoulos, Philip A. Bernstein, and Harry K. T. Wong. A language facility for designing databaseintensive applications. ACM Transactions on Database Systems, 5(2):185– 207, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-2/p185-mylopoulos/p185-mylopoulos.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
2/p185-mylopoulos/.

Lozinskii:1980:CRR

[101] Eliezer L. Lozinskii. Construction of relations in relational databases. ACM Transactions on Database Systems, 5 (2):208–224, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1980-5-2/p208-lozinskii/p208-
lozinskii.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1980-5-2/p208-lozinskii/.

Stonebraker:1980:RDS

[102] Michael Stonebraker. Retrospection on a database system. ACM Transactions on Database Systems, 5(2):225– 240, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1980-5-2/p225-stonebraker/
p225-stonebraker.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1980-5-2/p225-stonebraker/.

Reprinted in [1093].

Beeri:1980:MPF

[103] Catriel Beeri. On the membership problem for functional and multivalued dependencies in relational databases. ACM Transactions on Database Systems, 5(3):241–259, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1980-5-3/p241-beeri/p241-beeri.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1980-5-
                                        3/p241-beeri/.

Klug:1980:CCR

[104] A. Klug. Calculating constraints on relational expressions. ACM Transactions on Database Systems, 5(3): 260–290, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1980-5-3/p260-klug/p260-klug.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1980-5-
                                        3/p260-klug/.

Denning:1980:SSD

[105] Dorothy E. Denning. Secure statistical databases with random sample queries. ACM Transactions on Database Systems, 5(3):291–315, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1980-5-3/p291-denning/p291-
                                        denning.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1980-
                                        5-3/p291-denning/.

Beck:1980:SMS

[106] Leland L. Beck. A security mechanism for statistical database. ACM Transactions on Database Systems, 5 (3):316–338, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1980-5-3/p316-beck/p316-beck.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
3/p316-beck/.

Lee:1980:QTF

[107] D. T. Lee and C. K. Wong. Quintary trees: a file structure for multidimensional database systems. ACM Transactions on Database Systems, 5(3):339–353, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1980-5-3/p339-lee/p339-lee.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1980-5-3/p339-lee/

.

Kung:1980:CMB

[108] H. T. Kung and Philip L. Lehman. Concurrent manipulation of binary search trees. ACM Transactions on Database Systems, 5(3):354–382, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-3/p354-kung/p354-kung.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1980-5-3/p354-kung/

.

Denning:1980:CLQ

[109] D. E. Denning. Corrigenda on linear queries in statistical databases. ACM Transactions on Database Systems, 5 (3):383, September 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hsiao:1980:TFT

[110] David K. Hsiao. TODS — the first three years (1976&ndash1978). ACM Transactions on Database Systems, 5 (4):385–403, December 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1980-5-4/p385-hsiao/p385-hsiao.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
4/p385-hsiao/.

Armstrong:1980:DFD

[111] W. W. Armstrong and C. Delobel. Decompositions and functional dependencies in relations. ACM Transactions on Database Systems, 5(4):404–430, December 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-4/p404-armstrong/p404-armstrong.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
4/p404-armstrong/.

Hammer:1980:RMS

[112] Michael Hammer and David Shipman. Reliability mechanisms for SDD-1: a system for distributed databases. ACM Transactions on Database Systems, 5(4):431–466, December 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1980-5-4/p431-hammer/p431-
hammer.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1980-
5-4/p431-hammer/.

Schloer:1980:SSD

[113] Jan Schlöer. Security of statistical databases: multidimensional transformation. ACM Transactions on Database Systems, 5(4):467–492, December 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1980-5-4/p467-schler/.

Herot:1980:SMD

[114] Christopher F. Herot. Spatial management of data. ACM Transactions on Database Systems, 5(4):493–513, December 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1980-5-4/p493-herot/p493-herot.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1980-5-
4/p493-herot/.

Zaniolo:1981:DRD

[115] Carlo Zaniolo and Michel A. Melkanoff. On the design of relational database schemata. ACM Transactions on Database Systems, 6(1):1– 47, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-1/p1-zaniolo/p1-zaniolo.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
1/p1-zaniolo/.

Lien:1981:HSR

[116] Y. Edmund Lien. Hierarchical schemata for relational databases. ACM Transactions on Database Systems, 6(1):48–69, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1981-6-1/p48-lien/p48-lien.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
1/p48-lien/.

Chamberlin:1981:SRT

[117] D. D. Chamberlin, M. M. Astrahan, W. F. King, R. A. Lorie, J. W. Mehl, T. G. Price, M. Schkolnick, P. Griffiths Selinger, D. R. Slutz, B. W. Wade, and R. A. Yost. Support for repetitive transactions and ad hoc queries in System R. ACM Transactions on Database Systems, 6(1):70– 94, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-1/p70-chamberlin/p70-chamberlin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
1/p70-chamberlin/.  Also pub-

lished in/as: IBM Research Report RJ2551(33151), May. 1979.

Schlorer:1981:SSD

[118] Jan Schlörer. Security of statistical databases: Multidimensional transformation. ACM Transactions on Database Systems, 6(1):95–112, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-1/p95-schlorer/p95-
schlorer.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1981-6-1/p95-schlorer/.

Chin:1981:SDD

[119] Francis Y. Chin and Gultekin Özsoyoğlu. Statistical database design. ACM Transactions on Database Systems, 6 (1):113–139, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1981-6-1/p113-chin/p113-chin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
1/p113-chin/.

Shipman:1981:FDM

[120] David W. Shipman. The functional data model and the data language DAPLEX. ACM Transactions on Database Systems, 6(1):140– 173, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-1/p140-shipman/p140-
shipman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1981-
6-1/p140-shipman/. Reprinted in

[1093].

Rosenberg:1981:TSO

[121] Arnold L. Rosenberg and Lawrence Snyder. Time- and space-optimality in B-Trees. ACM Transactions on Database Systems, 6(1):174–193, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-1/p174-rosenberg/p174-rosenberg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
1/p174-rosenberg/.

Scholl:1981:NFO

[122] Michel Scholl. New file organizations based on dynamic hashing. ACM Transactions on Database Systems, 6 (1):194–211, March 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1981-6-1/p194-scholl/p194-
                                        scholl.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1981-
                                        6-1/p194-scholl/.

Kung:1981:OMC

[123] H. T. Kung and John T. Robinson. On optimistic methods for concurrency control. ACM Transactions on Database Systems, 6(2):213– 226, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1981-6-2/p213-kung/p213-kung.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1981-6-2/p213-kung/

. Reprinted in [1093].

Boral:1981:PAS

[124] Haran Boral and David J. De-Witt. Processor allocation strategies for multiprocessor database machines. ACM Transactions on Database Systems, 6(2):227–254, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1981-6-2/p227-boral/p227-boral.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1981-6-
                                        2/p227-boral/.

Su:1981:TDT

[125] Stanley Y. W. Su, Herman Lam, and Der Her Lo. Transformation of data traversals and operations in application programs to account for semantic changes of databases. ACM Transactions on Database Systems, 6(2):255– 294, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-2/p255-su/p255-su.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1981-6-2/p255-su/.

Clemons:1981:DES

[126] Eric K. Clemons. Design of an external schema facility to define and process recursive structures. ACM Transactions on Database Systems, 6(2):295– 311, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-2/p295-clemons/p295-
clemons.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1981-
6-2/p295-clemons/.

Davida:1981:DES

[127] George I. Davida, David L. Wells, and John B. Kam. A database encryption system with subkeys. ACM Transactions on Database Systems, 6(2):312– 328, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-2/p312-davida/p312-
davida.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1981-
6-2/p312-davida/.

Ling:1981:ITN

[128] Tok Wang Ling, Frank W. Tompa, and Tiko Kameda. An improved third normal form for relational databases. ACM Transactions on Database Systems, 6(2):329–346, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1981-6-2/p329-ling/p329-ling.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1981-6-2/p329-ling/

.

McLean:1981:CSC

[129] Gordon McLean, Jr. Comments on SDD-1 concurrency control mechanisms. ACM Transactions on Database Systems, 6(2):347–350, June 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1981-6-2/p347-mclean/p347-mclean.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1981-6-
                                        2/p347-mclean/.

Hammer:1981:DDS

[130] Michael Hammer and Dennis Mc Leod. Database description with SDM: a semantic database model. ACM Transactions on Database Systems, 6(3): 351–386, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1981-6-3/p351-hammer/p351-
                                        hammer.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1981-
                                        6-3/p351-hammer/.  Reprinted in

[1093]. Also published in [1095].

Fagin:1981:NFR

[131] Ronald Fagin. A normal form for relational databases that is based on domains and keys. ACM Transactions on Database Systems, 6(3):387–415, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-3/p387-fagin/p387-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
3/p387-fagin/.

Hong:1981:AHS

[132] Y. C. Hong and Stanley Y. W. Su. Associative hardware and software techniques for integrity control. ACM Transactions on Database Systems, 6(3):416–440, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1981-6-3/p416-hong/p416-hong.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1981-6-3/p416-hong/

.

March:1981:FMS

[133] Salvatore T. March, Dennis G. Severance, and Michael Wilens. Frame memory: a storage architecture to support rapid design and implementation of efficient databases. ACM Transactions on Database Systems, 6(3):441–463, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-3/p441-march/p441-march.
pdf;   http://www.acm.org/pubs/

citations/journals/tods/1981-6-
3/p441-march/.

vandeRiet:1981:HLP

[134] Reind P. van de Riet, Anthony I. Wasserman, Martin L. Kersten, and Wiebren de Jonge. High-level programming features for improving the efficiency of a relational database system. ACM Transactions on Database Systems, 6(3):464–485, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-3/p464-van_de_riet/
p464-van_de_riet.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1981-6-3/p464-van_de_riet/.

Also published in/as: UCSF, Lab. of Med. Inf. Science, Tech. Rpt. 44, Feb. 1980.

Culik:1981:DMT

[135] K. Culik II, Th. Ottmann, and D. Wood. Dense multiway trees. ACM Transactions on Database Systems, 6 (3):486–512, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/citations/journals/
tods/1981-6-3/p486-culic/.

Comer:1981:AHF

[136] Douglas Comer. Analysis of a heuristic for full trie minimization. ACM Transactions on Database Systems, 6(3):513–537, September 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1981-6-3/p513-comer/p513-comer.

pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
3/p513-comer/.

Kent:1981:CAU

[137] W. Kent. Consequences of assuming a universal relation. ACM Transactions on Database Systems, 6(4):539–556, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-4/p539-kent/p539-kent.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1981-6-4/p539-kent/

. See remark [203].

Bancilhon:1981:USR

[138] F. B. Bancilhon and N. Spyratos. Update semantics of relational views. ACM Transactions on Database Systems, 6(4):557–575, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/articles/
journals/tods/1981-6-4/p557-bancilhon/
p557-bancilhon.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1981-6-4/p557-bancilhon/. See

comment [303].

Baroody:1981:OOA

[139] A. James Baroody, Jr. and David J. DeWitt. An object-oriented approach to database system implementation. ACM Transactions on Database Systems, 6(4):576–601, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1981-6-4/p576-baroody/p576-
baroody.pdf;   http://www.acm.org/

pubs/citations/journals/tods/1981-
6-4/p576-baroody/.

Bernstein:1981:QPS

[140] Philip A. Bernstein, Nathan Goodman, Eugene Wong, Christopher L. Reeve, and James B. Rothnie, Jr. Query processing in a system for distributed databases (SDD-1). ACM Transactions on Database Systems, 6(4):602–625, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-4/p602-bernstein/p602-bernstein.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
4/p602-bernstein/.

Welty:1981:HFC

[141] Charles Welty and David W. Stemple. Human factors comparison of a procedural and a nonprocedural query language. ACM Transactions on Database Systems, 6(4):626–649, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1981-6-4/p626-welty/p626-welty.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
4/p626-welty/.

Lehman:1981:ELC

[142] Philip L. Lehman and S. Bing Yao. Efficient locking for concurrent operations on B-Trees. ACM Transactions on Database Systems, 6(4):650–670, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/

tods/1981-6-4/p650-lehman/p650-
lehman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1981-
6-4/p650-lehman/.

Larson:1981:AIS

[143] Per-Åke Larson. Analysis of indexsequential files with overflow chaining. ACM Transactions on Database Systems, 6(4):671–680, December 1981. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1981-6-4/p671-larson/p671-larson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1981-6-
4/p671-larson/.

Comer:1981:EKD

[144] D. Comer. Extended K-d tree database organization: a dynamic multiattribute file corresponds to leaves in the tree. ACM Transactions on Database Systems, 6(3):??, September 1981. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zaniolo:1982:DRN

[145] C. Zaniolo. Database relations with null values. ACM Transactions on Database Systems, 1(1):??, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Katz:1982:DCD

[146] R. H. Katz and E. Wong. Decompiling CODASYL DML into relational queries. ACM Transactions on Database Systems, 7(1):1– 23, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-1/p1-katz/p1-katz.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1982-7-1/p1-katz/.

Zaniolo:1982:FAD

[147] Carlo Zaniolo and Michel A. Melkanoff. A formal approach to the definition and the design of conceptual schemata for database systems. ACM Transactions on Database Systems, 7(1):24– 59, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-1/p24-zaniolo/p24-zaniolo.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
1/p24-zaniolo/.

Batory:1982:OFD

[148] D. S. Batory. Optimal file designs and reorganization points. ACM Transactions on Database Systems, 7(1):60– 81, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-1/p60-batory/p60-batory.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-

1/p60-batory/. Also published in/as: University of Toronto, TR-CSRG-110, 1980.

Du:1982:DAC

[149] H. C. Du and J. S. Sobolewski. Disk allocation for Cartesian product files on multiple-disk systems. ACM Transactions on Database Systems, 7(1):82– 101, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-1/p82-du/p82-du.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1982-7-1/p82-du/.

Dahl:1982:DSD

[150] Verónica Dahl. On database systems development through logic. ACM Transactions on Database Systems, 7 (1):102–123, March 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1982-7-1/p102-dahl/p102-dahl.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
1/p102-dahl/.

Addis:1982:RBL

[151] T. R. Addis. A relation-based language interpreter for a content addressable file store. ACM Transactions on Database Systems, 7(2):125– 163, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-2/p125-addis/p125-addis.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
2/p125-addis/.

Buneman:1982:ITD

[152] Peter Buneman, Robert E. Frankel, and Rishiyur Nikhil. An implementation technique for database query languages. ACM Transactions on Database Systems, 7(2):164– 186, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/

                                        tods/1982-7-2/p164-buneman/p164-
                                        buneman.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1982-
                                        7-2/p164-buneman/.

Obermarck:1982:DDD

[153] Ron Obermarck. Distributed deadlock detection algorithm. ACM Transactions on Database Systems, 7(2):187– 208, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1982-7-2/p187-obermarck/p187-obermarck.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1982-7-
                                        2/p187-obermarck/.

Garcia-Molina:1982:ROT

[154] Héctor García-Molina and Gio Wiederhold. Read-only transactions in a distributed database. ACM Transactions on Database Systems, 7(2):209– 234, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1982-7-2/p209-garcia-molina/p209-
                                        garcia-molina.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tods/1982-7-2/p209-garcia-molina/

.

Shneiderman:1982:AAR

[155] Ben Shneiderman and Glenn Thomas. An architecture for automatic relational database system conversion. ACM Transactions on Database Systems, 7(2):235–257, June 1982. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

                                        http://www.acm.org/pubs/articles/
                                        journals/tods/1982-7-2/p235-shneiderman/

p235-shneiderman.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1982-7-2/p235-shneiderman/.

Roussopoulos:1982:VIR

[156] Nicholas Roussopoulos. View indexing in relational databases. ACM Transactions on Database Systems, 7(2):258– 290, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-2/p258-roussopoulos/p258-
roussopoulos.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1982-7-2/p258-roussopoulos/

.

Jacobs:1982:IRL

[157] Barry E. Jacobs, Alan R. Aronson, and Anthony C. Klug. On interpretations of relational languages and solutions to the implied constraint problem. ACM Transactions on Database Systems, 7(2):291– 315, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-2/p291-jacobs/p291-
jacobs.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1982-
7-2/p291-jacobs/.

Chamberlin:1982:HFC

[158] Donald D. Chamberlin. On “human factors comparison of a procedural and a nonprocedural query language”. ACM Transactions on Database Systems, 7(2):316–317, June 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Traiger:1982:TCD

[159] Irving L. Traiger, Jim Gray, Cesare A. Galtieri, and Bruce G. Lindsay. Transactions and consistency in distributed database systems. ACM Transactions on Database Systems, 7(3):323–342, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-3/p323-traiger/p323-
traiger.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1982-
7-3/p323-traiger/.

Fagin:1982:SUR

[160] Ronald Fagin, Alberto O. Mendelzon, and Jeffrey D. Ullman. A simplified universal relation assumption and its properties. ACM Transactions on Database Systems, 7(3):343–360, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-3/p343-fagin/p343-fagin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
3/p343-fagin/.

Klug:1982:DVD

[161] Anthony Klug and Rod Price. Determining View dependencies using tableaux. ACM Transactions on Database Systems, 7(3):361–380, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-3/p361-klug/p361-klug.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1982-7-3/p361-klug/

.

Dayal:1982:CTU

[162] Umeshwar Dayal and Philip A. Bernstein. On the correct translation of update operations on relational views. ACM Transactions on Database Systems, 7(3):381–416, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-3/p381-dayal/p381-dayal.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
3/p381-dayal/.

Griffith:1982:TPR

[163] Robert L. Griffith. Three principles of representation for semantic networks. ACM Transactions on Database Systems, 7(3):417–442, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-3/p417-griffith/p417-griffith.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
3/p417-griffith/.

Kim:1982:OSL

[164] Won Kim. On optimizing an SQLlike nested query. ACM Transactions on Database Systems, 7(3):443–469, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-3/p443-kim/p443-kim.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
3/p443-kim/.

Wong:1982:SAI

[165] Eugene Wong. A statistical approach to incomplete information in database systems. ACM Transactions on Database Systems, 7(3):470–488, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1982-7-3/p470-wong/p470-wong.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1982-7-3/p470-wong/

.

Zaniolo:1982:NNF

[166] Carlo Zaniolo. A new normal form for the design of relational database schemata. ACM Transactions on Database Systems, 7(3):489–499, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1982-7-3/p489-zaniolo/p489-
                                        zaniolo.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1982-
                                        7-3/p489-zaniolo/.

Lam:1982:CSA

[167] K. Lam and C. T. Yu. A clustered search algorithm incorporating arbitrary term dependencies. ACM Transactions on Database Systems, 7(3):500–508, September 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1982-7-3/p500-lam/p500-lam.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1982-7-3/p500-lam/

.

Batory:1982:UMP

[168] D. S. Batory and C. C. Gotlieb. A unifying model of physical databases. ACM Transactions on Database Systems, 7(4):509–539, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/articles/
journals/tods/1982-7-4/p509-batory/
p509-batory.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1982-7-4/p509-batory/.

Aghili:1982:PGD

[169] Houtan Aghili and Dennis G. Severance. Practical guide to the design of differential files for recovery of online databases. ACM Transactions on Database Systems, 7(4):540–565, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-4/p540-aghili/p540-
aghili.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1982-
7-4/p540-aghili/.

Larson:1982:PAL

[170] Per-Åke Larson. Performance analysis of linear hashing with partial expansions. ACM Transactions on Database Systems, 7(4):566–587, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-4/p566-larson/p566-
larson.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1982-
7-4/p566-larson/.

Babb:1982:JNF

[171] E. Babb. Joined normal form: a storage encoding for relational databases. ACM Transactions on Database Systems, 7(4):588–614, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1982-7-4/p588-babb/p588-babb.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1982-7-4/p588-babb/

.

Heyman:1982:MMD

[172] Daniel P. Heyman. Mathematical models of database degradation. ACM Transactions on Database Systems, 7(4):615–631, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1982-7-4/p615-heyman/p615-
                                        heyman.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1982-
                                        7-4/p615-heyman/.

Korth:1982:DFU

[173] Henry F. Korth. Deadlock freedom using edge locks. ACM Transactions on Database Systems, 7(4):632–652, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1982-7-4/p632-korth/p632-korth.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1982-7-
                                        4/p632-korth/.

Goodman:1982:TQS

[174] Nathan Goodman and Oded Shmueli.

Tree queries: a simple class of relational queries. ACM Transactions on Database Systems, 7(4):653–677, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1982-7-4/p653-goodman/p653-
goodman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1982-
7-4/p653-goodman/.

Kerschberg:1982:QOS

[175] Larry Kerschberg, Peter D. Ting, and S. Bing Yao. Query optimization in star computer networks. ACM Transactions on Database Systems, 7(4):678–711, December 1982. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1982-7-4/p678-kerschberg/p678-kerschberg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1982-7-
4/p678-kerschberg/.

Maier:1983:MOS

[176] David Maier and Jeffrey D. Ullman. Maximal objects and the semantics of universal relation databases. ACM Transactions on Database Systems, 8(1):1–14, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1983-8-1/p1-maier/p1-maier.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-

1/p1-maier/. Also published in/as: SUNY, Stony Brook, CS, TR 80/016, 1980.

Haskin:1983:OCH

[177] Roger L. Haskin and Lee A. Hollaar. Operational characteristics of a hardware-based pattern matcher. ACM Transactions on Database Systems, 8(1):15–40, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1983-8-1/p15-haskin/p15-haskin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
1/p15-haskin/.

Sicherman:1983:AQR

[178] George L. Sicherman, Wiebren De Jonge, and Reind P. Van De Riet. Answering queries without revealing secrets. ACM Transactions on Database Systems, 8(1):41–59, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-1/p41-sicherman/p41-
sicherman.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1983-8-1/p41-sicherman/. Also pub-

lished in/as: reprinted in deJonge thesis, Jun. 1985.

deJonge:1983:CSD

[179] Wiebren de Jonge. Compromising statistical databases responding to queries about means. ACM Transactions on Database Systems, 8(1):60– 80, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-1/p60-de_jonge/p60-
de_jonge.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1983-8-1/p60-de_jonge/. Also pub-

lished in/as: reprinted in Jun. 1985 thesis.

Graham:1983:FD

[180] Marc H. Graham. Functions in databases. ACM Transactions on Database Systems, 8(1):81–109, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1983-8-1/p81-graham/p81-graham.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
1/p81-graham/.

Katz:1983:RCG

[181] R. H. Katz and E. Wong. Resolving conflicts in global storage design through replication. ACM Transactions on Database Systems, 8(1):110– 135, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-1/p110-katz/p110-katz.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1983-8-1/p110-katz/

.

Lomet:1983:BIE

[182] David B. Lomet. Bounded index exponential hashing. ACM Transactions on Database Systems, 8(1):136– 165, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-1/p136-lomet/p136-lomet.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
1/p136-lomet/.

Stonebraker:1983:PER

[183] Michael Stonebraker, John Woodfill, Jeff Ranstrom, Marguerite Murphy, Marc Meyer, and Eric Allman. Performance enhancements to a relational database system. ACM Transactions on Database Systems, 8(2):167– 185, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-2/p167-stonebraker/
p167-stonebraker.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1983-8-2/p167-stonebraker/.

Garcia-Molina:1983:USK

[184] Héctor García-Molina. Using semantic knowledge for transaction processing in a distributed database. ACM Transactions on Database Systems, 8(2):186– 213, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-2/p186-garcia-molina/p186-
garcia-molina.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1983-8-2/p186-garcia-molina/

.

Clifford:1983:FST

[185] James Clifford and David S. Warren. Formal semantics for time in databases. ACM Transactions on Database Systems, 8(2):214–254, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1983-8-2/p214-clifford/p214-clifford.
pdf;   http://www.acm.org/pubs/

citations/journals/tods/1983-8-
2/p214-clifford/.

Bitton:1983:DRE

[186] Dina Bitton and David J. De-Witt. Duplicate record elimination in large data files. ACM Transactions on Database Systems, 8(2):255– 265, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-2/p255-bitton/p255-
bitton.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1983-
8-2/p255-bitton/.

Sagiv:1983:CGC

[187] Yehoshua Sagiv. A characterization of globally consistent databases and their correct access paths. ACM Transactions on Database Systems, 8(2):266– 286, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-2/p266-sagiv/p266-sagiv.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
2/p266-sagiv/.

Ullman:1983:CTJ

[188] Jeffrey D. Ullman. Corrigendum: The theory of joins in relational databases. ACM Transactions on Database Systems, 8(2):287, June 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [79].

Cardenas:1983:PRA

[189] Alfonso F. Cardenas, Farid Alavian, and Algirdas Avizienis. Performance of recovery architectures in parallel associative database processors. ACM Transactions on Database Systems, 8(3):291–323, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1983-8-3/p291-cardenas/p291-cardenas.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
3/p291-cardenas/.

Bitton:1983:PAE

[190] Dina Bitton, Haran Boral, David J. DeWitt, and W. Kevin Wilkinson. Parallel algorithms for the execution of relational database operations. ACM Transactions on Database Systems, 8(3):324–353, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-3/p324-bitton/p324-
bitton.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1983-
8-3/p324-bitton/.

Eager:1983:ARD

[191] Derek L. Eager and Kenneth C. Sevcik. Achieving robustness in distributed database systems. ACM Transactions on Database Systems, 8(3):354–381, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-3/p354-eager/p354-eager.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
3/p354-eager/.

Trueblood:1983:MMM

[192] Robert P. Trueblood, H. Rex Hartson, and Johannes J. Martin. MUL-TISAFE — A modular multiprocessing approach to secure database management. ACM Transactions on Database Systems, 8(3):382–409, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-3/p382-trueblood/p382-trueblood.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
3/p382-trueblood/.

Ito:1983:HFO

[193] Tetsuro Ito and Makoto Kizawa. Hierarchical file organization and its application to similar-string matching. ACM Transactions on Database Systems, 8(3):410–433, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1983-8-3/p410-ito/p410-ito.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1983-8-3/p410-ito/

.

Kolodner:1983:IRS

[194] Janet L. Kolodner. Indexing and retrieval strategies for natural language fact retrieval. ACM Transactions on Database Systems, 8(3):434–464, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-3/p434-kolodner/p434-kolodner.
pdf;   http://www.acm.org/pubs/

citations/journals/tods/1983-8-
3/p434-kolodner/.

Bernstein:1983:MCC

[195] Philip A. Bernstein and Nathan Goodman. Multiversion concurrency control — theory and algorithms. ACM Transactions on Database Systems, 8(4):465–483, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-4/p465-bernstein/p465-bernstein.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
4/p465-bernstein/.

Lynch:1983:MAN

[196] Nancy A. Lynch. Multilevel atomicity — A new correctness criterion for database concurrency control. ACM Transactions on Database Systems, 8(4):484–502, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1983-8-4/p484-lynch/p484-lynch.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
4/p484-lynch/.

Hecht:1983:SMF

[197] Matthew S. Hecht and John D. Gabbe. Shadowed management of free disk pages with a linked list. ACM Transactions on Database Systems, 8(4):503–514, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/

1983-8-4/p503-hecht/p503-hecht.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
4/p503-hecht/.

Malhotra:1983:EIA

[198] A. Malhotra, H. M. Markowitz, and D. P. Pazel. EAS-E: An integrated approach to application development. ACM Transactions on Database Systems, 8(4):515–542, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-4/p515-malhotra/p515-malhotra.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
4/p515-malhotra/.

Moran:1983:CDO

[199] Shlomo Moran. On the complexity of designing optimal partial-match retrieval systems. ACM Transactions on Database Systems, 8(4):543–551, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-4/p543-moran/p543-moran.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1983-8-
4/p543-moran/.

Ramamohanarao:1983:PMR

[200] K. Ramamohanarao, John W. Lloyd, and James A. Thom. Partialmatch retrieval using hashing and descriptors. ACM Transactions on Database Systems, 8(4):552–576, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1983-8-4/p552-ramamohanarao/p552-
                                        ramamohanarao.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tods/1983-8-4/p552-ramamohanarao/

.

Ceri:1983:CQE

[201] S. Ceri and G. Pelagatti. Correctness of query execution strategies in distributed databases. ACM Transactions on Database Systems, 8(4):577–607, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1983-8-4/p577-ceri/p577-ceri.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1983-8-4/p577-ceri/

.

Biskup:1983:FCR

[202] Joachim Biskup. A foundation of Codd’s relational maybe operators. ACM Transactions on Database Systems, 8(4):608–636, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1983-8-4/p608-biskup/p608-
                                        biskup.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1983-
                                        8-4/p608-biskup/.

Ullman:1983:KCA

[203] Jeffrey D. Ullman. On Kent’s “Consequences of assuming a universal relation” (Technical correspondence). ACM Transactions on Database Systems, 8(4):637–643, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1983-8-4/p637-ullman/p637-
ullman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1983-
8-4/p637-ullman/. See [137].

Kent:1983:URR

[204] William Kent. The universal relation revisited (technical correspondence). ACM Transactions on Database Systems, 8(4):644–648, December 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1983-8-4/p644-kent/p644-kent.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1983-8-4/p644-kent/

.

Kaplan:1984:DPN

[205] S. Jerrold Kaplan. Designing a portable natural language database query system. ACM Transactions on Database Systems, 9(1):1–19, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1984-9-1/p1-kaplan/p1-kaplan.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1984-9-1/p1-kaplan/

.

Reiss:1984:PDS

[206] Steven P. Reiss. Practical dataswapping: The first steps. ACM Transactions on Database Systems, 9(1):20– 37, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-1/p20-reiss/p20-reiss.pdf;

                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1984-9-1/p20-reiss/

.

Nievergelt:1984:GFA

[207] J. Nievergelt, Hans Hinterberger, and Kenneth C. Sevcik. The grid file: An adaptable, symmetric multikey file structure. ACM Transactions on Database Systems, 9(1):38– 71, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1984-9-1/p38-nievergelt/p38-nievergelt.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1984-9-
                                        1/p38-nievergelt/.

Buchanan:1984:DMS

[208] Jack R. Buchanan, Richard D. Fennell, and Hanan Samet. A database management system for the Federal Courts. ACM Transactions on Database Systems, 9(1):72–88, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1984-9-1/p72-buchanan/p72-
                                        buchanan.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1984-9-1/p72-buchanan/.

Papadimitriou:1984:CCM

[209] Christos H. Papadimitriou and Paris C. Kanellakis. On concurrency control by multiple versions. ACM Transactions on Database Systems, 9(1):89– 99, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/

1984-9-1/p89-papadimitriou/p89-
papadimitriou.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1984-9-1/p89-papadimitriou/

.

Shultz:1984:RTA

[210] Roger K. Shultz and Roy J. Zingg. Response time analysis of multiprocessor computers for database support. ACM Transactions on Database Systems, 9 (1):100–132, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1984-9-1/p100-shultz/p100-
shultz.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-1/p100-shultz/.

Valduriez:1984:JSA

[211] Patrick Valduriez and Georges Gardarin. Join and semijoin algorithms for a multiprocessor database machine. ACM Transactions on Database Systems, 9(1):133–161, March 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/articles/
journals/tods/1984-9-1/p133-valduriez/
p133-valduriez.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1984-9-1/p133-valduriez/.

Christodoulakis:1984:ICA

[212] S. Christodoulakis. Implications of certain assumptions in database performance evaluation. ACM Transactions on Database Systems, 9(2):163– 186, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/

1984-9-2/p163-christodoulakis/p163-
christodoulakis.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1984-9-2/p163-christodoulakis/

.

Effelsberg:1984:LIP

[213] Wolfgang Effelsberg and Mary E. S. Loomis. Logical, internal, and physical reference behavior in CODASYL database systems. ACM Transactions on Database Systems, 9(2):187– 213, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-2/p187-effelsberg/p187-effelsberg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
2/p187-effelsberg/.

Kim:1984:PPR

[214] Won Kim, Daniel Gajski, and David J. Kuck. A parallel pipelined relational query processor. ACM Transactions on Database Systems, 9(2):214– 242, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1984-9-2/p214-kim/p214-kim.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
2/p214-kim/.

Al-Suwaiyel:1984:ATC

[215] M. Al-Suwaiyel and E. Horowitz. Algorithms for trie compaction. ACM Transactions on Database Systems, 9 (2):243–263, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/citations/journals/
tods/1984-9-2/p243-ai-suwaiyel/.

Mendelzon:1984:DST

[216] Alberto O. Mendelzon. Database states and their tableaux. ACM Transactions on Database Systems, 9(2):264– 282, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-2/p264-mendelzon/p264-mendelzon.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
2/p264-mendelzon/.

Maier:1984:FUR

[217] David Maier, Jeffrey D. Ullman, and Moshe Y. Vardi. On the foundations of the universal relation model. ACM Transactions on Database Systems, 9(2):283–308, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1984-9-2/p283-maier/p283-maier.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
2/p283-maier/.

Katsuno:1984:ECF

[218] Hirofumi Katsuno. An extension of conflict-free multi-valued dependency sets. ACM Transactions on Database Systems, 9(2):309–326, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1984-9-2/p309-katsuno/p309-
katsuno.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-2/p309-katsuno/.

Korth:1984:SUD

[219] Henry F. Korth, Gabriel M. Kuper, Joan Feigenbaum, Allen Van Gelder, and Jeffrey D. Ullman. System/U: a database system based on the universal relation assumption. ACM Transactions on Database Systems, 9(3):331–347, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1984-9-3/p331-korth/p331-korth.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
3/p331-korth/.

Wald:1984:RQI

[220] Joseph A. Wald and Paul G. Sorenson. Resolving the query inference problem using Steiner trees. ACM Transactions on Database Systems, 9(3):348–368, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-3/p348-wald/p348-wald.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1984-9-3/p348-wald/

.

Ramamohanarao:1984:RLH

[221] K. Ramamohanarao and R. Sacks-Davis. Recursive linear hashing. ACM Transactions on Database Systems, 9(3):369–391, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1984-9-3/p369-ramamohanarao/p369-
ramamohanarao.pdf;   http://www.
acm.org/pubs/citations/journals/

tods/1984-9-3/p369-ramamohanarao/

.

Cooper:1984:ATU

[222] Robert B. Cooper and Martin K. Solomon. The average time until bucket overflow. ACM Transactions on Database Systems, 9(3):392–408, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1984-9-3/p392-cooper/p392-
cooper.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-3/p392-cooper/.

March:1984:SER

[223] Salvatore T. March and Gary D. Scudder. On the selection of efficient record segmentations and backup strategies for large shared databases. ACM Transactions on Database Systems, 9(3):409–438, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-3/p409-march/p409-march.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
3/p409-march/.

Manber:1984:CCD

[224] Udi Manber and Richard E. Ladner. Concurrency control in a dynamic search structure. ACM Transactions on Database Systems, 9(3):439–455, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1984-9-3/p439-manbar/. Also

published in/as: ACM Proc. on Database Systems, Boston, Apr. 1982, pp. 268–282.

Davidson:1984:OCP

[225] Susan B. Davidson. Optimism and consistency in partitioned distributed database systems. ACM Transactions on Database Systems, 9(3):456–481, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1984-9-3/p456-davidson/p456-davidson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1984-9-
                                        3/p456-davidson/.

Ibaraki:1984:ONO

[226] Toshihide Ibaraki and Tiko Kameda. On the optimal nesting order for computing N-relational joins. ACM Transactions on Database Systems, 9(3): 482–502, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1984-9-3/p482-ibaraki/p482-
                                        ibaraki.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1984-
                                        9-3/p482-ibaraki/.

Elhardt:1984:DCH

[227] Klaus Elhardt and Rudolf Bayer. A database cache for high performance and fast restart in database systems. ACM Transactions on Database Systems, 9(4):503–525, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1984-9-4/p503-elhardt/p503-

elhardt.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-4/p503-elhardt/.

Reuter:1984:PAR

[228] Andreas Reuter. Performance analysis of recovery techniques. ACM Transactions on Database Systems, 9(4): 526–559, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1984-9-4/p526-reuter/p526-
reuter.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-4/p526-reuter/.

Effelsberg:1984:PDB

[229] Wolfgang Effelsberg and Theo Haerder. Principles of database buffer management. ACM Transactions on Database Systems, 9(4):560–595, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-4/p560-effelsberg/p560-effelsberg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1984-9-
4/p560-effelsberg/.

Bernstein:1984:ACC

[230] Philip A. Bernstein and Nathan Goodman. An algorithm for concurrency control and recovery in replicated distributed databases. ACM Transactions on Database Systems, 9(4):596–615, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1984-9-4/p596-bernstein/p596-bernstein.
pdf;   http://www.acm.org/pubs/

                                        citations/journals/tods/1984-9-
                                        4/p596-bernstein/.

Chen:1984:ANV

[231] Wen Chin Chen and Jeffrey Scott Vitter. Analysis of new variants of coalesced hashing. ACM Transactions on Database Systems, 9(4):616–645, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1984-9-4/p616-chen/p616-chen.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1984-9-4/p616-chen/

.

Deogun:1984:OCF

[232] J. S. Deogun, V. V. Raghavan, and T. K. W. Tsou. Organization of clustered files for consecutive retrieval. ACM Transactions on Database Systems, 9(4):646–671, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1984-9-4/p646-deogun/p646-
                                        deogun.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1984-
                                        9-4/p646-deogun/.

Traub:1984:SSS

[233] J. F. Traub, Y. Yemini, and H. Wozniakowski. The statistical security of a statistical database. ACM Transactions on Database Systems, 9(4): 672–679, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Navathe:1984:VPA

[234] Shamkant Navathe, Stefano Ceri, Gio Wiederhold, and Jinglie Dou.

Vertical partitioning algorithms for database design. ACM Transactions on Database Systems, 9(4):680–710, December 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1984-9-4/p680-navathe/p680-
navathe.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1984-
9-4/p680-navathe/. Also published

in/as: Stanford Un., TR-CS-82-957, Jan. 1983, revised Aug. 1983.

Maier:1984:DFG

[235] D. Maier. Databases in the fifth generation project: Is Prolog a database language? ACM Transactions on Database Systems, 9(2):??, June 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Also published in/as: ACM SIGMOD, 1984.

Skeen:1984:IAP

[236] D. Skeen and D. D. Wright. Increasing availability in partitioned database systems. ACM Transactions on Database Systems, ??(??):290–299, April 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Franaszek:1985:LCT

[237] Peter Franaszek and John T. Robinson. Limitations of concurrency in transaction processing. ACM Transactions on Database Systems, 10(1):1– 28, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-1/p1-franaszek/p1-
franaszek.pdf;   http://www.acm.
org/pubs/citations/journals/tods/

                                        1985-10-1/p1-franaszek/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/3160.html.

Sacca:1985:DPC

[238] Domenico Sacca and Gio Wiederhold. Database partitioning in a cluster of processors. ACM Transactions on Database Systems, 10(1):29– 56, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1985-10-1/p29-sacca/p29-sacca.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1985-10-1/p29-sacca/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/3161.html.  Also

published in [1090, 242–247], and IBM Research Report No. RJ-4076, 1983.

Pramanik:1985:UGT

[239] Sakti Pramanik and David Ittner. Use of graph-theoretic models for optimal relational database accesses to perform join. ACM Transactions on Database Systems, 10(1):57– 74, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1985-10-1/p57-pramanik/p57-
                                        pramanik.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1985-10-1/p57-pramanik/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/3325.html.

Larson:1985:LHO

[240] Per-Åke Larson. Linear hashing with overflow-handling by linear probing. ACM Transactions on Database Systems, 10(1):75–89, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-1/p75-larson/p75-larson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1985-10-
1/p75-larson/;   http://www.acm.
org/pubs/toc/Abstracts/tods/3324.
html.

Veklerov:1985:ADH

[241] Eugene Veklerov. Analysis of dynamic hashing with deferred splitting. ACM Transactions on Database Systems, 10(1):90–96, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1985-10-1/p90-veklerov/p90-
veklerov.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1985-10-1/p90-veklerov/.

Palvia:1985:EBS

[242] Prashant Palvia. Expressions for batched searching of sequential and hierarchical files. ACM Transactions on Database Systems, 10(1):97– 106, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-1/p97-palvia/p97-palvia.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1985-10-
1/p97-palvia/;   http://www.acm.
org/pubs/toc/Abstracts/tods/3326.
html.

Bever:1985:DHS

[243] Martin Bever and Peter C. Lockemann. Database hosting in stronglytyped programming languages. ACM Transactions on Database Systems, 10 (1):107–126, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1985-10-1/p107-bever/p107-
                                        bever.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1985-
                                        10-1/p107-bever/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        3327.html.

Chen:1985:AAS

[244] Wen Chin Chen and Jeffrey Scott Vitter. Addendum to: “Analysis of Some New Variants of Coalesced Hashing” [ACM trans. database systems 9 (1984), no. 4, 616–645]. ACM Transactions on Database Systems, 10(1):127, March 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Vitter:1985:EIO

[245] Jeffrey Scott Vitter. An efficient I/O interface for optical disks. ACM Transactions on Database Systems, 10(2):129–162, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1985-10-2/p129-vitter/p129-
                                        vitter.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1985-
                                        10-2/p129-vitter/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        3862.html.

Schkolnick:1985:ECU

[246] M. Schkolnick and P. Tiberio. Estimating the cost of updates in a relational database. ACM Transactions on Database Systems, 10(2):163– 179, June 1985. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-2/p163-schkolnick/p163-
schkolnick.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1985-10-2/p163-schkolnick/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/3863.html.

Yu:1985:ARC

[247] C. T. Yu, Cheing-Mei Suen, K. Lam, and M. K. Siu. Adaptive record clustering. ACM Transactions on Database Systems, 10(2):180–204, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-2/p180-yu/p180-yu.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1985-10-2/p180-yu/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/3861.html.

Katoh:1985:CTS

[248] Naoki Katoh, Toshihide Ibaraki, and Tiko Kameda. Cautious transaction schedulers with admission control. ACM Transactions on Database Systems, 10(2):205–229, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1985-10-2/p205-katoh/p205-katoh.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1985-10-
2/p205-katoh/;   http://www.acm.
org/pubs/toc/Abstracts/tods/3860.
html.

Albano:1985:GST

[249] Antonio Albano, Luca Cardelli, and Renzo Orsini. Galileo: a strongly-typed, interactive conceptual language. ACM Transactions on Database Systems, 10(2):230–260, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-2/p230-albano/p230-
albano.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-2/p230-albano/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

3859.html. Also published in [1095].

Huang:1985:HBT

[250] Shou-Hsuan Stephen Huang. Heightbalanced trees of order (β, γ, δ). ACM Transactions on Database Systems, 10(2):261–284, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1985-10-2/p261-huang/p261-
huang.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-2/p261-huang/.

Piwowarski:1985:CBS

[251] Marek Piwowarski. Comments on batched searching of sequential and tree-structured files. ACM Transactions on Database Systems, 10(2):285– 287, June 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-2/p285-piwowarski/p285-
piwowarski.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1985-10-2/p285-piwowarski/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/214294.html. See [16, 168].

Ullman:1985:ILQ

[252] Jeffrey D. Ullman. Implementation of logical query languages for databases. ACM Transactions on Database Systems, 10(3):289–321, September 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-3/p289-ullman/p289-
ullman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-3/p289-ullman/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

3980.html. Sep., YEAR = 1985.

Anonymous:1985:SPA

[253] Anonymous. Selected papers from the 1985 ACM SIGMOD Conference. ACM Transactions on Database Systems, 10 (3):289–346, September 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Batory:1985:MCV

[254] D. S. Batory and Won Kim. Modeling concepts for VLSI CAD objects. ACM Transactions on Database Systems, 10(3):322–346, September 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-3/p322-batory/p322-
batory.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-3/p322-batory/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

4018.html. Also published in/as: ACM-SIGMOD 1985.

Subieta:1985:SQL

[255] Kazimierz Subieta. Semantics of query languages for network databases. ACM Transactions on Database Systems, 10 (3):347–394, September 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1985-10-3/p347-subieta/p347-
subieta.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-3/p347-subieta/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
214293.html.

Liew:1985:DDP

[256] Chong K. Liew, Uinam J. Choi, and Chung J. Liew. A data distortion by probability distribution. ACM Transactions on Database Systems, 10(3):395–411, September 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1985-10-3/p395-liew/p395-liew.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1985-10-3/p395-liew/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/4017.html.

Tay:1985:LPC

[257] Y. C. Tay, Nathan Goodman, and Rajan Suri. Locking performance in centralized databases. ACM Transactions on Database Systems, 10(4):415–462, December 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-4/p415-tay/p415-tay.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1985-10-4/p415-tay/

;   http://www.acm.org/pubs/toc/
Abstracts/tods/4880.html.

Batory:1985:MSA

[258] D. S. Batory. Modeling the storage architectures of commercial database systems. ACM Transactions on Database Systems, 10(4):463–528, December 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1985-10-4/p463-batory/p463-
batory.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1985-
10-4/p463-batory/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
5392.html.

Agrawal:1985:ICC

[259] Rakesh Agrawal and David J. Dewitt. Integrated concurrency control and recovery mechanisms: Design and performance evaluation. ACM Transactions on Database Systems, 10(4):529–564, December 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-4/p529-agrawal/p529-agrawal.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1985-10-
4/p529-agrawal/;   http://www.acm.
org/pubs/toc/Abstracts/tods/4958.
html.

Borgida:1985:LFF

[260] Alexander Borgida. Language features for flexible handling of exceptions in information systems. ACM Transactions on Database Systems, 10(4):565–603, December 1985. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1985-10-4/p565-borgida/p565-borgida.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1985-10-
4/p565-borgida/;   http://www.acm.
org/pubs/toc/Abstracts/tods/4995.

html. Also published in/as: Rutgers Un., TR-LCSR-70, rev. Mar. 1985.

Hagmann:1986:PAS

[261] Robert Brian Hagmann and Domenico Ferrari. Performance analysis of several back-end database architectures. ACM Transactions on Database Systems, 11(1):1–26, March 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1986-11-1/p1-hagmann/p1-hagmann.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
1/p1-hagmann/;   http://www.acm.
org/pubs/toc/Abstracts/tods/5242.
html.

Garcia-Molina:1986:ABA

[262] Héctor García-Molina, Frank Pittelli, and Susan Davidson. Applications of Byzantine agreement in database systems. ACM Transactions on Database Systems, 11(1):27– 47, March 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-1/p27-molina/p27-molina.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
1/p27-molina/;   http://www.acm.
org/pubs/toc/Abstracts/tods/5243.
html.

Segev:1986:OJO

[263] Arie Segev. Optimization of join operations in horizontally partitioned database systems. ACM Transactions on Database Systems, 11(1):48– 80, March 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-1/p48-segev/p48-segev.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1986-11-1/p48-segev/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/5241.html.

Gyssens:1986:CJD

[264] Marc Gyssens. On the complexity of join dependencies. ACM Transactions on Database Systems, 11(1):81– 108, March 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1986-11-1/p81-gyssens/p81-
gyssens.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1986-
11-1/p81-gyssens/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
5237.html.

Sacco:1986:FTE

[265] Giovanni Maria Sacco. Fragmentation: a technique for efficient query processing. ACM Transactions on Database Systems, 11(2):113– 133, June 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-2/p113-sacco/p113-sacco.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
2/p113-sacco/;   http://www.acm.

org/pubs/toc/Abstracts/tods/5638.

html. Also published in/as: University of Torino, TR., Aug. 1983.

Beeri:1986:IAL

[266] Catriel Beeri and Michael Kifer. An integrated approach to logical design of relational database schemes. ACM Transactions on Database Systems, 11(2):134–158, June 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1986-11-2/p134-beeri/p134-beeri.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
2/p134-beeri/;   http://www.acm.
org/pubs/toc/Abstracts/tods/214291.
html.

Mendelson:1986:IIC

[267] Haim Mendelson and Aditya N. Saharia. Incomplete information costs and database design. ACM Transactions on Database Systems, 11(2):159– 185, June 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-2/p159-mendelson/p159-mendelson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
2/p159-mendelson/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
5678.html.

Ginsburg:1986:CTS

[268] Seymour Ginsburg and Katsumi Tanaka. Computation-tuple sequences and object histories. ACM Transactions on Database Systems, 11(2):186– 212, June 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-2/p186-ginsburg/p186-ginsburg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
2/p186-ginsburg/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
5924.html.

Garg:1986:OPK

[269] Anil K. Garg and C. C. Gotlieb. Orderpreserving key transformations. ACM Transactions on Database Systems, 11(2):213–234, June 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1986-11-2/p213-garg/p213-garg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
2/p213-garg/;   http://www.acm.
org/pubs/toc/Abstracts/tods/5923.
html.

Shapiro:1986:JPD

[270] Leonard D. Shapiro. Join processing in database systems with large main memories. ACM Transactions on Database Systems, 11(3):239–264, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-3/p239-shapiro/p239-shapiro.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
3/p239-shapiro/;   http://www.acm.
org/pubs/toc/Abstracts/tods/6315.
html.

Gavish:1986:SQO

[271] Bezalel Gavish and Arie Segev. Set query optimization in distributed database systems. ACM Transactions on Database Systems, 11(3):265–293, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1986-11-3/p265-gavish/p265-
gavish.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1986-
11-3/p265-gavish/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
6488.html.

Lafortune:1986:STM

[272] Stéphane Lafortune and Eugene Wong. A state transition model for distributed query processing. ACM Transactions on Database Systems, 11(3):294–322, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-3/p294-lafortune/p294-lafortune.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
3/p294-lafortune/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
6460.html.

Lozinskii:1986:POI

[273] Eliezer L. Lozinskii. A problemoriented inferential database system. ACM Transactions on Database Systems, 11(3):323–356, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1986-11-3/p323-lozinskii/p323-lozinskii.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
3/p323-lozinskii/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
6419.html.

Osborn:1986:DRD

[274] Sylvia L. Osborn and T. E. Heaven. The design of a relational database system with abstract data types for domains. ACM Transactions on Database Systems, 11(3):357–373, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1986-11-3/p357-osborn/p357-
osborn.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1986-
11-3/p357-osborn/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
6461.html.

Gawlick:1986:RIW

[275] Dieter Gawlick. Report on the international workshop on high-performance transaction systems. ACM Transactions on Database Systems, 11(4): 375–377, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1986-11-4/p375-gawlick/p375-
gawlick.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1986-
11-4/p375-gawlick/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
17346.html.

Mohan:1986:TMR

[276] C. Mohan, B. Lindsay, and R. Obermarck. Transaction management in the R* distributed database management system. ACM Transactions on Database Systems, 11(4):378–396, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/

                                        1986-11-4/p378-mohan/p378-mohan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1986-11-
                                        4/p378-mohan/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/7266.
                                        html.

Bayer:1986:CTR

[277] Rudolf Bayer. Consistency of transactions and random batch. ACM Transactions on Database Systems, 11(4): 397–404, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1986-11-4/p397-bayer/p397-
                                        bayer.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1986-
                                        11-4/p397-bayer/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        214287.html.

ONeil:1986:ETM

[278] Patrick E. O’Neil. The Escrow transactional method. ACM Transactions on Database Systems, 11(4):405–430, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1986-11-4/p405-o_neil/p405-
                                        o_neil.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1986-
                                        11-4/p405-o_neil/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        7265.html.

Hsu:1986:PTP

[279] Meichun Hsu and Arvola Chan. Partitioned two-phase locking. ACM Transactions on Database Systems, 11(4): 431–446, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1986-11-4/p431-hsu/p431-hsu.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1986-11-4/p431-hsu/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/7477.html.

Luk:1986:EEL

[280] W. S. Luk and Steve Kloster. ELFS: English language from SQL. ACM Transactions on Database Systems, 11 (4):447–472, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sacco:1986:BMR

[281] Giovanni Maria Sacco and Mario Schkolnick. Buffer management in relational database systems. ACM Transactions on Database Systems, 11(4): 473–498, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1986-11-4/p473-sacco/p473-
sacco.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1986-
11-4/p473-sacco/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
7336.html.

Ariav:1986:TOD

[282] Gad Ariav. A temporally oriented data model. ACM Transactions on Database Systems, 11(4):499–527, December 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1986-11-4/p499-ariav/p499-ariav.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1986-11-
4/p499-ariav/;   http://www.acm.

                                        org/pubs/toc/Abstracts/tods/7350.
                                        html.

Albano:1986:OSG

[283] Antonio Albano, S. Alfo, Luca Cardelli, and Renzo Orsini. An overview of SIDEREUS: a graphical database schema editor for Galileo. ACM Transactions on Database Systems, 11 (??):568–571, ???? 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Durand:1986:FMS

[284] Charles Durand. Forward multidimensional search with applications to information retrieval. ACM Transactions on Database Systems, ??(??):??, September 1986. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Submitted.

Spyratos:1987:PMD

[285] Nicolas Spyratos. The partition model: a deductive database model. ACM Transactions on Database Systems, 12(1):1–37, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1987-12-1/p1-spyratos/p1-spyratos.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        1/p1-spyratos/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/22718.

html. Also published in/as: Institut National de la Recherche en Informatique et Automatique, TR-286, Apr. 1983.

Wu:1987:ASM

[286] C. T. Wu and Walter A. Burkhard. Associative searching in multiple storage units. ACM Transactions on Database Systems, 12(1):38–64, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1987-12-1/p38-wu/p38-wu.pdf; http:
//www.acm.org/pubs/citations/journals/
tods/1987-12-1/p38-wu/;   http:/
/www.acm.org/pubs/toc/Abstracts/

tods/12048.html. Discusses interpolation hashing, a multidimensional variant of linear hashing.

Lomet:1987:PEF

[287] David B. Lomet. Partial expansions for file organizations with an index. ACM Transactions on Database Systems, 12(1):65–84, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1987-12-1/p65-lomet/p65-lomet.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1987-12-1/p65-lomet/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/12049.html.

Fedorowicz:1987:DPE

[288] Jane Fedorowicz. Database performance evaluation in an indexed file environment. ACM Transactions on Database Systems, 12(1):85– 110, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1987-12-1/p85-fedorowicz/p85-fedorowicz.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1987-12-
1/p85-fedorowicz/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
13675.html.

Ozsoyoglu:1987:NNF

[289] Z. Meral Özsoyoğlu and Li-Yan Yuan. A new normal form for nested relations. ACM Transactions on Database Systems, 12(1):111–136, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-1/p111-ozsoyoglu/p111-ozsoyoglu.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        1/p111-ozsoyoglu/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        13676.html.

Christodoulakis:1987:ARP

[290] Stavros Christodoulakis. Analysis of retrieval performance for records and objects using optical disk technology. ACM Transactions on Database Systems, 12(2):137–169, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1987-12-2/p137-christodoulakis/
                                        p137-christodoulakis.pdf;   http:
                                        //www.acm.org/pubs/citations/journals/
                                        tods/1987-12-2/p137-christodoulakis/
                                        ; http://www.acm.org/pubs/toc/Abstracts/
                                        tods/23015.html.

Herlihy:1987:DQA

[291] Maurice Herlihy. Dynamic quorum adjustment for partitioned data. ACM Transactions on Database Systems, 12(2):170–194, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1987-12-2/p170-herlihy/p170-
                                        herlihy.pdf;   http://www.acm.org/

pubs/citations/journals/tods/1987-
12-2/p170-herlihy/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
22953.html.

Ellis:1987:CLH

[292] Carla Schlatter Ellis. Concurrency in linear hashing. ACM Transactions on Database Systems, 12(2):195– 217, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1987-12-2/p195-ellis/p195-ellis.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1987-12-
2/p195-ellis/;   http://www.acm.
org/pubs/toc/Abstracts/tods/22954.

html. Also published in [1091].

Valduriez:1987:JI

[293] Patrick Valduriez. Join indices. ACM Transactions on Database Systems, 12(2):218–246, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1987-12-2/p218-valduriez/p218-
valduriez.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1987-12-2/p218-valduriez/;   http:
//www.acm.org/pubs/toc/Abstracts/

tods/22955.html. Compares join indices with inverted indices, clustered indices, B+ trees, linked lists, and hybrid hash techniques.

Snodgrass:1987:TQL

[294] Richard Snodgrass. The Temporal Query Language TQUEL. ACM Transactions on Database Systems, 12(2):247–298, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1987-12-2/p247-snodgrass/p247-
                                        snodgrass.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1987-12-2/p247-snodgrass/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/22956.html.

Wong:1987:MIR

[295] S. K. M. Wong, W. Ziarko, V. V. Raghavan, and P. C. N. Wong. On modeling of information retrieval concepts in vector spaces. ACM Transactions on Database Systems, 12(2):299– 321, June 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-2/p299-wong/p299-wong.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1987-12-2/p299-wong/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/22957.html.

Rybinski:1987:FOL

[296] Henryk Rybiński. On first-order-logic databases. ACM Transactions on Database Systems, 12(3):325–349, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-3/p325-rybinski/p325-rybinski.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        3/p325-rybinski/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        27630.html.

Stonebraker:1987:EDS

[297] Michael Stonebraker, Jeff Anton, and Eric Hanson. Extending a database system with procedures. ACM Transactions on Database Systems, 12(3):350–376, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1987-12-3/p350-stonebraker/
p350-stonebraker.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1987-12-3/p350-stonebraker/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/27631.html. Also

published in/as: UCB/ERL memo M85/59, 1985.

Ozsoyoglu:1987:RMM

[298] Z. Meral Özsoyoğlu and Li-Yan Yuan. Reduced MVDs and minimal covers. ACM Transactions on Database Systems, 12(3):377–394, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1987-12-3/p377-ozsoyoolu/;
http://www.acm.org/pubs/toc/Abstracts/
tods/214286.html.

Faloutsos:1987:OSE

[299] Christos Faloutsos and Stavros Christodoulakis. Optimal signature extraction and information loss. ACM Transactions on Database Systems, 12(3):395–428, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1987-12-3/p395-faloutsos/p395-faloutsos.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1987-12-
3/p395-faloutsos/;   http://www.

                                        acm.org/pubs/toc/Abstracts/tods/
                                        214285.html.

Ibaraki:1987:SC

[300] Toshihide Ibaraki, Tiko Kameda, and Toshimi Minoura. Serializability with constraints. ACM Transactions on Database Systems, 12(3):429–452, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-3/p429-ibaraki/p429-ibaraki.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        3/p429-ibaraki/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/214284.
                                        html.

Wolfson:1987:OLC

[301] Ouri Wolfson. The overhead of locking (and commit) protocols in distributed databases. ACM Transactions on Database Systems, 12(3):453–471, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-3/p453-wolfson/p453-wolfson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        3/p453-wolfson/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/28053.
                                        html.

Su:1987:CBD

[302] Stanley Y. W. Su, Jozo Dujmovic, D. S. Batory, S. B. Navathe, and Richard Elnicki. A cost-benefit decision model: Analysis, comparison, and selection of data management systems. ACM Transactions on Database Systems, 12(3):472–520, September 1987. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1987-12-3/p472-su/p472-su.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1987-12-3/p472-su/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/33403.html.

Keller:1987:CBS

[303] Arthur M. Keller. Comment on Bancilhon and Spyratos’ “Update semantics and relational views”. ACM Transactions on Database Systems, 12(3): 521–523, September 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1987-12-3/p521-keller/p521-
keller.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1987-
12-3/p521-keller/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
214296.html. See [138].

Abiteboul:1987:IFS

[304] Serge Abiteboul and Richard Hull. IFO: a formal semantic database model. ACM Transactions on Database Systems, 12(4):525–565, December 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1987-12-4/p525-abiteboul/p525-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1987-12-
4/p525-abiteboul/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

32205.html. Also published in/as: ACM SIGACT-SIGMOD Symp. on Principles of Database Systems, 1984 (short version). Also published in/as: USC CSD, TR-84-304, Apr. 1984.

Ozsoyoglu:1987:ERA

[305] G. Özsoyoğlu, Z. M. Özsoyoğlu, and V. Matos. Extending relational algebra and relational calculus with set-valued attributes and aggregate functions. ACM Transactions on Database Systems, 12(4):566–592, December 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1987-12-4/p566-ozsoyoglu/p566-ozsoyoglu.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1987-12-
                                        4/p566-ozsoyoglu/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        32219.html.

Palley:1987:URM

[306] Michael A. Palley and Jeffrey S. Simonoff. The use of regression methodology for the compromise of confidential information in statistical databases. ACM Transactions on Database Systems, 12(4):593–608, December 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1987-12-4/p593-palley/p593-
                                        palley.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1987-
                                        12-4/p593-palley/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        42174.html.

Agrawal:1987:CCP

[307] Rakesh Agrawal, Michael J. Carey, and Miron Livny. Concurrency control performance modeling: Alternatives and implications. ACM Transactions on Database Systems, 12(4):609–654, December 1987. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1987-12-4/p609-agrawal/p609-agrawal.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1987-12-
4/p609-agrawal/;   http://www.acm.
org/pubs/toc/Abstracts/tods/32220.

html. Also published in/as: ACM SIGMOD Conf. on the Management of Data, 1985.

Sacks-Davis:1987:MAM

[308] R. Sacks-Davis, A. Kent, and K. Ramamohanarao. Multikey access methods based on superimposed coding techniques. ACM Transactions on Database Systems, 12(4):655–696, December 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1987-12-4/p655-sacks-davis/
p655-sacks-davis.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1987-12-4/p655-sacks-davis/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/32222.html.

Elhardt:1987:SQO

[309] K. Elhardt. Support for query optimization by optimal predicate splitting. ACM Transactions on Database Systems, ??(??):??, March 1987. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Accepted.

Vianu:1988:DFO

[310] Victor Vianu. A dynamic framework for object projection views. ACM Transactions on Database Systems, 13(1):1–22, March 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1988-13-1/p1-vianu/p1-vianu.
                                        pdf; http://www.acm.org/pubs/citations/
                                        journals/tods/1988-13-1/p1-vianu/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/42202.html.

Sellis:1988:MQO

[311] Timos K. Sellis. Multiple-query optimization. ACM Transactions on Database Systems, 13(1):23–52, March 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1988-13-1/p23-sellis/p23-sellis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1988-13-
                                        1/p23-sellis/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/42203.
                                        html.

Shasha:1988:CSS

[312] Dennis Shasha and Nathan Goodman. Concurrent search structure algorithms. ACM Transactions on Database Systems, 13(1):53–90, March 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1988-13-1/p53-shasha/p53-shasha.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1988-13-
                                        1/p53-shasha/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/42204.
                                        html.

Finkelstein:1988:PDD

[313] S. J. Finkelstein, M. Schkolnick, and P. Tiberio. Physical database design for relational databases. ACM Transactions on Database Systems, 13(1):91– 128, March 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1988-13-1/p91-finkelstein/
p91-finkelstein.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1988-13-1/p91-finkelstein/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/42205.html. Also

published in/as: IBM Research Report No. RJ5034, Nov. 1986, preprint IBM Aug. 1985.

Raju:1988:FFD

[314] K. V. S. V. N. Raju and Arun K. Majumdar. Fuzzy functional dependencies and lossless join decomposition of fuzzy relational database systems. ACM Transactions on Database Systems, 13(2):129–166, June 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1988-13-2/p129-raju/p129-raju.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1988-13-2/p129-raju/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/42344.html.

Winslett:1988:MBA

[315] Marianne Winslett. A model-based approach to updating databases with incomplete information. ACM Transactions on Database Systems, 13(2):167– 196, June 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1988-13-2/p167-winslett/p167-winslett.
pdf;   http://www.acm.org/pubs/

                                        citations/journals/tods/1988-13-
                                        2/p167-winslett/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        42386.html.

Eich:1988:DCC

[316] Margaret H. Eich and David L. Wells. Database concurrency control using data flow graphs. ACM Transactions on Database Systems, 13(2):197– 227, June 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1988-13-2/p197-eich/p197-eich.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1988-13-2/p197-eich/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/42345.html.

Batory:1988:ICE

[317] D. S. Batory, T. Y. Leung, and T. E. Wise. Implementation concepts for an extensible data model and data language. ACM Transactions on Database Systems, 13(3):231–262, September 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1988-13-3/p231-batory/p231-
                                        batory.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1988-
                                        13-3/p231-batory/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/

45062.html. Also published in/as: Un. Texas, TR-86-24, Oct. 1986.

Apers:1988:DAD

[318] Peter M. G. Apers. Data allocation in distributed database systems. ACM Transactions on Database Systems, 13(3):263–304, September 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1988-13-3/p263-apers/p263-apers.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1988-13-
3/p263-apers/;   http://www.acm.
org/pubs/toc/Abstracts/tods/45063.
html.

Storey:1988:MCU

[319] Veda C. Storey and Robert C. Goldstein. A methodology for creating user views in database design. ACM Transactions on Database Systems, 13(3):305–338, September 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1988-13-3/p305-storey/p305-
storey.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1988-
13-3/p305-storey/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
45064.html.

Diederich:1988:NMF

[320] Jim Diederich and Jack Milton. New methods and fast algorithms for database normalization. ACM Transactions on Database Systems, 13(3): 339–365, September 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1988-13-3/p339-diederich/p339-
diederich.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1988-13-3/p339-diederich/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/44499.html. Also published

in/as: UCD, Math, revised Jun. 1987.

Larson:1988:LHS

[321] Per-Åke Larson. Linear hashing with separators — A dynamic hashing scheme achieving one-access retrieval. ACM Transactions on Database Systems, 13(3):366–388, September 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1988-13-3/p366-larson/p366-
                                        larson.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1988-
                                        13-3/p366-larson/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        44500.html.

Roth:1988:EAC

[322] Mark A. Roth, Henry F. Korth, and Abraham Silberschatz. Extended algebra and calculus for nested relational databases. ACM Transactions on Database Systems, 13(4):389–417, December 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1988-13-4/p389-roth/p389-roth.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1988-13-4/p389-roth/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/49347.html.  See

comment [403].

Gadia:1988:HRM

[323] Shashi K. Gadia. A homogeneous relational model and query languages for temporal databases. ACM Transactions on Database Systems, 13(4): 418–448, December 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/

tods/1988-13-4/p418-gadia/p418-
gadia.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1988-
13-4/p418-gadia/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
50065.html.

Brosda:1988:URR

[324] Volkert Brosda and Gottfried Vossen. Update and retrieval in a relational database through a universal schema interface. ACM Transactions on Database Systems, 13(4):449–485, December 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1988-13-4/p449-brosda/p449-
brosda.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1988-
13-4/p449-brosda/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
49884.html.

Gottlob:1988:PUS

[325] Georg Gottlob, Paolo Paolini, and Roberto Zicari. Properties and update semantics of consistent views. ACM Transactions on Database Systems, 13 (4):486–524, December 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1988-13-4/p486-gottlob/p486-
gottlob.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1988-
13-4/p486-gottlob/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
50068.html.

Lomet:1988:SBD

[326] David B. Lomet. A simple bounded disorder file organization with good performance. ACM Transactions on Database Systems, 13(4):525–551, December 1988. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1988-13-4/p525-lomet/p525-lomet.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1988-13-
4/p525-lomet/;   http://www.acm.
org/pubs/toc/Abstracts/tods/50067.
html.

Freytag:1989:TRQ

[327] Johann Christoph Freytag and Nathan Goodman. On the translation of relational queries into iterative programs. ACM Transactions on Database Systems, 14(1):1–27, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1989-14-1/p1-freytag/p1-freytag.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
1/p1-freytag/.

Ahad:1989:ECP

[328] Rafiul Ahad, K. V. Bapa Rao, and Dennis McLeod. On estimating the cardinality of the projection of a database relation. ACM Transactions on Database Systems, 14(1):28– 40, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1989-14-1/p28-ahad/p28-ahad.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1989-14-1/p28-ahad/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/62034.html.

Ramesh:1989:VDT

[329] R. Ramesh, A. J. G. Babu, and J. Peter Kincaid. Variable-depth trie index optimization: Theory and experimental results. ACM Transactions on Database Systems, 14(1):41– 74, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1989-14-1/p41-ramesh/p41-ramesh.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
1/p41-ramesh/;   http://www.acm.
org/pubs/toc/Abstracts/tods/77249.
html.

Gladney:1989:DRD

[330] H. M. Gladney. Data replicas in distributed information services. ACM Transactions on Database Systems, 14(1):75–97, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1989-14-1/p75-gladney/p75-
gladney.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-1/p75-gladney/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

62035.html. Reviews: Computing Reviews, Vol. 30, No. 11, November 1989.

McLeish:1989:FRS

[331] Mary McLeish. Further results on the security of partitioned dynamic statistical databases. ACM Transactions on Database Systems, 14(1):98– 113, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-1/p98-mcleish/p98-

                                        mcleish.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1989-
                                        14-1/p98-mcleish/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        62036.html.

Bic:1989:ADD

[332] Lubomir Bic and Robert L. Hartmann. AGM: a dataflow database machine. ACM Transactions on Database Systems, 14(1):114–146, March 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1989-14-1/p114-bic/p114-bic.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tods/1989-14-1/p114-bic/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/62037.html.

Yu:1989:FER

[333] C. T. Yu, W. Meng, and S. Park. A framework for effective retrieval. ACM Transactions on Database Systems, 14(2):147–167, June 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1989-14-2/p147-yu/p147-yu.
                                        pdf; http://www.acm.org/pubs/citations/
                                        journals/tods/1989-14-2/p147-yu/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/63519.html.

Embley:1989:NNF

[334] David W. Embley. NFQL: The Natural Forms Query Language. ACM Transactions on Database Systems, 14(2):168–211, June 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1989-14-2/p168-embley/p168-

embley.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-2/p168-embley/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

64125.html. Also published in/as: Brigham Young Un., TR-CS-87-6, Mar. 1987.

Atzeni:1989:EOS

[335] Paolo Atzeni and Edward P. F. Chan. Efficient optimization of simple chase join expressions. ACM Transactions on Database Systems, 14(2):212– 230, June 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-2/p212-atzeni/p212-
atzeni.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-2/p212-atzeni/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
63520.html.

Ramakrishna:1989:FOU

[336] M. V. Ramakrishna and Per-Åke Larson. File organization using composite perfect hashing. ACM Transactions on Database Systems, 14(2):231– 263, June 1, 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-2/p231-ramakrishna/
p231-ramakrishna.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1989-14-2/p231-ramakrishna/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/63521.html.

ElAbaddi:1989:MAP

[337] Amr El Abaddi and Sam Toueg. Maintaining availability in partitioned replicated databases. ACM Transactions on Database Systems, 14(2):264– 290, June 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-2/p264-abbadi/p264-
abbadi.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-2/p264-abbadi/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

63501.html. Also published in/as: ACM SIGACT-SIGMOD Symp. on Principles of Database Systems, Cambridge MA, Mar. 1986.

Hudson:1989:CSA

[338] Scott E. Hudson and Roger King. Cactis: a self-adaptive, concurrent implementation of an objectoriented database management system. ACM Transactions on Database Systems, 14(3):291–321, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-3/p291-hudson/p291-
hudson.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-3/p291-hudson/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
68013.html.

Sheard:1989:AVD

[339] Tim Sheard and David Stemple. Automatic verification of database transaction safety. ACM Transactions on Database Systems, 14(3):322–368, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-3/p322-sheard/p322-

sheard.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-3/p322-sheard/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

68014.html. Also published in/as: U. Mass, COINS TR-88-29, Apr. 1988, also TR-86-30, 1986.

Blakeley:1989:UDR

[340] José A. Blakeley, Neil Coburn, and Per-Åke Larson. Updating derived relations: Detecting irrelevant and autonomously computable updates. ACM Transactions on Database Systems, 14(3):369–400, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1989-14-3/p369-blakeley/p369-blakeley.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
3/p369-blakeley/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

68015.html. Also published in/as: Indiana Un., TR-235, Nov. 1987. Also published in [1092].

Mackert:1989:ISU

[341] Lothar F. Mackert and Guy M. Lohman. Index scans using a finite LRU buffer: a validated I/O model. ACM Transactions on Database Systems, 14(3):401–424, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1989-14-3/p401-mackert/p401-mackert.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
3/p401-mackert/;   http://www.acm.
org/pubs/toc/Abstracts/tods/68016.

html. Also published in/as: IBM, TR-RC-4836, Sep. 1985.

Manolopoulos:1989:PTH

[342] Y. Manolopoulos and J. G. Kollias. Performance of a two-headed disk system when serving database queries under the scan policy. ACM Transactions on Database Systems, 14(3):425–442, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1989-14-3/p425-manolopoulos/p425-
                                        manolopoulos.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tods/1989-14-3/p425-manolopoulos/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/68017.html.

Dreizen:1989:ISR

[343] Howard M. Dreizen and Shi-Kuo Chang. Imprecise schema: a rationale for relations with embedded subrelations. ACM Transactions on Database Systems, 14(4):447–479, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1989-14-4/p447-dreizen/p447-dreizen.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1989-14-
                                        4/p447-dreizen/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/76903.
                                        html.

Motro:1989:IVC

[344] Amihai Motro. Integrity = validity + completeness. ACM Transactions on Database Systems, 14(4):480–502, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1989-14-4/p480-motro/p480-motro.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
4/p480-motro/;   http://www.acm.
org/pubs/toc/Abstracts/tods/76904.
html.

Farrag:1989:USK

[345] Abdel Aziz Farrag and M. Tamer Özsu. Using semantic knowledge of transactions to increase concurrency. ACM Transactions on Database Systems, 14(4):503–525, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1989-14-4/p503-farrag/p503-
farrag.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1989-
14-4/p503-farrag/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

76905.html. Accepted. Also published in/as: Un. Alberta, DCS, TR-85-11, Jul. 1985.

Ozsoyoglu:1989:QPT

[346] Gultekin Özsoyoğlu, Victor Matos, and Z. Meral Özsoyoğlu. Query processing techniques in the summarytable-by-example database query language. ACM Transactions on Database Systems, 14(4):526–573, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1989-14-4/p526-ozsoyoglu/p526-ozsoyoglu.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-
4/p526-ozsoyoglu/;   http://www.

                                        acm.org/pubs/toc/Abstracts/tods/
                                        76906.html.

Grady:1989:EJO

[347] Danièle Grady and Claude Puech. On the effect of join operations on relation sizes. ACM Transactions on Database Systems, 14(4):574–603, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1989-14-4/p574-grady/p574-grady.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1989-14-
                                        4/p574-grady/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/76907.
                                        html.

Lang:1989:UAB

[348] Sheau-Dong Lang, James R. Driscoll, and Jiann H. Jou. A unified analysis of batched searching of sequential and tree-structured files. ACM Transactions on Database Systems, 14 (4):604–618, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1989-14-4/p604-lang/p604-lang.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1989-14-
                                        4/p604-lang/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/76908.
                                        html.

Motro:1989:QDK

[349] A. Motro and Q. Yuan. Querying database knowledge. ACM Transactions on Database Systems, 14(4):??, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Also published in [1094].

Liu:1990:IMI

[350] Ken-Chih C. Liu and Rajshekhar Sunderraman. Indefinite and maybe information in relational databases. ACM Transactions on Database Systems, 15(1):1–39, March 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1990-15-1/p1-liu/p1-liu.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1990-15-1/p1-liu/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/77644.html.

Langerak:1990:VUR

[351] Rom Langerak. View updates in relational databases with an independent scheme. ACM Transactions on Database Systems, 15(1):40– 66, March 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1990-15-1/p40-langerak/p40-
langerak.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1990-15-1/p40-langerak/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/77645.html.

Whang:1990:QOM

[352] Kyu-Young Y. Whang and Ravi Krishnamurthy. Query optimization in a memory-resident domain relational calculus database system. ACM Transactions on Database Systems, 15(1):67– 95, March 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1990-15-1/p67-whang/p67-whang.pdf;
 http://www.acm.org/pubs/citations/

                                        journals/tods/1990-15-1/p67-whang/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/77646.html.

Herlihy:1990:AVA

[353] Maurice Herlihy. Apologizing versus asking permission: Optimistic concurrency control for abstract data types. ACM Transactions on Database Systems, 15(1):96–124, March 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1990-15-1/p96-herlihy/p96-
                                        herlihy.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1990-
                                        15-1/p96-herlihy/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        77647.html.

Wald:1990:EAF

[354] Joseph A. Wald and Paul G. Sorenson. Explaining ambiguity in a formal query language. ACM Transactions on Database Systems, 15(2):125– 161, June 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-2/p125-wald/p125-wald.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1990-15-2/p125-wald/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/78923.html.

Chakravarthy:1990:LBA

[355] Upen S. Chakravarthy, John Grant, and Jack Minker. Logic-based approach to semantic query optimization. ACM Transactions on Database Systems, 15(2):162–207, June 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1990-15-2/p162-chakravarthy/p162-
chakravarthy.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1990-15-2/p162-chakravarthy/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/78924.html.

Whang:1990:LTP

[356] Kyu-Young Whang, Brad T. Vander-Zanden, and Howard M. Taylor. A linear-time probabilistic counting algorithm for database applications. ACM Transactions on Database Systems, 15(2):208–229, June 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1990-15-2/p208-whang/p208-whang.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1990-15-
2/p208-whang/;   http://www.acm.
org/pubs/toc/Abstracts/tods/78925.

html. A probabilistic technique called linear counting, based on hashing, for counting the number of unique values in the presence of duplicates is presented in this paper.

Jajodia:1990:DVA

[357] Sushil Jajodia and David Mutchler. Dynamic voting algorithms for maintaining the consistency of a replicated database. ACM Transactions on Database Systems, 15(2):230– 280, June 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1990-15-2/p230-jajodia/p230-jajodia.
pdf;   http://www.acm.org/pubs/

                                        citations/journals/tods/1990-15-
                                        2/p230-jajodia/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/78926.
                                        html.

Dasgupta:1990:FCC

[358] Partha Dasgupta and Zvi M. Kedem. The five-color concurrency control protocol: Non-two-phase locking in general databases. ACM Transactions on Database Systems, 15(2):281– 307, June 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-2/p281-dasgupta/p281-dasgupta.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1990-15-
                                        2/p281-dasgupta/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        78927.html.

Moore:1990:DTA

[359] James C. Moore, William B. Richmond, and Andrew B. Whinston. A decision-theoretic approach to information retrieval. ACM Transactions on Database Systems, 15(3):311–340, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-3/p311-moore/p311-moore.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1990-15-
                                        3/p311-moore/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/88597.
                                        html.

Westland:1990:SOC

[360] J. Christopher Westland. Scaling up output capacity and performance results from information systems prototypes. ACM Transactions on Database Systems, 15(3):341–358, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1990-15-3/p341-westland/p341-westland.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1990-15-
3/p341-westland/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
87943.html.

Alonso:1990:DCI

[361] Rafael Alonso, Daniel Barbara, and Héctor García-Molina. Data caching issues in an information retrieval system. ACM Transactions on Database Systems, 15(3):359–384, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1990-15-3/p359-alonso/p359-
alonso.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1990-
15-3/p359-alonso/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
87848.html.

Kifer:1990:CTQ

[362] Michael Kifer and Eliezer L. Lozinskii. On compile-time query optimization in deductive databases by means of static filtering. ACM Transactions on Database Systems, 15(3):385–426, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1990-15-3/p385-kifer/p385-kifer.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1990-15-
3/p385-kifer/;   http://www.acm.

                                        org/pubs/toc/Abstracts/tods/87121.
                                        html.

Agrawal:1990:DTC

[363] Rakesh Agrawal, Shaul Dar, and H. V. Jagadish. Direct transitive closure algorithms: Design and performance evaluation. ACM Transactions on Database Systems, 15(3):427–458, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-3/p427-agrawal/p427-agrawal.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1990-15-
                                        3/p427-agrawal/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/88888.
                                        html.

Zhang:1990:NSC

[364] Weining Zhang, Clement T. Yu, and Daniel Troy. Necessary and sufficient conditions to linearize doubly recursive programs in logic databases. ACM Transactions on Database Systems, 15(3):459–482, September 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-3/p459-zhang/p459-zhang.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1990-15-
                                        3/p459-zhang/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/89237.
                                        html.

Can:1990:CEC

[365] Fazli Can and Esen A. Ozkarahan. Concepts and effectiveness of the covercoefficient-based clustering methodology for text databases. ACM Transactions on Database Systems, 15(4):

483–517, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1990-15-4/p483-can/p483-can.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1990-15-4/p483-can/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/99938.html.

Nakano:1990:TOR

[366] Ryohei Nakano. Translation with optimization from relational calculus to relational algebra having aggregate functions. ACM Transactions on Database Systems, 15(4):518–557, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1990-15-4/p518-nakano/p518-
nakano.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1990-
15-4/p518-nakano/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
99943.html.

Jagadish:1990:CTM

[367] H. V. Jagadish. A compression technique to materialize transitive closure. ACM Transactions on Database Systems, 15(4):558–598, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1990-15-4/p558-jagadish/p558-jagadish.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1990-15-
4/p558-jagadish/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
99944.html.

Omiecinski:1990:PAR

[368] Edward Omiecinski and Peter Scheuermann. A parallel algorithm for record clustering. ACM Transactions on Database Systems, 15(4):599–624, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-4/p599-omiecinski/p599-
                                        omiecinski.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1990-15-4/p599-omiecinski/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/99947.html.

Lomet:1990:HTM

[369] David B. Lomet and Betty Salzberg. The hB-Tree: a multiattribute indexing method with good guaranteed performance. ACM Transactions on Database Systems, 15(4):625–658, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1990-15-4/p625-lomet/p625-lomet.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1990-15-
                                        4/p625-lomet/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/99949.
                                        html.

Abiteboul:1991:RBL

[370] Serge Abiteboul and Stéphane Grumbach. A rule-based language with functions and sets. ACM Transactions on Database Systems, 16(1):1– 30, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1991-16-1/p1-abiteboul/p1-

abiteboul.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1991-16-1/p1-abiteboul/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/103141.html.

Mendelzon:1991:FDH

[371] Alberto O. Mendelzon and Peter T. Wood. Functional dependencies in Horn clause queries. ACM Transactions on Database Systems, 16(1):31– 55, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1991-16-1/p31-mendelzon/p31-mendelzon.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
1/p31-mendelzon/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
103142.html.

Meghini:1991:COF

[372] C. Meghini and C. Thanos. The complexity of operations on a fragmented relation. ACM Transactions on Database Systems, 16(1):56– 87, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1991-16-1/p56-meghini/p56-
meghini.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1991-
16-1/p56-meghini/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
103143.html.

Rabitti:1991:MAN

[373] Fausto Rabitti, Elisa Bertino, Won Kim, and Darrell Woelk. A model of authorization for next-generation database systems. ACM Transactions on Database Systems, 16(1):88– 131, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1991-16-1/p88-rabitti/p88-
rabitti.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1991-
16-1/p88-rabitti/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
103144.html.

Weikum:1991:PRS

[374] Gerhard Weikum. Principles and realization strategies of multilevel transaction management. ACM Transactions on Database Systems, 16(1):132– 180, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1991-16-1/p132-weikum/p132-
weikum.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1991-
16-1/p132-weikum/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
103145.html.

Wolfson:1991:MPR

[375] Ouri Wolfson and Amir Milo. The multicast policy and its relationship to replicated data placement. ACM Transactions on Database Systems, 16 (1):181–205, March 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1991-16-1/p181-wolfson/p181-
wolfson.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1991-
16-1/p181-wolfson/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

103146.html. See corrigendum in

Casanova:1991:STM

[376] M. A. Casanova, A. L. Furtado, and L. Tucherman. A software tool for modular database design. ACM Transactions on Database Systems, 16(2):209– 234, June 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1991-16-2/p209-casanova/p209-casanova.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
2/p209-casanova/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
103711.html.

VanGelder:1991:STR

[377] Allen Van Gelder and Rodney W. Topor. Safety and translation of relational calculus queries. ACM Transactions on Database Systems, 16(2):235– 278, June 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1991-16-2/p235-van_gelder/p235-
van_gelder.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1991-16-2/p235-van_gelder/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/103712.html.

Shasha:1991:OEQ

[378] Dennis Shasha and Tsong-Li L. Wang. Optimizing equijoin queries in distributed databases where relations are hash-partitioned. ACM Transactions on Database Systems, 16(2):279–308, June 1, 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1991-16-2/p279-shasha/p279-

                                        shasha.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1991-
                                        16-2/p279-shasha/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        103713.html.

Cesarini:1991:DHM

[379] F. Cesarini and G. Soda. A dynamic hash method with signature. ACM Transactions on Database Systems, 16(2):309–337, June 1, 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1991-16-2/p309-cesarini/p309-cesarini.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1991-16-
                                        2/p309-cesarini/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        103714.html.

King:1991:MRB

[380] Richard P. King, Nagui Halim, Héctor García-Molina, and Christos A. Polyzois. Management of a remote backup copy for disaster recovery. ACM Transactions on Database Systems, 16(2):338–368, June 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1991-16-2/p338-king/p338-king.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1991-16-2/p338-king/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/103715.html.

Gogolla:1991:TSV

[381] Martin Gogolla and Uwe Hohenstein. Towards a semantic view of an extended entity-relationship model. ACM Transactions on Database Systems, 16(3):369–416, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1991-16-3/p369-gogolla/p369-gogolla.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
3/p369-gogolla/;   http://www.acm.
org/pubs/toc/Abstracts/tods/111200.
html.

Sciore:1991:UAS

[382] Edward Sciore. Using annotations to support multiple kinds of versioning in an object- oriented database system. ACM Transactions on Database Systems, 16(3):417–438, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1991-16-3/p417-sciore/p417-
sciore.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1991-
16-3/p417-sciore/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
111205.html.

Karabeg:1991:SRC

[383] Dino Karabeg and Victor Vianu. Simplification rules and complete axiomatization for relational update transactions. ACM Transactions on Database Systems, 16(3):439–475, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1991-16-3/p439-karabeg/p439-karabeg.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
3/p439-karabeg/;   http://www.acm.

                                        org/pubs/toc/Abstracts/tods/111208.
                                        html.

Yu:1991:RTR

[384] Philip S. Yu, Avraham Leff, and Yann-Hang Lee. On robust transaction routing and load sharing. ACM Transactions on Database Systems, 16(3): 476–512, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1991-16-3/p476-yu/p476-yu.
                                        pdf; http://www.acm.org/pubs/citations/
                                        journals/tods/1991-16-3/p476-yu/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/111210.html.

Negri:1991:FSS

[385] M. Negri, G. Pelagatti, and L. Sbattella. Formal semantics of SQL queries. ACM Transactions on Database Systems, 16(3):513–534, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1991-16-3/p513-negri/p513-negri.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1991-16-
                                        3/p513-negri/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/111212.
                                        html.

Roussopoulos:1991:IAM

[386] Nicholas Roussopoulos. An incremental access method for ViewCache: Concept, algorithms, and cost analysis. ACM Transactions on Database Systems, 16(3):535–563, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/

1991-16-3/p535-roussopoulos/p535-
roussopoulos.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1991-16-3/p535-roussopoulos/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/111215.html.

Mukkamala:1991:NEC

[387] Ravi Mukkamala and Sushil Jajodia. A note on estimating the cardinality of the projection of a database relation. ACM Transactions on Database Systems, 16(3):564–566, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1991-16-3/p564-mukkamala/p564-mukkamala.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
3/p564-mukkamala/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
111218.html.

Wolfson:1991:CMP

[388] O. Wolfson and A. Milo. Corrigendum: “The Multicast Policy and its Relationship to Replicated Data Placement” [ACM Trans. Database Systems 16 (1991), no. 1, 181–205, by O. Wolfson and A. Milo] (MR 92a:68042). ACM Transactions on Database Systems, 16 (3):567, September 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [375].

Hernandez:1991:CTM

[389] Héctor J. Hernández and Edward P. F. Chan. Constant-time-maintainable BCNF database schemes. ACM Transactions on Database Systems, 16(4): 571–599, December 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1991-16-4/p571-hernandez/p571-
                                        hernandez.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1991-16-4/p571-hernandez/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/115301.html.

Hou:1991:SEA

[390] Wen-Chi Hou and Gultekin Özsoyoğlu. Statistical estimators for aggregate relational algebra queries. ACM Transactions on Database Systems, 16(4): 600–654, December 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1991-16-4/p600-hou/p600-hou.
                                        pdf; http://www.acm.org/pubs/citations/
                                        journals/tods/1991-16-4/p600-hou/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/115300.html.

Negri:1991:DJN

[391] M. Negri and G. Pelagatti. Distributive join: a new algorithm for joining relations. ACM Transactions on Database Systems, 16(4):655–669, December 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1991-16-4/p655-negri/p655-negri.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1991-16-
                                        4/p655-negri/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/115299.
                                        html.

Moerkotte:1991:RCC

[392] Guido Moerkotte and Peter C. Lockemann. Reactive consistency control in deductive databases. ACM Transactions on Database Systems, 16(4): 670–702, December 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1991-16-4/p670-moerkotte/p670-
moerkotte.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1991-16-4/p670-moerkotte/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/115298.html.

Carey:1991:CDT

[393] Michael J. Carey and Miron Livny. Conflict detection tradeoffs for replicated data. ACM Transactions on Database Systems, 16(4):703–746, December 1991. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1991-16-4/p703-carey/p703-carey.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1991-16-
4/p703-carey/;   http://www.acm.
org/pubs/toc/Abstracts/tods/115289.
html.

Cattell:1992:OOB

[394] R. G. G. Cattell and J. Skeen. Object operations benchmark. ACM Transactions on Database Systems, 17(1):1– 31, March 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-1/p1-cattell/p1-cattell.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
1/p1-cattell/;   http://www.acm.
org/pubs/toc/Abstracts/tods/128766.
html.

Weddell:1992:RAF

[395] Grant E. Weddell. Reasoning about functional dependencies generalized for semantic data models. ACM Transactions on Database Systems, 17(1):32– 64, March 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1992-17-1/p32-weddell/p32-
                                        weddell.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1992-
                                        17-1/p32-weddell/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        128767.html.

Paredaens:1992:CNA

[396] Jan Paredaens and Dirk Van Gucht. Converting nested algebra expressions into flat algebra expressions. ACM Transactions on Database Systems, 17(1):65–93, March 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1992-17-1/p65-paredaens/p65-
                                        paredaens.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1992-17-1/p65-paredaens/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/128768.html.

Mohan:1992:ATR

[397] C. Mohan, Don Haderle, Bruce Lindsay, Hamid Pirahesh, and Peter Schwarz. ARIES: a transaction recovery method supporting fine-granularity locking and partial rollbacks using write-ahead logging. ACM Transactions on Database Systems, 17(1):94– 162, March 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-1/p94-mohan/p94-mohan.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1992-17-1/p94-mohan/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/128770.html. Also

published in/as: IBM Almaden Res. Ctr, Res. R. No. RJ-6649, Jan. 1989, 45 pp.

Badrinath:1992:SBC

[398] B. R. Badrinath and Krithi Ramamritham. Semantics-based concurrency control: Beyond commutativity. ACM Transactions on Database Systems, 17(1):163–199, March 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-1/p163-badrinath/p163-badrinath.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
1/p163-badrinath/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
128771.html.

Wang:1992:CTM

[399] Ke Wang and Marc H. Graham. Constant-time maintainability: a generalization of independence. ACM Transactions on Database Systems, 17(2):201–246, June 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1992-17-2/p201-wang/p201-wang.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
2/p201-wang/;   http://www.acm.
org/pubs/toc/Abstracts/tods/128904.
html.

Becker:1992:RBO

[400] Ludger Becker and Ralf Hartmut Güting. Rule-based optimization and query processing in an extensible geometric database system. ACM Transactions on Database Systems, 17(2):247–303, June 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1992-17-2/p247-becker/p247-
                                        becker.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1992-
                                        17-2/p247-becker/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        128905.html.

Franaszek:1992:CCH

[401] Peter A. Franaszek, John T. Robinson, and Alexander Thomasian. Concurrency control for high contention environments. ACM Transactions on Database Systems, 17(2):304–345, June 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1992-17-2/p304-franaszek/p304-franaszek.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1992-17-
                                        2/p304-franaszek/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        128906.html.

Leng:1992:OWA

[402] Chun-Wu Roger Leng and Dik Lun Lee. Optimal weight assignment for signature generation. ACM Transactions on Database Systems, 17(2):346– 373, June 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/

1992-17-2/p346-leng/p346-leng.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1992-17-2/p346-leng/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/128907.html.

Tansel:1992:MRH

[403] Abdullah U. Tansel and Lucy Garnett. On M. A. Roth, H. F. Korth and A. Silberschatz: “Extended Algebra and Calculus for Nested Relational Databases” [ACM Trans. Database Systems 13 (1988), no. 4, 389–417]. ACM Transactions on Database Systems, 17(2):374– 383, June 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1992-17-2/p374-tansel/p374-
tansel.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1992-
17-2/p374-tansel/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
128908.html. See [322].

Bergamaschi:1992:TRC

[404] Sonia Bergamaschi and Claudio Sartori. On taxonomic reasoning in conceptual design. ACM Transactions on Database Systems, 17(3):385–422, September 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1992-17-3/p385-bergamaschi/
p385-bergamaschi.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1992-17-3/p385-bergamaschi/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/132272.html.

Markowitz:1992:REE

[405] Victor M. Markowitz and Arie Shoshani. Representing extended entity-relationship structures in relational databases: a modular approach. ACM Transactions on Database Systems, 17(3):423–464, September 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-3/p423-markowitz/p423-markowitz.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
3/p423-markowitz/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
132273.html.

Date:1992:SCG

[406] C. J. Date and Ronald Fagin. Simple conditions for guaranteeing higher normal forms in relational databases. ACM Transactions on Database Systems, 17(3):465–476, September 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-3/p465-date/p465-date.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1992-17-3/p465-date/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/132274.html.

Hsu:1992:PEC

[407] Meichun Hsu and Bin Zhang. Performance evaluation of cautious waiting. ACM Transactions on Database Systems, 17(3):477–512, September 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1992-17-3/p477-hsu/p477-hsu.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1992-17-3/p477-hsu/

;   http://www.acm.org/pubs/toc/
Abstracts/tods/132275.html.

Abbott:1992:SRT

[408] Robert K. Abbott and Héctor García-Molina. Scheduling real-time transactions: a performance evaluation. ACM Transactions on Database Systems, 17(3):513–560, September 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1992-17-3/p513-abbott/p513-
abbott.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1992-
17-3/p513-abbott/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
132276.html.

Siegel:1992:MAR

[409] Michael Siegel, Edward Sciore, and Sharon Salveter. A method for automatic rule derivation to support semantic query optimization. ACM Transactions on Database Systems, 17(4): 563–600, December 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1992-17-4/p563-siegel/p563-
siegel.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1992-
17-4/p563-siegel/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
146932.html.

Kamel:1992:IDC

[410] Nabil Kamel and Roger King. Intelligent database caching through the use of page answers and page traces. ACM Transactions on Database Systems, 17(4):601–646, December 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1992-17-4/p601-kamel/p601-kamel.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
4/p601-kamel/;   http://www.acm.
org/pubs/toc/Abstracts/tods/146933.
html.

Maiocchi:1992:ADT

[411] Roberto Maiocchi, Barbara Pernici, and Federico Barbic. Automatic deduction of temporal information. ACM Transactions on Database Systems, 17(4):647–688, December 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-4/p647-maiocchi/p647-maiocchi.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
4/p647-maiocchi/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
146934.html.

Agrawal:1992:GTQ

[412] D. Agrawal and A. El Abbadi. The generalized tree quorum protocol: An efficient approach for managing replicated data. ACM Transactions on Database Systems, 17(4):689–717, December 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1992-17-4/p689-agrawal/p689-agrawal.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1992-17-
4/p689-agrawal/;   http://www.acm.
org/pubs/toc/Abstracts/tods/146935.
html.

Atzeni:1992:URD

[413] Paolo Atzeni and Riccardo Torlone. Updating relational databases through weak instance interfaces. ACM Transactions on Database Systems, 17(4): 718–745, December 1992. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1992-17-4/p718-atzeni/p718-
atzeni.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1992-
17-4/p718-atzeni/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
146936.html.

Ishikawa:1993:MLI

[414] Hiroshi Ishikawa, Fumio Suzuki, Fumihiko Kozakura, Akifumi Makinouchi, Mika Miyagishima, Yoshio Izumida, Masaaki Aoshima, and Yasuo Yamane. The model, language, and implementation of an objectoriented multimedia knowledge base management system. ACM Transactions on Database Systems, 18(1):1– 50, March 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-1/p1-ishikawa/p1-ishikawa.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
1/p1-ishikawa/;   http://www.acm.
org/pubs/toc/Abstracts/tods/151285.
html.

Johnson:1993:PCB

[415] Theodore Johnson and Dennis Shasha. The performance of current B-Tree algorithms. ACM Transactions on Database Systems, 18(1):51–101, March 1993. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1993-18-1/p51-johnson/p51-
johnson.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1993-
18-1/p51-johnson/;   http://www.
acm.org/pubs/toc/Abstracts/tods/

151286.html. “Current” in the title should be “Concurrent”.

Kumar:1993:CAT

[416] Akhil Kumar and Arie Segev. Cost and availability tradeoffs in replicated data concurrency control. ACM Transactions on Database Systems, 18(1):102– 131, March 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-1/p102-kumar/p102-kumar.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
1/p102-kumar/;   http://www.acm.
org/pubs/toc/Abstracts/tods/151287.
html.

Abdel-Ghaffar:1993:ODA

[417] Khaled A. S. Abdel-Ghaffar and Amr El Abbadi. Optimal disk allocation for partial match queries. ACM Transactions on Database Systems, 18(1):132– 156, March 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-1/p132-abdel-ghaffar/p132-
abdel-ghaffar.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1993-18-1/p132-abdel-ghaffar/
; http://www.acm.org/pubs/toc/Abstracts/
tods/151288.html.

Matsliach:1993:PAF

[418] Gabriel Matsliach. Performance analysis of file organizations that use multibucket data leaves with partial expansions. ACM Transactions on Database Systems, 18(1):157–180, March 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-1/p157-matsliach/p157-matsliach.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
1/p157-matsliach/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
151289.html.

Chomicki:1993:FRI

[419] Jan Chomicki and Tomasz Imieliński. Finite representation of infinite query answers. ACM Transactions on Database Systems, 18(2):181–223, June 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1993-18-2/p181-chomicki/p181-chomicki.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
2/p181-chomicki/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
151635.html.

Hou:1993:PTC

[420] Wen-Chi Hou and Gultekin Özsoyoğlu. Processing time-constrained aggregate queries in CASE-DB. ACM Transactions on Database Systems, 18(2):224– 261, June 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-2/p224-hou/p224-hou.pdf;

                                        http://www.acm.org/pubs/citations/
                                        journals/tods/1993-18-2/p224-hou/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/151636.html.

Drenick:1993:SQO

[421] P. E. Drenick and E. J. Smith. Stochastic query optimization in distributed databases. ACM Transactions on Database Systems, 18(2):262– 288, June 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1993-18-2/p262-drenick/p262-drenick.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1993-18-
                                        2/p262-drenick/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/151637.
                                        html.

Rothermel:1993:OCP

[422] Kurt Rothermel and Stefan Pappe. Open commit protocols tolerating commission failures. ACM Transactions on Database Systems, 18(2):289– 332, June 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1993-18-2/p289-rothermel/p289-rothermel.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1993-18-
                                        2/p289-rothermel/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        151638.html.

Rahm:1993:EPE

[423] Erhard Rahm. Empirical performance evaluation of concurrency and coherency control protocols for database sharing systems. ACM Transactions on Database Systems, 18(2):333– 377, June 1993. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-2/p333-rahm/p333-rahm.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1993-18-2/p333-rahm/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/151639.html.

Kuper:1993:LDM

[424] Gabriel M. Kuper and Moshe Y. Vardi. The logical data model. ACM Transactions on Database Systems, 18(3): 379–413, September 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1993-18-3/p379-kuper/p379-
kuper.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1993-
18-3/p379-kuper/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
155274.html.

Levene:1993:SNE

[425] Mark Levene and George Loizou. Semantics for null extended nested relations. ACM Transactions on Database Systems, 18(3):414–459, September 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1993-18-3/p414-levene/p414-
levene.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1993-
18-3/p414-levene/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
155275.html.

Agrawal:1993:COS

[426] Divyakant Agrawal, Amr El Abbadi, and Ambuj K. Singh. Consistency and orderability: Semantics-based correctness criteria for databases. ACM Transactions on Database Systems, 18 (3):460–486, September 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1993-18-3/p460-agrawal/p460-
agrawal.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1993-
18-3/p460-agrawal/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
155276.html.

Sagiv:1993:SQT

[427] Yehoshua Sagiv and Oded Shmueli. Solving queries by tree projections. ACM Transactions on Database Systems, 18(3):487–511, September 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1993-18-3/p487-sagiv/p487-sagiv.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
3/p487-sagiv/;   http://www.acm.
org/pubs/toc/Abstracts/tods/155277.
html.

Ioannidis:1993:TCA

[428] Yannis Ioannidis, Raghu Ramakrishnan, and Linda Winger. Transitive closure algorithms based on graph traversal. ACM Transactions on Database Systems, 18(3):512–576, September 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-3/p512-ioannidis/p512-ioannidis.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
3/p512-ioannidis/;   http://www.

acm.org/pubs/toc/Abstracts/tods/
155273.html.

Thomasian:1993:TPL

[429] Alexander Thomasian. Two-phase locking performance and its thrashing behavior. ACM Transactions on Database Systems, 18(4):579–625, December 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1993-18-4/p579-thomasian/p579-thomasian.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
4/p579-thomasian/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
169720.html.

Qian:1993:DSD

[430] Xiaolei Qian. The deductive synthesis of database transactions. ACM Transactions on Database Systems, 18 (4):626–677, December 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1993-18-4/p626-qian/p626-qian.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1993-18-
4/p626-qian/;   http://www.acm.
org/pubs/toc/Abstracts/tods/169716.
html.

Malvestuto:1993:USA

[431] Francesco M. Malvestuto. A universalscheme approach to statistical databases containing homogeneous summary tables. ACM Transactions on Database Systems, 18(4):678–708, December 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/

                                        1993-18-4/p678-malvestuto/p678-
                                        malvestuto.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1993-18-4/p678-malvestuto/;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/169712.html.

Ioannidis:1993:OHL

[432] Yannis E. Ioannidis and Stavros Christodoulakis. Optimal histograms for limiting worst-case error propagation in the size of join results. ACM Transactions on Database Systems, 18(4):709–748, December 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1993-18-4/p709-ioannidis/p709-ioannidis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1993-18-
                                        4/p709-ioannidis/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        169708.html.

Anonymous:1993:AI

[433] Anonymous. 1993 author index. ACM Transactions on Database Systems, 18 (4):749–750, December 1993. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kim:1994:CS

[434] Won Kim. Charter and scope. ACM Transactions on Database Systems, 19(1):1–??, March 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Olivier:1994:TSO

[435] Martin S. Olivier and Sebastiaan H. von Solms. A taxonomy for secure object-oriented databases. ACM Transactions on Database Systems, 19(1):3–46, March 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1994-19-1/p3-olivier/p3-olivier.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1994-19-
1/p3-olivier/;   http://www.acm.
org/pubs/toc/Abstracts/tods/174640.
html.

Tendick:1994:MRP

[436] Patrick Tendick and Norman Matloff. A modified random perturbation method for database security. ACM Transactions on Database Systems, 19(1):47–63, March 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1994-19-1/p47-tendick/p47-tendick.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1994-19-
1/p47-tendick/;   http://www.acm.
org/pubs/toc/Abstracts/tods/174641.
html.

Clifford:1994:CHR

[437] James Clifford, Albert Croker, and Alexander Tuzhilin. On completeness of historical relational query languages. ACM Transactions on Database Systems, 19(1):64– 116, March 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1994-19-1/p64-clifford/p64-
clifford.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1994-19-1/p64-clifford/;   http:

                                        //www.acm.org/pubs/toc/Abstracts/
                                        tods/174642.html.

Salem:1994:AL

[438] Kenneth Salem, Héctor García-Molina, and Jeannie Shands. Altruistic locking. ACM Transactions on Database Systems, 19(1):117–165, March 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1994-19-1/p117-salem/p117-salem.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1994-19-
                                        1/p117-salem/;   http://www.acm.
                                        org/pubs/toc/Abstracts/tods/174639.
                                        html.

Rosenthal:1994:TTR

[439] Arnon Rosenthal and David Reiner. Tools and transformations — rigorous and otherwise — for practical database design. ACM Transactions on Database Systems, 19(2):167– 211, June 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1994-19-2/p167-rosenthal/p167-rosenthal.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1994-19-
                                        2/p167-rosenthal/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        176568.html.

Bright:1994:ARS

[440] M. W. Bright, A. R. Hurson, and S. Pakzad. Automated resolution of semantic heterogeneity in multidatabases. ACM Transactions on Database Systems, 19(2):212–253, June 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1994-19-2/p212-bright/p212-
bright.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1994-
19-2/p212-bright/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
176569.html.

Sciore:1994:USV

[441] Edward Sciore, Michael Siegel, and Arnon Rosenthal. Using semantic values to facilitate interoperability among heterogeneous information systems. ACM Transactions on Database Systems, 19(2):254– 290, June 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1994-19-2/p254-sciore/p254-
sciore.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1994-
19-2/p254-sciore/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
176570.html.

Subrahmanian:1994:AKB

[442] V. S. Subrahmanian. Amalgamating knowledge bases. ACM Transactions on Database Systems, 19(2):291– 331, June 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1994-19-2/p291-subrahmanian/p291-
subrahmanian.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1994-19-2/p291-subrahmanian/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/176571.html.

Yan:1994:ISS

[443] Tak W. Yan and Héctor García-Molina. Index structures for selective dissemination of information under the Boolean model. ACM Transactions on Database Systems, 19(2):332– 364, June 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1994-19-2/p332-yan/p332-yan.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1994-19-2/p332-yan/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/176573.html.

Ceri:1994:AGP

[444] Stefano Ceri, Piero Fraternali, Stefano Paraboschi, and Letizia Tanca. Automatic generation of production rules for integrity maintenance. ACM Transactions on Database Systems, 19(3):367–422, September 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1994-19-3/p367-ceri/p367-ceri.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1994-19-3/p367-ceri/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/185828.html.

Polyzois:1994:ERB

[445] Christos A. Polyzois and Héctor García-Molina. Evaluation of remote backup algorithms for transactionprocessing systems. ACM Transactions on Database Systems, 19(3):423–449, September 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1994-19-3/p423-polyzois/p423-polyzois.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1994-19-

3/p423-polyzois/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
185836.html.

Chrysanthis:1994:SET

[446] Panos K. Chrysanthis and Krithi Ramamritham. Synthesis of extended transaction models using ACTA. ACM Transactions on Database Systems, 19(3):450–491, September 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1994-19-3/p450-chrysanthis/
p450-chrysanthis.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1994-19-3/p450-chrysanthis/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/185843.html.

Korth:1994:FAC

[447] Henry F. Korth and Greg Speegle. Formal aspects of concurrency control in long-duration transaction systems using the NT/PV model. ACM Transactions on Database Systems, 19(3):492–535, September 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1994-19-3/p492-korth/p492-korth.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1994-19-
3/p492-korth/;   http://www.acm.
org/pubs/toc/Abstracts/tods/185854.
html.

Goldman:1994:QCN

[448] Kenneth J. Goldman and Nancy Lynch. Quorum consensus in nestedtransaction systems. ACM Transactions on Database Systems, 19(4):

537–585, December 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1994-19-4/p537-goldman/p537-
goldman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1994-
19-4/p537-goldman/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
195666.html.

Krishnakumar:1994:BIT

[449] Narayanan Krishnakumar and Arthur J. Bernstein. Bounded ignorance: a technique for increasing concurrency in a replicated system. ACM Transactions on Database Systems, 19(4):586–625, December 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1994-19-4/p586-krishnakumar/p586-
krishnakumar.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1994-19-4/p586-krishnakumar/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/195670.html.

Winslett:1994:FQL

[450] Marianne Winslett, Kenneth Smith, and Xiaolei Qian. Formal query languages for secure relational databases. ACM Transactions on Database Systems, 19(4):626–662, December 1994. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/articles/
journals/tods/1994-19-4/p626-winslett/
p626-winslett.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1994-19-4/p626-winslett/;
http://www.acm.org/pubs/toc/Abstracts/
tods/195675.html.

Kim:1995:CS

[451] Won Kim. Charter and scope. ACM Transactions on Database Systems, 20(1):1–2, March 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Aiken:1995:SAT

[452] Alexander Aiken, Joseph M. Hellerstein, and Jennifer Widom. Static analysis techniques for predicting the behavior of active database rules. ACM Transactions on Database Systems, 20(1):3–41, March 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1995-20-1/p3-aiken/p3-aiken.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1995-20-1/p3-aiken/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/202107.html.

Chen:1995:DUR

[453] Weidong Chen. Declarative updates of relational databases. ACM Transactions on Database Systems, 20(1):42– 70, March 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1995-20-1/p42-chen/p42-chen.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1995-20-1/p42-chen/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/202110.html.

Jagadish:1995:IDM

[454] H. V. Jagadish. The INCINER-ATE data model. ACM Transactions on Database Systems, 20(1):71– 110, March 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1995-20-1/p71-jagadish/p71-
jagadish.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1995-20-1/p71-jagadish/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/202113.html.

Baekgaard:1995:ICN

[455] Lars Baekgaard and Leo Mark. Incremental computation of nested relational query expressions. ACM Transactions on Database Systems, 20(2):111–148, June 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1995-20-2/p111-baekgaard/p111-
baekgaard.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1995-20-2/p111-baekgaard/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/210198.html.

Chomicki:1995:ECT

[456] Jan Chomicki. Efficient checking of temporal integrity constraints using bounded history encoding. ACM Transactions on Database Systems, 20(2):149–186, June 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1995-20-2/p149-chomicki/p149-
chomicki.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1995-20-2/p149-chomicki/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/210200.html.

Graefe:1995:FAU

[457] Goetz Graefe and Richard L. Cole. Fast algorithms for universal quantification in large databases. ACM Transactions on Database Systems, 20(2):187– 236, June 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1995-20-2/p187-graefe/p187-
graefe.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1995-
20-2/p187-graefe/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
210202.html.

Chen:1995:QED

[458] Weidong Chen. Query evaluation in deductive databases with alternating fixpoint semantics. ACM Transactions on Database Systems, 20(3):239–287, September 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1995-20-3/p239-chen/p239-chen.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1995-20-3/p239-chen/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/211416.html.

Ioannidis:1995:CCQ

[459] Yannis E. Ioannidis and Raghu Ramakrishnan. Containment of conjunctive queries: Beyond relations as sets. ACM Transactions on Database Systems, 20(3):288–324, September 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1995-20-3/p288-ioannidis/p288-ioannidis.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1995-20-
3/p288-ioannidis/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
211419.html.

Shasha:1995:TCA

[460] Dennis Shasha, François Llirbat, Eric Simon, and Patrick Valduriez. Transaction chopping: Algorithms and performance studies. ACM Transactions on Database Systems, 20(3):325–363, September 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1995-20-3/p325-shasha/p325-
                                        shasha.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1995-
                                        20-3/p325-shasha/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        211427.html.

Chen:1995:EML

[461] I.-Min A. Chen, Richard Hull, and Dennis McLeod. An execution model for limited ambiguity rules and its application to derived data update. ACM Transactions on Database Systems, 20(4):365–413, December 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1995-20-4/p365-chen/p365-chen.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1995-20-4/p365-chen/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/219039.html.

Fraternali:1995:SAD

[462] Piero Fraternali and Letizia Tanca. A structured approach for the definition of the semantics of active databases. ACM Transactions on Database Systems, 20(4):414–471, December 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1995-20-4/p414-fraternali/p414-
fraternali.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1995-20-4/p414-fraternali/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/219042.html.

Srivastava:1995:SOD

[463] Divesh Srivastava, S. Sudarshan, Raghu Ramakrishnan, and Jeffrey F. Naughton. Space optimization in deductive databases. ACM Transactions on Database Systems, 20(4):472–516, December 1995. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1995-20-4/p472-srivastava/p472-
srivastava.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1995-20-4/p472-srivastava/;   http:
//www.acm.org/pubs/toc/Abstracts/
tods/219056.html.

Baralis:1996:MTA

[464] Elena Baralis, Stefano Ceri, and Stefano Paraboschi. Modularization techniques for active rules design. ACM Transactions on Database Systems, 21(1):1–29, March 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1996-21-1/p1-baralis/p1-baralis.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1996-21-
1/p1-baralis/;   http://www.acm.
org/pubs/toc/Abstracts/tods/227605.
html.

Buneman:1996:PTI

[465] Peter Buneman and Atsushi Ohori.

Polymorphism and type inference in database programming. ACM Transactions on Database Systems, 21(1):30– 76, March 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1996-21-1/p30-buneman/p30-
buneman.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1996-
21-1/p30-buneman/;   http://www.
acm.org/pubs/toc/Abstracts/tods/
227609.html.

Mok:1996:NFP

[466] Wai Yin Mok, Yiu-Kai Ng, and David W. Embley. A normal form for precisely characterizing redundancy in nested relations. ACM Transactions on Database Systems, 21(1):77– 106, March 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1996-21-1/p77-mok/p77-mok.
pdf; http://www.acm.org/pubs/citations/
journals/tods/1996-21-1/p77-mok/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/227612.html.

Mumick:1996:MC

[467] Inderpal Singh Mumick, Sheldon J. Finkelstein, Hamid Pirahesh, and Raghu Ramakrishnan. Magic conditions. ACM Transactions on Database Systems, 21(1):107–155, March 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1996-21-1/p107-mumick/p107-
mumick.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1996-
21-1/p107-mumick/;   http://www.

acm.org/pubs/toc/Abstracts/tods/
227624.html.

Liu:1996:BBS

[468] Ling Liu and Robert Meersman. The building blocks for specifying communication behavior of complex objects: An activity-driven approach. ACM Transactions on Database Systems, 21(2):157–207, June 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1996-21-2/p157-liu/p157-liu.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1996-21-2/p157-liu/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/232622.html.

Ross:1996:TRE

[469] Kenneth A. Ross. Tail recursion elimination in deductive databases. ACM Transactions on Database Systems, 21(2):208–237, June 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1996-21-2/p208-ross/p208-ross.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1996-21-
2/p208-ross/;   http://www.acm.
org/pubs/toc/Abstracts/tods/232628.
html.

Bell:1996:IDD

[470] Colin Bell, Anil Nerode, Raymond T. Ng, and V. S. Subrahmanian. Implementing deductive databases by mixed integer programming. ACM Transactions on Database Systems, 21(2):238– 269, June 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1996-21-2/p238-bell/p238-bell.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1996-21-2/p238-bell/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/232691.html.

Guo:1996:SSI

[471] Sha Guo, Wei Sun, and Mark A. Weiss. Solving satisfiability and implication problems in database systems. ACM Transactions on Database Systems, 21(2):270–293, June 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1996-21-2/p270-guo/p270-guo.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tods/1996-21-2/p270-guo/
                                        ;   http://www.acm.org/pubs/toc/
                                        Abstracts/tods/232692.html.

Ciaccia:1996:DKB

[472] Paolo Ciaccia, Paolo Tiberio, and Pavel Zezula. Declustering of key-based partitioned signature files. ACM Transactions on Database Systems, 21(3): 295–338, September 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1996-21-3/p295-ciaccia/p295-
                                        ciaccia.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1996-
                                        21-3/p295-ciaccia/;   http://www.
                                        acm.org/pubs/toc/Abstracts/tods/
                                        232755.html.

Dey:1996:PRM

[473] Debabrata Dey and Sumit Sarkar. A probabilistic relational model and algebra. ACM Transactions on Database Systems, 21(3):339–369, September 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1996-21-3/p339-dey/p339-dey.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1996-21-3/p339-dey/
;   http://www.acm.org/pubs/toc/
Abstracts/tods/232796.html.

Ghandeharizadeh:1996:HED

[474] Shahram Ghandeharizadeh, Richard Hull, and Dean Jacobs. Heraclitus: Elevating deltas to be firstclass citizens in a database programming language. ACM Transactions on Database Systems, 21(3):370–426, September 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1996-21-3/p370-ghandeharizadeh/
p370-ghandeharizadeh.pdf;   http:
//www.acm.org/pubs/citations/journals/
tods/1996-21-3/p370-ghandeharizadeh/
; http://www.acm.org/pubs/toc/Abstracts/
tods/232801.html.

Kuo:1996:MVD

[475] Dean Kuo. Model and verification of a data manager based on ARIES. ACM Transactions on Database Systems, 21(4):427–479, December 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1996-21-4/p427-kuo/p427-kuo.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1996-21-4/p427-kuo/

.

Litwin:1996:LSD

[476] Witold Litwin, Marie-Anna Neimat, and Donovan A. Schneider. LH* — a scalable, distributed data structure. ACM Transactions on Database Systems, 21(4):480–525, December 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1996-21-4/p480-litwin/p480-
                                        litwin.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1996-
                                        21-4/p480-litwin/.

Raschid:1996:SUR

[477] Louiqa Raschid and Jorge Lobo. Semantics for update rule programs and implementation in a relational database management system. ACM Transactions on Database Systems, 21(4):526–572, December 1996. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1996-21-4/p526-raschid/p526-raschid.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1996-21-
                                        4/p526-raschid/.

Keen:1997:EEL

[478] John S. Keen and William J. Dally. Extended ephemeral logging: log storage management for applications with long lived transactions. ACM Transactions on Database Systems, 22(1):1– 42, March 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1997-22-1/p1-keen/p1-keen.
                                        pdf;   http://www.acm.org/pubs/

citations/journals/tods/1997-22-
1/p1-keen/.

Galindo-Legaria:1997:OSR

[479] César Galindo-Legaria and Arnon Rosenthal. Outerjoin simplification and reordering for query optimization. ACM Transactions on Database Systems, 22(1):43–74, March 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1997-22-1/p43-galindo-legaria/p43-
galindo-legaria.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1997-22-1/p43-galindo-legaria/

.

Peters:1997:AMD

[480] Randel J. Peters and M. Tamer Özsu. An axiomatic model of dynamic schema evolution in objectbase systems. ACM Transactions on Database Systems, 22(1):75–114, March 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-1/p75-peters/p75-peters.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1997-22-
1/p75-peters/.

Wang:1997:LDT

[481] X. Sean Wang, Claudio Bettini, Alexander Brodsky, and Sushil Jajodia. Logical design for temporal databases with multiple granularities. ACM Transactions on Database Systems, 22(2):115–170, June 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        1997-22-2/p115-wang/p115-wang.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/1997-22-2/p115-wang/

.

Clifford:1997:SND

[482] James Clifford, Curtis Dyreson, Tomás Isakowitz, Christian S. Jensen, and Richard T. Snodgrass. On the semantics of “now” in databases. ACM Transactions on Database Systems, 22(2):171–214, June 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        tods/1997-22-2/p171-clifford/p171-
                                        clifford.pdf;   http://www.acm.
                                        org/pubs/citations/journals/tods/
                                        1997-22-2/p171-clifford/.

Ammann:1997:AFM

[483] Paul Ammann, Sushil Jajodia, and Indrakshi Ray. Applying formal methods to semantic-based decomposition of transactions. ACM Transactions on Database Systems, 22(2):215– 254, June 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1997-22-2/p215-ammann/p215-
                                        ammann.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1997-
                                        22-2/p215-ammann/.

Wolfson:1997:ADR

[484] Ouri Wolfson, Sushil Jajodia, and Yixiu Huang. An adaptive data replication algorithm. ACM Transactions on Database Systems, 22(2):255– 314, June 1997. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-2/p255-wolfson/p255-wolfson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1997-22-
2/p255-wolfson/.

Franklin:1997:TCS

[485] Michael J. Franklin, Michael J. Carey, and Miron Livny. Transactional clientserver cache consistency: alternatives and performance. ACM Transactions on Database Systems, 22(3):315–363, September 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-3/p315-franklin/p315-franklin.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1997-22-
3/p315-franklin/;   http://www.
acm.org:80/pubs/citations/journals/
tods/1997-22-3/p315-franklin/.

Eiter:1997:DD

[486] Thomas Eiter, Georg Gottlob, and Heikki Mannila. Disjunctive Datalog. ACM Transactions on Database Systems, 22(3):364–418, September 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1997-22-3/p364-eiter/p364-eiter.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1997-22-
3/p364-eiter/;   http://www.acm.
org:80/pubs/citations/journals/
tods/1997-22-3/p364-eiter/.

Lakshmanan:1997:PFP

[487] Laks V. S. Lakshmanan, Nicola Leone, Robert Ross, and V. S. Subrahmanian. ProbView: a flexible probabilistic database system. ACM Transactions on Database Systems, 22(3):419–469, September 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-3/p419-lakshmanan/p419-
lakshmanan.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
1997-22-3/p419-lakshmanan/;   http:
//www.acm.org:80/pubs/citations/
journals/tods/1997-22-3/p419-lakshmanan/

.

Storey:1997:DDC

[488] Veda C. Storey, Roger H. L. Chiang, Debabrata Dey, Robert C. Goldstein, and Shankar Sundaresan. Database design with common sense business reasoning and learning. ACM Transactions on Database Systems, 22(4):471–512, December 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1997-22-4/p471-storey/p471-
storey.pdf;   http://www.acm.org/
pubs/citations/journals/tods/1997-
22-4/p471-storey/;   http://www.
acm.org:80/pubs/citations/journals/
tods/1997-22-4/p471-storey/.

Tari:1997:ONF

[489] Zahir Tari, John Stokes, and Stefano Spaccapietra. Object normal forms and dependency constraints for objectoriented schemata. ACM Transactions on Database Systems, 22(4):513–569, December 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-4/p513-tari/p513-tari.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/1997-22-4/p513-tari/

.

Zaharioudakis:1997:AFG

[490] Markos Zaharioudakis, Michael J. Carey, and Michael J. Franklin. Adaptive, fine-grained sharing in a clientserver OODBMS: a callback-based approach. ACM Transactions on Database Systems, 22(4):570–627, December 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1997-22-4/p570-zaharioudakis/p570-
zaharioudakis.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1997-22-4/p570-zaharioudakis/

.

Anonymous:1997:AI

[491] Anonymous. 1997 author index. ACM Transactions on Database Systems, 22 (4):628–??, ???? 1997. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Dyreson:1998:SVT

[492] Curtis E. Dyreson and Richard T. Snodgrass. Supporting valid-time indeterminacy. ACM Transactions on Database Systems, 23(1):1–57, March 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1998-23-1/p1-dyreson/p1-dyreson.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1998-23-
1/p1-dyreson/;   http://www.acm.

                                        org:80/pubs/citations/journals/
                                        tods/1998-23-1/p1-dyreson/.

Revesz:1998:SQL

[493] Peter Z. Revesz. Safe query languages for constraint databases. ACM Transactions on Database Systems, 23(1):58– 99, March 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1998-23-1/p58-revesz/p58-revesz.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1998-23-
                                        1/p58-revesz/;   http://www.acm.
                                        org:80/pubs/citations/journals/
                                        tods/1998-23-1/p58-revesz/.

Stolboushkin:1998:SSD

[494] Alexei P. Stolboushkin and Michael A. Taitslin. Safe stratified datalog with integer order does not have syntax. ACM Transactions on Database Systems, 23(1):100–109, March 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1998-23-1/p100-stolboushkin/p100-
                                        stolboushkin.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tods/1998-23-1/p100-stolboushkin/

.

Anonymous:1998:TR

[495] Anonymous. TODS referees. ACM Transactions on Database Systems, 23 (1):110–111, March 1998.

Anonymous:1998:C

[496] Anonymous. Corrigenda. ACM Transactions on Database Systems, 23(1): 112, March 1998.

Hellerstein:1998:OTQ

[497] Joseph M. Hellerstein. Optimization techniques for queries with expensive methods. ACM Transactions on Database Systems, 23(2):113– 157, June 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/1998-23-2/p113-hellerstein/
p113-hellerstein.pdf;   http://
www.acm.org/pubs/citations/journals/
tods/1998-23-2/p113-hellerstein/
;   http://www.acm.org:80/pubs/
citations/journals/tods/1998-23-
2/p113-hellerstein/.

Liu:1998:MAP

[498] Xiangning Liu, Abdelsalam Helal, and Weimin Du. Multiview access protocols for large-scale replication. ACM Transactions on Database Systems, 23(2):158–198, June 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/tods/
1998-23-2/p158-liu/p158-liu.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1998-23-2/p158-liu/
;   http://www.acm.org:80/pubs/
citations/journals/tods/1998-23-
2/p158-liu/.

Mehrotra:1998:ECM

[499] Sharad Mehrotra, Rajeev Rastogi, Henry F. Korth, and Abraham Silberschatz. Ensuring consistency in multidatabases by preserving twolevel serializability. ACM Transactions on Database Systems, 23(2):199– 230, June 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1998-23-2/p199-mehrotra/p199-mehrotra.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1998-23-
                                        2/p199-mehrotra/;   http://www.
                                        acm.org:80/pubs/citations/journals/
                                        tods/1998-23-2/p199-mehrotra/.

Bertino:1998:ACM

[500] Elisa Bertino, Claudio Bettini, Elena Ferrari, and Pierangela Samarati. An access control model supporting periodicity constraints and temporal reasoning. ACM Transactions on Database Systems, 23(3):231–285, September 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        tods/1998-23-3/p231-bertino/; http:
                                        //www.acm.org:80/pubs/citations/
                                        journals/tods/1998-23-3/p231-bertino/

.

Castano:1998:CSA

[501] S. Castano, V. De Antonellis, M. G. Fugini, and B. Pernici. Conceptual schema analysis: techniques and applications. ACM Transactions on Database Systems, 23(3):286–333, September 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1998-23-3/p286-castano/p286-castano.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1998-23-
                                        3/p286-castano/;   http://www.acm.
                                        org:80/pubs/citations/journals/
                                        tods/1998-23-3/p286-castano/.

Formica:1998:EMC

[502] A. Formica, H. D. Groger, and M. Missikoff. An efficient method for checking object-oriented database schema correctness. ACM Transactions on Database Systems, 23(3):334–369, September 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1998-23-3/p334-formica/p334-formica.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/1998-23-
3/p334-formica/;   http://www.acm.
org:80/pubs/citations/journals/
tods/1998-23-3/p334-formica/.

Konopnicki:1998:IGW

[503] David Konopnicki and Oded Shmueli. Information gathering in the World-Wide Web: the W3QL query language and the W3QS system. ACM Transactions on Database Systems, 23(4): 369–410, December 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/
tods/1998-23-4/p369-konopnicki/
p369-konopnicki.pdf;   http://www.
acm.org/pubs/citations/journals/
tods/1998-23-4/p369-konopnicki/
;   http://www.acm.org:80/pubs/
citations/journals/tods/1998-23-
4/p369-konopnicki/.

Sistla:1998:TTC

[504] A. Prasad Sistla, Ouri Wolfson, Yelena Yesha, and Robert Sloan. Towards a theory of cost management for digital libraries and electronic commerce. ACM Transactions on Database Systems, 23(4):411–452, December 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        tods/1998-23-4/p411-sistla/p411-
                                        sistla.pdf;   http://www.acm.org/
                                        pubs/citations/journals/tods/1998-
                                        23-4/p411-sistla/;   http://www.
                                        acm.org:80/pubs/citations/journals/
                                        tods/1998-23-4/p411-sistla/.

Zobel:1998:IFV

[505] Justin Zobel, Alistair Moffat, and Kotagiri Ramamohanarao. Inverted files versus signature files for text indexing. ACM Transactions on Database Systems, 23(4):453–490, December 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        1998-23-4/p453-zobel/p453-zobel.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/1998-23-
                                        4/p453-zobel/;   http://www.acm.
                                        org:80/pubs/citations/journals/
                                        tods/1998-23-4/p453-zobel/.

Datta:1999:BPS

[506] Anindya Datta, Debra E. Vandermeer, Aslihan Celik, and Vijay Kumar. Broadcast protocols to support efficient retrieval from databases by mobile users. ACM Transactions on Database Systems, 24(1):1– 79, March 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        tods/1999-24-1/p1-datta/.

Levene:1999:DDI

[507] Mark Levene and George Loizou. Database design for incomplete relations. ACM Transactions on Database Systems, 24(1):80–126, March 1999.

CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/citations/
journals/tods/1999-24-1/p80-levene/

.

Wijsen:1999:TFC

[508] Jef Wijsen. Temporal FDs on complex objects. ACM Transactions on Database Systems, 24(1):127–176, March 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-1/p127-wijsen/.

Chaudhuri:1999:OQU

[509] Surajit Chaudhuri and Kyuseok Shim. Optimization of queries with userdefined predicates. ACM Transactions on Database Systems, 24(2):177– 228, June 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-2/p177-chaudhuri/.

Gravano:1999:GTS

[510] Luis Gravano, Héctor García-Molina, and Anthony Tomasic. GlOSS: text-source discovery over the Internet. ACM Transactions on Database Systems, 24(2):229–264, June 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-2/p229-gravano/.

Hjaltason:1999:DBS

[511] Gísli R. Hjaltason and Hanan Samet. Distance browsing in spatial databases. ACM Transactions on Database Systems, 24(2):265–318, June 1999. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/citations/
journals/tods/1999-24-2/p265-hjaltason/

.

Alagic:1999:TCO

[512] Suad Alagíc. Type-checking OQL queries in the ODMG type systems. ACM Transactions on Database Systems, 24(3):319–360, September 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-3/p319-alagic/.

Bozkaya:1999:ILM

[513] Tolga Bozkaya and Meral Ozsoyoglu. Indexing large metric spaces for similarity search queries. ACM Transactions on Database Systems, 24(3): 361–404, September 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/citations/journals/
tods/1999-24-3/p361-bozkaya/.

Casati:1999:SIE

[514] Fabio Casati, Stefano Ceri, Stefano Paraboschi, and Guiseppe Pozzi. Specification and implementation of exceptions in workflow management systems. ACM Transactions on Database Systems, 24(3):405–451, September 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-3/p405-casati/.

Dey:1999:IDD

[515] Debabrata Dey, Veda C. Storey, and Terence M. Barron. Improving database design through the analysis of relationships. ACM Transactions on Database Systems, 24(4):453–486, December 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
1999-24-4/p453-dey/p453-dey.pdf;
http://www.acm.org/pubs/citations/
journals/tods/1999-24-4/p453-dey/

.

Muralidhar:1999:SRD

[516] Krishnamurty Muralidhar and Rathindra Sarathy. Security of random data perturbation methods. ACM Transactions on Database Systems, 24(4): 487–493, December 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/citations/journals/
tods/1999-24-4/p487-muralidhar/.

Wand:1999:OAR

[517] Yair Wand, Veda C. Storey, and Ron Weber. An ontological analysis of the relationship construct in conceptual modeling. ACM Transactions on Database Systems, 24(4):494–528, December 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-4/p494-wand/.

Yan:1999:SID

[518] Tak W. Yan and Hector Garcia-Molina. The SIFT information dissemination system. ACM Transactions on Database Systems, 24(4):529–565, December 1999. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/1999-24-4/p529-yan/.

Morris:19xx:DON

[519] K. Morris, J. D. Ullman, and A. Van-Gelder. Design overview of the NAIL! system. ACM Transactions on Database Systems, ??(??):??, ???? 19xx. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Also published in/as: Proc. of International Conference on Logic Programming, BCS 3, 1986. Also published in/as: Stanford Un., CSD, TR-CS-86-1108.

Spiegler:19xx:DRA

[520] I. Spiegler and Y. Noff. Dynamic recovery as an alternative to data base restoration. ACM Transactions on Database Systems, ??(??):??, ???? 19xx. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). Submitted, March 1987.

Guting:2000:FRQ

[521] Ralf Hartmut Güting, Michael H. Böhlen, Martin Erwig, Christian S. Jensen, Nikos A. Lorentzos, Markus Schneider, and Michalis Vazirgiannis. A foundation for representing and querying moving objects. ACM Transactions on Database Systems, 25(1):1– 42, March 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
2000-25-1/p1-guting/p1-guting.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/2000-25-1/p1-guting/

.

Kossmann:2000:IDP

[522] Donald Kossmann and Konrad Stocker. Iterative dynamic programming: a new class of query optimization algorithms. ACM Transactions on Database Systems, 25(1):43–82, March 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/2000-25-1/p43-kossmann/p43-
kossmann.pdf;   http://www.acm.
org/pubs/citations/journals/tods/
2000-25-1/p43-kossmann/.

Lerner:2000:MCT

[523] Barbara Staudt Lerner. A model for compound type changes encountered in schema evolution. ACM Transactions on Database Systems, 25(1):83– 127, March 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
2000-25-1/p83-lerner/p83-lerner.
pdf;   http://www.acm.org/pubs/
citations/journals/tods/2000-25-
1/p83-lerner/.

Bohm:2000:CMQ

[524] Christian Böhm. A cost model for query processing in high dimensional data spaces. ACM Transactions on Database Systems, 25(2):129– 178, June 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/
2000-25-2/p129-bohm/p129-bohm.pdf;
 http://www.acm.org/pubs/citations/
journals/tods/2000-25-2/p129-bohm/

.

Cui:2000:TLV

[525] Yingwei Cui, Jennifer Widom, and Janet L. Wiener. Tracing the lineage of view data in a warehousing environment. ACM Transactions on Database Systems, 25(2):179– 227, June 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/tods/
                                        2000-25-2/p179-cui/p179-cui.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/tods/2000-25-2/p179-cui/

.

Parsons:2000:EIT

[526] Jeffrey Parsons and Yair Wand. Emancipating instances from the tyranny of classes in information modeling. ACM Transactions on Database Systems, 25(2):228–268, June 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        2000-25-2/p228-parsons/p228-parsons.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/2000-25-
                                        2/p228-parsons/.

Baralis:2000:AAS

[527] Elena Baralis and Jennifer Widom. An algebraic approach to static analysis of active database rules. ACM Transactions on Database Systems, 25(3): 269–332, September 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        tods/2000-25-3/p269-baralis/.

Kemme:2000:NAD

[528] Bettina Kemme and Gustavo Alonso.

A new approach to developing and implementing eager database replication protocols. ACM Transactions on Database Systems, 25(3):333–379, September 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/2000-25-3/p333-kemme/.

Meo:2000:TDV

[529] Rosa Meo. Theory of dependence values. ACM Transactions on Database Systems, 25(3):380–406, September 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/citations/journals/
tods/2000-25-3/p380-meo/.

Bohlen:2000:TSM

[530] Michael H. Böhlen, Christian S. Jensen, and Richard T. Snodgrass. Temporal statement modifiers. ACM Transactions on Database Systems, 25 (4):407–456, 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/
tods/2000-25-4/p407-bohlen/p407-
bohlen.pdf;   http://www.acm.org/
pubs/citations/journals/tods/2000-
25-4/p407-bohlen/.

Fegaras:2000:OOQ

[531] Leonidas Fegaras and David Maier. Optimizing object queries using an effective calculus. ACM Transactions on Database Systems, 25(4): 457–516, 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.

org/pubs/articles/journals/tods/

                                        2000-25-4/p457-fegaras/p457-fegaras.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/2000-25-
                                        4/p457-fegaras/.

Kossmann:2000:CII

[532] Donald Kossmann, Michael J. Franklin, Gerhard Drasch, and Wig Ag. Cache investment: integrating query optimization and distributed data placement. ACM Transactions on Database Systems, 25(4):517–558, 2000. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/tods/
                                        2000-25-4/p517-kossmann/p517-kossmann.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/tods/2000-25-
                                        4/p517-kossmann/.

Andries:2001:AUM

[533] Marc Andries, Luca Cabibbo, Jan Paredaens, and Jan van den Bussche. Applying an update method to a set of receivers. ACM Transactions on Database Systems, 26(1):1–40, March 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Dekhtyar:2001:PTD

[534] Alex Dekhtyar, Robert Ross, and V. S. Subrahmanian. Probabilistic temporal databases, I: algebra. ACM Transactions on Database Systems, 26(1):41– 95, March 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hsu:2001:RBP

[535] Windsor W. Hsu, Alan Jay Smith, and Honesty C. Young. I/O reference behavior of production database workloads and the TPC benchmarks — an analysis at the logical level. ACM Transactions on Database Systems, 26 (1):96–143, March 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chomicki:2001:QAD

[536] Jan Chomicki, David Toman, and Michael H. Böhlen. Querying ATSQL databases with temporal logic. ACM Transactions on Database Systems, 26 (2):145–178, June 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fukuda:2001:DMO

[537] Takeshi Fukuda, Yasuhiko Morimoto, Shimichi Morishita, and Takeshi Tokuyama. Data mining with optimized two-dimensional association rules. ACM Transactions on Database Systems, 26(2):179–213, June 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jajodia:2001:FSM

[538] Sushil Jajodia, Pierangela Samarati, Maria Luisa Sapino, and V. S. Subrahmanian. Flexible support for multiple access control policies. ACM Transactions on Database Systems, 26 (2):214–260, June 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Eiter:2001:POB

[539] Thomas Eiter, James J. Lu, Thomas Lukasiewicz, and V. S. Subrahmanian. Probabilistic object bases. ACM Transactions on Database Systems, 26(3): 264–312, September 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2001:AQU

[540] Chen Li and Edward Chang. Answering queries with useful bindings. ACM Transactions on Database Systems, 26(3):313–343, September 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ng:2001:ERD

[541] Wilfred Ng. An extension of the relational data model to incorporate ordered domains. ACM Transactions on Database Systems, 26(3):344–383, September 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kotidis:2001:CDV

[542] Yannis Kotidis and Nick Roussopoulos. A case for dynamic view management. ACM Transactions on Database Systems, 26(4):388–423, December 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mamoulis:2001:MSJ

[543] Nikos Mamoulis and Dimitris Papadias. Multiway spatial joins. ACM Transactions on Database Systems, 26 (4):424–475, December 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lakshmanan:2001:SES

[544] Laks V. S. Lakshmanan, Fereidoon Sadri, and Subbu N. Subramanian. SchemaSQL: An extension to SQL for multidatabase interoperability. ACM Transactions on Database Systems, 26 (4):476–519, December 2001. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Suciu:2002:DQE

[545] Dan Suciu. Distributed query evaluation on semistructured data. ACM Transactions on Database Systems, 27(1):1–62, March 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schuldt:2002:AIT

[546] Heiko Schuldt, Gustavo Alonso, Catriel Beeri, and Hans-Jörg Schek. Atomicity and isolation for transactional processes. ACM Transactions on Database Systems, 27(1):63–116, March 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2002:LFD

[547] Mengchi Liu, Gillian Dobbie, and Tok Wang Ling. A logical foundation for deductive object-oriented databases. ACM Transactions on Database Systems, 27(1):117–151, March 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bruno:2002:TSQ

[548] Nicolas Bruno, Surajit Chaudhuri, and Luis Gravano. Top-k selection queries over relational databases: Mapping strategies and performance evaluation. ACM Transactions on Database Systems, 27(2):153–187, June 2002. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chakrabarti:2002:LAD

[549] Kaushik Chakrabarti, Eamonn Keogh, Sharad Mehrotra, and Michael Pazzani. Locally adaptive dimensionality reduction for indexing large time series databases. ACM Transactions on Database Systems, 27(2):188–228, June 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Calders:2002:SDM

[550] Toon Calders, Raymond T. Ng, and Jef Wijsen. Searching for dependencies at multiple abstraction levels. ACM Transactions on Database Systems, 27 (3):229–260, September 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gibbons:2002:FIM

[551] Phillip B. Gibbons, Yossi Matias, and Viswanath Poosala. Fast incremental maintenance of approximate histograms. ACM Transactions on Database Systems, 27(3):261–298, September 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tao:2002:CMO

[552] Yufei Tao, Dimitris Papadias, and Jun Zhang. Cost models for overlapping and multiversion structures. ACM Transactions on Database Systems, 27 (3):299–342, September 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

May:2002:UGS

[553] Wolfgang May and Bertram Ludäscher. Understanding the global semantics of referential actions using logic rules. ACM Transactions on Database Systems, 27(4):343–397, December 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ciaccia:2002:SMS

[554] Paolo Ciaccia and Marco Patella. Searching in metric spaces with userdefined and approximate distances. ACM Transactions on Database Systems, 27(4):398–437, December 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fernandez:2002:SFP

[555] Mary Fernández, Yana Kadiyska, Dan Suciu, Atsuyuki Morishima, and Wang-Chiew Tan. SilkRoute: a framework for publishing relational data in XML. ACM Transactions on Database Systems, 27(4):438–493, December 2002. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Beneventano:2003:DLS

[556] Domenico Beneventano, Sonia Bergamaschi, and Claudio Sartori. Description logics for semantic query optimization in object-oriented database systems. ACM Transactions on Database Systems, 28(1):1–50, March 2003. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Karp:2003:SAF

[557] Richard M. Karp, Scott Shenker, and Christos H. Papadimitriou. A simple algorithm for finding frequent elements in streams and bags. ACM Transactions on Database Systems, 28 (1):51–55, March 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Melnik:2003:AAS

[558] Sergey Melnik and Hector Garcia-Molina. Adaptive algorithms for set containment joins. ACM Transactions on Database Systems, 28(1):56– 99, March 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tao:2003:SQD

[559] Yufei Tao and Dimitris Papadias. Spatial queries in dynamic environments. ACM Transactions on Database Systems, 28(2):101–139, June 2003. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gunopulos:2003:DAM

[560] Dimitrios Gunopulos, Roni Khardon, Heikki Mannila, Sanjeev Saluja, Hannu Toivonen, and Ram Sewak Sharma. Discovering all most specific sentences. ACM Transactions on Database Systems, 28(2):140–174, June 2003. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lechtenborger:2003:CRV

[561] Jens Lechtenbörger and Gottfried Vossen. On the computation of relational view complements. ACM Transactions on Database Systems, 28(2):175–208, June 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wijesekera:2003:RPF

[562] Duminda Wijesekera, Sushil Jajodia, Francesco Parisi-Presicce, and Åsa Hagström. Removing permissions in the flexible authorization framework. ACM Transactions on Database Systems, 28(3):209–229, September 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jacox:2003:ISJ

[563] Edwin H. Jacox and Hanan Samet. Iterative spatial join. ACM Transactions on Database Systems, 28(3): 230–256, September 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jimenez-Peris:2003:QAD

[564] Ricardo Jiménez-Peris, M. Patiño-Martínez, Gustavo Alonso, and Bettina Kemme. Are quorums an alternative for data replication? ACM Transactions on Database Systems, 28 (3):257–294, September 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tao:2003:APS

[565] Yufei Tao, Jimeng Sun, and Dimitris Papadias. Analysis of predictive spatio-temporal queries. ACM Transactions on Database Systems, 28(4): 295–336, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lakshmanan:2003:EDM

[566] Laks V. S. Lakshmanan, Carson Kai-Sang Leung, and Raymond T. Ng. Efficient dynamic mining of constrained frequent sets. ACM Transactions on Database Systems, 28(4):337–389, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cho:2003:EPR

[567] Junghoo Cho and Hector Garcia-Molina. Effective page refresh policies for Web crawlers. ACM Transactions on Database Systems, 28(4):

390–426, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chomicki:2003:PFR

[568] Jan Chomicki. Preference formulas in relational queries. ACM Transactions on Database Systems, 28(4): 427–466, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Diao:2003:PSP

[569] Yanlei Diao, Mehmet Altinel, Michael J. Franklin, Hao Zhang, and Peter Fischer. Path sharing and predicate evaluation for high-performance XML filtering. ACM Transactions on Database Systems, 28(4):467–516, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hjaltason:2003:IDS

[570] Gisli R. Hjaltason and Hanan Samet. Index-driven similarity search in metric spaces. ACM Transactions on Database Systems, 28(4):517–580, December 2003. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kolaitis:2004:F

[571] Phokion Kolaitis and Michael J. Franklin. Foreword. ACM Transactions on Database Systems, 29(1):1, March 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Buneman:2004:ASD

[572] Peter Buneman, Sanjeev Khanna, Keishi Tajima, and Wang-Chiew Tan.

Archiving scientific data. ACM Transactions on Database Systems, 29(1):2– 42, March 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Garofalakis:2004:PWS

[573] Minos Garofalakis and Phillip B. Gibbons. Probabilistic wavelet synopses. ACM Transactions on Database Systems, 29(1):43–90, March 2004. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Grust:2004:AXE

[574] Torsten Grust, Maurice Van Keulen, and Jens Teubner. Accelerating XPath evaluation in any RDBMS. ACM Transactions on Database Systems, 29 (1):91–131, March 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ross:2004:SCM

[575] Kenneth A. Ross. Selection conditions in main memory. ACM Transactions on Database Systems, 29(1):132–161, March 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arasu:2004:CMR

[576] Arvind Arasu, Brian Babcock, Shivnath Babu, Jon McAlister, and Jennifer Widom. Characterizing memory requirements for queries over continuous data streams. ACM Transactions on Database Systems, 29(1):162–194, March 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2004:NFX

[577] Marcelo Arenas and Leonid Libkin. A normal form for XML documents. ACM Transactions on Database Systems, 29(1):195–232, March 2004. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Amer-Yahia:2004:DAO

[578] Sihem Amer-Yahia and Sophie Cluet. A declarative approach to optimize bulk loading into databases. ACM Transactions on Database Systems, 29 (2):233–281, June 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sadri:2004:EOS

[579] Reza Sadri, Carlo Zaniolo, Amir Zarkesh, and Jafar Adibi. Expressing and optimizing sequence queries in database systems. ACM Transactions on Database Systems, 29(2):282–318, June 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Marian:2004:ETQ

[580] Amélie Marian, Nicolas Bruno, and Luis Gravano. Evaluating top-k queries over Web-accessible databases. ACM Transactions on Database Systems, 29 (2):319–362, June 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yu:2004:CAM

[581] Ting Yu, Divesh Srivastava, Laks V. S. Lakshmanan, and H. V. Jagadish. A compressed accessibility map for XML. ACM Transactions on Database Systems, 29(2):363–402, June 2004.

CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Datta:2004:PBA

[582] Anindya Datta, Kaushik Dutta, Helen Thomas, Debra Vandermeer, and Krithi Ramamritham. Proxy-based acceleration of dynamically generated content on the World Wide Web: An approach and implementation. ACM Transactions on Database Systems, 29 (2):403–443, June 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Vincent:2004:SFD

[583] Millist W. Vincent, Jixue Liu, and Chengfei Liu. Strong functional dependencies and their application to normal forms in XML. ACM Transactions on Database Systems, 29(3): 445–462, September 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Trajcevski:2004:MUM

[584] Goce Trajcevski, Ouri Wolfson, Klaus Hinrichs, and Sam Chamberlain. Managing uncertainty in moving objects databases. ACM Transactions on Database Systems, 29(3):463–507, September 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chaudhuri:2004:EPM

[585] Surajit Chaudhuri, Vivek Narasayya, and Sunita Sarawagi. Extracting predicates from mining models for efficient query evaluation. ACM Transactions on Database Systems, 29(3): 508–544, September 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Babu:2004:ECR

[586] Shivnath Babu, Utkarsh Srivastava, and Jennifer Widom. Exploiting k-constraints to reduce memory overhead in continuous queries over data streams. ACM Transactions on Database Systems, 29(3):545–580, September 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2004:QWM

[587] Gültekin Özsoyoğlu, Ismail Sengör Altingövde, Abdullah Al-Hamdani, Selma Ayşe Özel, Özgür Ulusoy, and Zehra Meral özsoyoğlu. Querying Web metadata: Native score management and text support in databases. ACM Transactions on Database Systems, 29 (4):581–634, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Baralis:2004:ECR

[588] Elena Baralis and Silvia Chiusano. Essential classification rule sets. ACM Transactions on Database Systems, 29 (4):635–674, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chen:2004:MBV

[589] Songting Chen, Bin Liu, and Elke A. Rundensteiner. Multiversion-based view maintenance over distributed data sources. ACM Transactions on Database Systems, 29(4):675–709, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Balmin:2004:IVX

[590] Andrey Balmin, Yannis Papakonstantinou, and Victor Vianu. Incremental validation of XML documents. ACM Transactions on Database Systems, 29 (4):710–751, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Green:2004:PXS

[591] Todd J. Green, Ashish Gupta, Gerome Miklau, Makoto Onizuka, and Dan Suciu. Processing XML streams with deterministic automata and stream indexes. ACM Transactions on Database Systems, 29(4):752–788, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Samet:2004:DPG

[592] Hanan Samet. Decoupling partitioning and grouping: Overcoming shortcomings of spatial indexing with bucketing. ACM Transactions on Database Systems, 29(4):789–830, December 2004. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Milo:2005:EIX

[593] Tova Milo, Serge Abiteboul, Bernd Amann, Omar Benjelloun, and Fred Dang Ngoc. Exchanging intensional XML data. ACM Transactions on Database Systems, 30(1):1–40, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Papadias:2005:PSC

[594] Dimitris Papadias, Yufei Tao, Greg Fu, and Bernhard Seeger. Progressive skyline computation in database systems. ACM Transactions on Database Systems, 30(1):41–82, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Witkowski:2005:ASM

[595] Andrew Witkowski, Srikanth Bellamkonda, Tolga Bozkaya, Nathan Folkert, Abhinav Gupta, John Haydu, Lei Sheng, and Sankar Subramanian. Advanced SQL modeling in RDBMS. ACM Transactions on Database Systems, 30(1):83–121, March 2005. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Madden:2005:TAQ

[596] Samuel R. Madden, Michael J. Franklin, Joseph M. Hellerstein, and Wei Hong. TinyDB: an acquisitional query processing system for sensor networks. ACM Transactions on Database Systems, 30(1):122–173, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2005:DEG

[597] Ronald Fagin, Phokion G. Kolaitis, and Lucian Popa. Data exchange: getting to the core. ACM Transactions on Database Systems, 30(1):174–210, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pu:2005:CDS

[598] Ken Q. Pu and Alberto O. Mendelzon. Concise descriptions of subsets of structured sets. ACM Transactions on Database Systems, 30(1):211–248, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cormode:2005:WHW

[599] Graham Cormode and S. Muthukrishnan. What’s hot and what’s not: tracking most frequent items dynamically. ACM Transactions on Database Systems, 30(1):249–278, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Garofalakis:2005:XSP

[600] Minos Garofalakis and Amit Kumar. XML stream processing using tree-edit distance embeddings. ACM Transactions on Database Systems, 30(1):279– 332, March 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Geerts:2005:TUB

[601] Floris Geerts, Bart Goethals, and Jan Van Den Bussche. Tight upper bounds on the number of candidate patterns. ACM Transactions on Database Systems, 30(2):333–363, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jagadish:2005:IAB

[602] H. V. Jagadish, Beng Chin Ooi, Kian-Lee Tan, Cui Yu, and Rui Zhang. iDistance: An adaptive B^{+}-tree based indexing method for nearest neighbor search. ACM Transactions on Database Systems, 30(2):364–397, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Braga:2005:XXU

[603] Daniele Braga, Alessandro Campi, and Stefano Ceri. XQBE (XQuery By Example): a visual interface to the standard XML query language. ACM Transactions on Database Systems, 30 (2):398–443, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gottlob:2005:EAP

[604] Georg Gottlob, Christoph Koch, and Reinhard Pichler. Efficient algorithms for processing XPath queries. ACM Transactions on Database Systems, 30 (2):444–491, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fekete:2005:MSI

[605] Alan Fekete, Dimitrios Liarokapis, Elizabeth O’Neil, Patrick O’Neil, and Dennis Shasha. Making snapshot isolation serializable. ACM Transactions on Database Systems, 30(2):492–528, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Papadias:2005:ANN

[606] Dimitris Papadias, Yufei Tao, Kyriakos Mouratidis, and Chun Kit Hui. Aggregate nearest neighbor queries in spatial databases. ACM Transactions on Database Systems, 30(2):529–576, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Peng:2005:XSX

[607] Feng Peng and Sudarshan S. Chawathe. XSQ: a streaming XPath engine. ACM Transactions on Database Systems, 30(2):577–623, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wyss:2005:RLM

[608] Catharine M. Wyss and Edward L. Robertson. Relational languages for metadata integration. ACM Transactions on Database Systems, 30(2):624– 660, June 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2005:GMD

[609] Rui Zhang, Panos Kalnis, Beng Chin Ooi, and Kian-Lee Tan. Generalized multidimensional data mapping and query processing. ACM Transactions on Database Systems, 30(3): 661–697, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pang:2005:IMS

[610] Chaoyi Pang, Guozhu Dong, and Kotagiri Ramamohanarao. Incremental maintenance of shortest distance and transitive closure in first-order logic and SQL. ACM Transactions on Database Systems, 30(3):698–721, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wijsen:2005:DRU

[611] Jef Wijsen. Database repairing using updates. ACM Transactions on Database Systems, 30(3):722–768, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Litwin:2005:LHA

[612] Witold Litwin, Rim Moussa, and Thomas Schwarz. LH*RS—a highlyavailable scalable distributed data structure. ACM Transactions on Database Systems, 30(3):769–811, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

He:2005:STC

[613] Zhen He, Byung Suk Lee, and Robert Snapp. Self-tuning cost modeling of user-defined functions in an objectrelational DBMS. ACM Transactions on Database Systems, 30(3):812–853, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hurtado:2005:CSI

[614] Carlos A. Hurtado, Claudio Gutierrez, and Alberto O. Mendelzon. Capturing summarizability with integrity constraints in OLAP. ACM Transactions on Database Systems, 30(3): 854–886, September 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Suciu:2005:F

[615] Dan Suciu and Gerhard Weikum. Foreword. ACM Transactions on Database Systems, 30(4):887, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Garofalakis:2005:WSG

[616] Minos Garofalakis and Amit Kumar. Wavelet synopses for general error metrics. ACM Transactions on Database Systems, 30(4):888–928, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Marx:2005:CX

[617] Maarten Marx. Conditional XPath. ACM Transactions on Database Systems, 30(4):929–959, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yan:2005:GIB

[618] Xifeng Yan, Philip S. Yu, and Jiawei Han. Graph indexing based on discriminative frequent structure analysis. ACM Transactions on Database Systems, 30(4):960–993, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2005:CSM

[619] Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Composing schema mappings: Secondorder dependencies to the rescue. ACM Transactions on Database Systems, 30 (4):994–1055, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bowman:2005:OQS

[620] Ivan T. Bowman and Kenneth Salem. Optimization of query streams using semantic prefetching. ACM Transactions on Database Systems, 30(4): 1056–1101, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kaushik:2005:SQO

[621] Raghav Kaushik, Jeffrey F. Naughton, Raghu Ramakrishnan, and Venkatesan T. Chakravarthy. Synopses for query optimization: a spacecomplexity perspective. ACM Transactions on Database Systems, 30(4):

1102–1127, December 2005. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wu:2006:OBI

[622] Kesheng Wu, Ekow J. Otoo, and Arie Shoshani. Optimizing bitmap indices with efficient compression. ACM Transactions on Database Systems, 31(1):1–38, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schneider:2006:TRB

[623] Markus Schneider and Thomas Behr. Topological relationships between complex spatial objects. ACM Transactions on Database Systems, 31(1):39– 81, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jaluta:2006:BTC

[624] Ibrahim Jaluta, Seppo Sippu, and Eljas Soisalon-Soininen. B-tree concurrency control and recovery in page-server database systems. ACM Transactions on Database Systems, 31(1):82–132, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gray:2006:CTC

[625] Jim Gray and Leslie Lamport. Consensus on transaction commit. ACM Transactions on Database Systems, 31 (1):133–160, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Guha:2006:IXD

[626] Sudipto Guha, H. V. Jagadish, Nick Koudas, Divesh Srivastava, and Ting Yu. Integrating XML data sources using approximate joins. ACM Transactions on Database Systems, 31(1):161– 207, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Abiteboul:2006:RQX

[627] Serge Abiteboul, Luc Segoufin, and Victor Vianu. Representing and querying XML with incomplete information. ACM Transactions on Database Systems, 31(1):208–254, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pelanis:2006:IPP

[628] Mindaugas Pelanis, Simonas Šaltenis, and Christian S. Jensen. Indexing the past, present, and anticipated future positions of moving objects. ACM Transactions on Database Systems, 31 (1):255–298, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rao:2006:SXD

[629] Praveen Rao and Bongki Moon. Sequencing XML data and query twigs for fast pattern matching. ACM Transactions on Database Systems, 31 (1):299–345, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

He:2006:ACS

[630] Bin He and Kevin Chen-Chuan Chang. Automatic complex schema matching across Web query interfaces: a correlation mining approach. ACM Transactions on Database Systems, 31 (1):346–395, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Guha:2006:ASA

[631] Sudipto Guha, Nick Koudas, and Kyuseok Shim. Approximation and streaming algorithms for histogram construction problems. ACM Transactions on Database Systems, 31(1):396– 438, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Qian:2006:DIM

[632] Gang Qian, Qiang Zhu, Qiang Xue, and Sakti Pramanik. Dynamic indexing for multidimensional non-ordered discrete data spaces using a datapartitioning approach. ACM Transactions on Database Systems, 31(2):439– 484, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Iwerks:2006:MNS

[633] Glenn S. Iwerks, Hanan Samet, and Kenneth P. Smith. Maintenance of Knn and spatial join queries on continuously moving points. ACM Transactions on Database Systems, 31(2):485– 536, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pentaris:2006:QOD

[634] Fragkiskos Pentaris and Yannis Ioannidis. Query optimization in distributed networks of autonomous database systems. ACM Transactions on Database Systems, 31(2):537–583, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lin:2006:SLT

[635] Xuemin Lin, Qing Liu, Yidong Yuan, Xiaofang Zhou, and Hongjun Lu. Summarizing level-two topological relations in large spatial datasets. ACM Transactions on Database Systems, 31 (2):584–630, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bright:2006:APB

[636] Laura Bright, Avigdor Gal, and Louiqa Raschid. Adaptive pull-based policies for wide area data delivery. ACM Transactions on Database Systems, 31 (2):631–671, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cohen:2006:RQA

[637] Sara Cohen, Werner Nutt, and Yehoshua Sagiv. Rewriting queries with arbitrary aggregation functions using views. ACM Transactions on Database Systems, 31(2):672–715, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kalashnikov:2006:DID

[638] Dmitri V. Kalashnikov and Sharad Mehrotra. Domain-independent data cleaning via analysis of entityrelationship graph. ACM Transactions on Database Systems, 31(2):716–767, June 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Eiter:2006:ISI

[639] Thomas Eiter and Leonid Libkin. Introduction to special ICDT section.

ACM Transactions on Database Systems, 31(3):769, September 2006. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Martens:2006:ECX

[640] Wim Martens, Frank Neven, Thomas Schwentick, and Geert Jan Bex. Expressiveness and complexity of XML Schema. ACM Transactions on Database Systems, 31(3):770–813, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Shaft:2006:TNN

[641] Uri Shaft and Raghu Ramakrishnan. Theory of nearest neighbors indexability. ACM Transactions on Database Systems, 31(3):814–838, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Braganholo:2006:PFA

[642] Vanessa P. Braganholo, Susan B. Davidson, and Carlos A. Heuser. PATAXÓ: a framework to allow updates through XML views. ACM Transactions on Database Systems, 31 (3):839–886, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Harizopoulos:2006:IIC

[643] Stavros Harizopoulos and Anastassia Ailamaki. Improving instruction cache performance in OLTP. ACM Transactions on Database Systems, 31(3): 887–920, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Shao:2006:TNV

[644] Feng Shao, Antal Novak, and Jayavel Shanmugasundaram. Triggers over nested views of relational data. ACM Transactions on Database Systems, 31 (3):921–967, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

May:2006:SQU

[645] Norman May, Sven Helmer, and Guido Moerkotte. Strategies for query unnesting in XML databases. ACM Transactions on Database Systems, 31(3): 968–1013, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Polyzotis:2006:XSX

[646] Neoklis Polyzotis and Minos Garofalakis. XSKETCH synopses for XML data graphs. ACM Transactions on Database Systems, 31(3): 1014–1063, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sugumaran:2006:RDO

[647] Vijayan Sugumaran and Veda C. Storey. The role of domain ontologies in database design: An ontology management and conceptual modeling environment. ACM Transactions on Database Systems, 31(3):1064–1094, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Metwally:2006:IES

[648] Ahmed Metwally, Divyakant Agrawal, and Amr El Abbadi. An integrated efficient solution for computing frequent and top-k elements in data streams. ACM Transactions on Database Systems, 31(3):1095–1133, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See comments in [760].

Chaudhuri:2006:PIR

[649] Surajit Chaudhuri, Gautam Das, Vagelis Hristidis, and Gerhard Weikum. Probabilistic information retrieval approach for ranking of database query results. ACM Transactions on Database Systems, 31(3):1134–1168, September 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Calders:2006:EPA

[650] Toon Calders, Laks V. S. Lakshmanan, Raymond T. Ng, and Jan Paredaens. Expressive power of an algebra for data mining. ACM Transactions on Database Systems, 31(4): 1169–1214, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Koch:2006:CNX

[651] Christoph Koch. On the complexity of nonrecursive XQuery and functional query languages on complex values. ACM Transactions on Database Systems, 31(4):1215–1256, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ilyas:2006:ARA

[652] Ihab F. Ilyas, Walid G. Aref, Ahmed K. Elmagarmid, Hicham G. Elmongui, Rahul Shah, and Jeffrey Scott Vitter. Adaptive rank-aware query optimization in relational databases. ACM Transactions on Database Systems, 31 (4):1257–1304, December 2006. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jiao:2006:MSS

[653] Yishan Jiao. Maintaining stream statistics over multiscale sliding windows. ACM Transactions on Database Systems, 31(4):1305–1334, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pei:2006:TMS

[654] Jian Pei, Yidong Yuan, Xuemin Lin, Wen Jin, Martin Ester, Qing Liu, Wei Wang, Yufei Tao, Jeffrey Xu Yu, and Qing Zhang. Towards multidimensional subspace skyline analysis. ACM Transactions on Database Systems, 31(4): 1335–1381, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jermaine:2006:SMS

[655] Christopher Jermaine, Alin Dobra, Subramanian Arumugam, Shantanu Joshi, and Abhijit Pol. The Sort-Merge-Shrink join. ACM Transactions on Database Systems, 31(4): 1382–1416, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Afrati:2006:FSS

[656] Foto Afrati and Jennifer Widom. Foreword to special section on SIGMOD/ PODS 2005. ACM Transactions on Database Systems, 31(4):1417, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yan:2006:FBS

[657] Xifeng Yan, Feida Zhu, Philip S. Yu, and Jiawei Han. Feature-based similarity search in graph structures. ACM Transactions on Database Systems, 31(4):1418–1453, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fuxman:2006:PDE

[658] Ariel Fuxman, Phokion G. Kolaitis, Renée J. Miller, and Wang-Chiew Tan. Peer data exchange. ACM Transactions on Database Systems, 31(4): 1454–1498, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheng:2006:DMM

[659] David Cheng, Ravi Kannan, Santosh Vempala, and Grant Wang. A divideand-merge methodology for clustering. ACM Transactions on Database Systems, 31(4):1499–1525, December 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Snodgrass:2007:ESV

[660] Richard T. Snodgrass. Editorial: Single- versus double-blind reviewing. ACM Transactions on Database Systems, 32(1):1:1–1:??, March 2007. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pourabbas:2007:EEJ

[661] Elaheh Pourabbas and Arie Shoshani. Efficient estimation of joint queries from multiple OLAP databases. ACM Transactions on Database Systems, 32 (1):2:1–2:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Olteanu:2007:FNS

[662] Dan Olteanu. Forward node-selecting queries over trees. ACM Transactions on Database Systems, 32(1):3:1–3:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nash:2007:CMG

[663] Alan Nash, Philip A. Bernstein, and Sergey Melnik. Composition of mappings given by embedded dependencies. ACM Transactions on Database Systems, 32(1):4:1–4:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hwang:2007:OTK

[664] Seung won Hwang and Kevin Chen chuan Chang. Optimizing top-k queries for middleware access: a unified costbased approach. ACM Transactions on Database Systems, 32(1):5:1–5:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ceri:2007:MCV

[665] Stefano Ceri, Francesco Di Giunta, and Pier Luca Lanzi. Mining constraint violations. ACM Transactions on Database Systems, 32(1):6:1–6:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jacox:2007:SJT

[666] Edwin H. Jacox and Hanan Samet. Spatial join techniques. ACM Transactions on Database Systems, 32(1):7:1– 7:??, March 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Athitsos:2007:QSE

[667] Vassilis Athitsos, Marios Hadjieleftheriou, George Kollios, and Stan Sclaroff. Query-sensitive embeddings. ACM Transactions on Database Systems, 32 (2):8:1–8:??, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chaudhuri:2007:OSS

[668] Surajit Chaudhuri, Gautam Das, and Vivek Narasayya. Optimized stratified sampling for approximate query processing. ACM Transactions on Database Systems, 32(2):9:1–9:??, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Deligiannakis:2007:EWM

[669] Antonios Deligiannakis, Minos Garofalakis, and Nick Roussopoulos. Extended wavelets for multiple measures. ACM Transactions on Database Systems, 32(2):10:1–10:??, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rusu:2007:PRN

[670] Florin Rusu and Alin Dobra. Pseudorandom number generation for sketchbased estimations. ACM Transactions on Database Systems, 32(2):11:1–11:48, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mazeika:2007:ESA

[671] Arturas Mazeika, Michael H. Böhlen, Nick Koudas, and Divesh Srivastava. Estimating the selectivity of approximate string queries. ACM Transactions on Database Systems, 32(2):12:1–12:??, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zeng:2007:CCC

[672] Zhiping Zeng, Jianyong Wang, Lizhu Zhou, and George Karypis. Out-of-core coherent closed quasi-clique mining from large dense graph databases. ACM Transactions on Database Systems, 32(2):13:1–13:??, June 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ipeirotis:2007:MMC

[673] Panagiotis G. Ipeirotis, Alexandros Ntoulas, Junghoo Cho, and Luis Gravano. Modeling and managing changes in text databases. ACM Transactions on Database Systems, 32(3):14:1–14:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tao:2007:RSM

[674] Yufei Tao, Xiaokui Xiao, and Reynold Cheng. Range search on multidimensional uncertain data. ACM Transactions on Database Systems, 32(3): 15:1–15:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sinha:2007:MRB

[675] Rishi Rakesh Sinha and Marianne Winslett. Multi-resolution bitmap indexes for scientific data. ACM Transactions on Database Systems, 32(3): 16:1–16:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chen:2007:IHJ

[676] Shimin Chen, Anastassia Ailamaki, Phillip B. Gibbons, and Todd C. Mowry. Improving hash join performance through prefetching. ACM Transactions on Database Systems, 32 (3):17:1–17:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cao:2007:SQO

[677] Bin Cao and Antonio Badia. SQL query optimization through nested relational algebra. ACM Transactions on Database Systems, 32(3):18:1–18:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mamoulis:2007:ETA

[678] Nikos Mamoulis, Man Lung Yiu, Kit Hung Cheng, and David W. Cheung. Efficient top-k aggregation of ranked inputs. ACM Transactions on Database Systems, 32(3):19:1–19:??, August 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chaudhuri:2007:IAS

[679] Surajit Chaudhuri. Introduction to ACM SIGMOD 2006 conference papers. ACM Transactions on Database Systems, 32(4):20:1–20:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ipeirotis:2007:TQO

[680] Panagiotis G. Ipeirotis, Eugene Agichtein, Pranay Jain, and Luis Gravano. Towards a query optimizer for text-centric tasks. ACM Transactions on Database Systems, 32(4): 21:1–21:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Petropoulos:2007:EIQ

[681] Michalis Petropoulos, Alin Deutsch, Yannis Papakonstantinou, and Yannis Katsis. Exporting and interactively querying Web service-accessed sources: The CLIDE System. ACM Transactions on Database Systems, 32(4): 22:1–22:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sharfman:2007:GAM

[682] Izchak Sharfman, Assaf Schuster, and Daniel Keren. A geometric approach to monitoring threshold functions over distributed data streams. ACM Transactions on Database Systems, 32(4): 23:1–23:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

VandenBussche:2007:IPS

[683] Jan Van den Bussche. Introduction to the PODS 2006 special section. ACM Transactions on Database Systems, 32(4):24:1–24:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2007:ISM

[684] Ronald Fagin. Inverting schema mappings. ACM Transactions on Database Systems, 32(4):25:1–25:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bender:2007:APM

[685] Michael A. Bender and Haodong Hu. An adaptive packed-memory array. ACM Transactions on Database Systems, 32(4):26:1–26:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ioannidis:2007:IES

[686] Yannis Ioannidis. Introduction to the EDBT 2006 special section. ACM Transactions on Database Systems, 32 (4):27:1–27:??, November 2007. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bruno:2007:PDR

[687] Nicolas Bruno and Surajit Chaudhuri. Physical design refinement: The ‘merge-reduce’ approach. ACM Transactions on Database Systems, 32(4): 28:1–28:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Skopal:2007:UFF

[688] Tomáš Skopal. Unified framework for fast exact and approximate search in dissimilarity spaces. ACM Transactions on Database Systems, 32(4): 29:1–29:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2007:NGN

[689] Yunyao Li, Huahai Yang, and H. V. Jagadish. NaLIX: a generic natural language search environment for XML data. ACM Transactions on Database Systems, 32(4):30:1–30:??, November 2007. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hristidis:2008:ABK

[690] Vagelis Hristidis, Heasoo Hwang, and Yannis Papakonstantinou. Authoritybased keyword search in databases. ACM Transactions on Database Systems, 33(1):1:1–1:40, March 2008. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Aggarwal:2008:SDM

[691] Charu C. Aggarwal and Philip S. Yu. On static and dynamic methods for condensation-based privacy-preserving data mining. ACM Transactions on Database Systems, 33(1):2:1–2:39, March 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Balazinska:2008:FTB

[692] Magdalena Balazinska, Hari Balakrishnan, Samuel R. Madden, and Michael Stonebraker. Fault-tolerance in the Borealis distributed stream processing system. ACM Transactions on Database Systems, 33(1):3:1–3:44, March 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2008:IPX

[693] Wenfei Fan and Philip Bohannon. Information preserving XML schema embedding. ACM Transactions on Database Systems, 33(1):4:1–4:44, March 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sharaf:2008:AMP

[694] Mohamed A. Sharaf, Panos K. Chrysanthis, Alexandros Labrinidis, and Kirk Pruhs. Algorithms and metrics for processing multiple heterogeneous continuous queries. ACM Transactions on Database Systems, 33 (1):5:1–5:44, March 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2008:CFD

[695] Wenfei Fan, Floris Geerts, Xibei Jia, and Anastasios Kementsietsidis. Conditional functional dependencies for capturing data inconsistencies. ACM Transactions on Database Systems, 33 (2):6:1–6:48, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jacox:2008:MSS

[696] Edwin H. Jacox and Hanan Samet. Metric space similarity joins. ACM Transactions on Database Systems, 33 (2):7:1–7:38, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

He:2008:COD

[697] Bingsheng He and Qiong Luo. Cacheoblivious databases: Limitations and opportunities. ACM Transactions on Database Systems, 33(2):8:1–8:42, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cormode:2008:ACQ

[698] Graham Cormode and Minos Garofalakis. Approximate continuous querying over distributed streams. ACM Transactions on Database Systems, 33 (2):9:1–9:??, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Eiter:2008:RLQ

[699] Thomas Eiter, Michael Fink, Gianluigi Greco, and Domenico Lembo. Repair localization for query answering from inconsistent databases. ACM Transactions on Database Systems, 33 (2):10:1–10:??, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2008:QIS

[700] Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Quasi-inverses of schema mappings. ACM Transactions on Database Systems, 33(2):11:1–11:??, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2008:CTA

[701] Donghui Zhang, Alexander Markowetz, Vassilis J. Tsotras, Dimitrios Gunopulos, and Bernhard Seeger. On computing temporal aggregates with range predicates. ACM Transactions on Database Systems, 33(2):12:1–12:??, June 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Soliman:2008:PTR

[702] Mohamed A. Soliman, Ihab F. Ilyas, and Kevin Chen-Chuan Chang. Probabilistic top-k and ranking-aggregate queries. ACM Transactions on Database Systems, 33(3):13:1–13:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ke:2008:CPM

[703] Yiping Ke, James Cheng, and Wilfred Ng. Correlated pattern mining in quantitative databases. ACM Transactions on Database Systems, 33(3): 14:1–14:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rusu:2008:SSJ

[704] Florin Rusu and Alin Dobra. Sketches for size of join estimation. ACM Transactions on Database Systems, 33(3): 15:1–15:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xu:2008:CBS

[705] Fei Xu, Christopher Jermaine, and Alin Dobra. Confidence bounds for sampling-based group by estimates. ACM Transactions on Database Systems, 33(3):16:1–16:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

LeFevre:2008:WAA

[706] Kristen LeFevre, David J. DeWitt, and Raghu Ramakrishnan. Workloadaware anonymization techniques for large-scale datasets. ACM Transactions on Database Systems, 33(3): 17:1–17:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Karras:2008:HSO

[707] Panagiotis Karras and Nikos Mamoulis. Hierarchical synopses with optimal error guarantees. ACM Transactions on Database Systems, 33(3):18:1–18:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lester:2008:EOI

[708] Nicholas Lester, Alistair Moffat, and Justin Zobel. Efficient online index construction for text databases. ACM Transactions on Database Systems, 33 (3):19:1–19:??, August 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2008:FTS

[709] Z. Meral Özsoyoğlu. Foreword to TODS SIGMOD/PODS/ICDT 2007 special issue. ACM Transactions on Database Systems, 33(4):20:1–20:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ooi:2008:IAS

[710] Beng Chin Ooi. Introduction to ACM SIGMOD 2007 special section. ACM Transactions on Database Systems, 33 (4):21:1–21:??, November 2008. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Melnik:2008:CMB

[711] Sergey Melnik, Atul Adya, and Philip A. Bernstein. Compiling mappings to bridge applications and databases. ACM Transactions on Database Systems, 33(4):22:1–22:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jermaine:2008:SAQ

[712] Chris Jermaine, Subramanian Arumugam, Abhijit Pol, and Alin Dobra. Scalable approximate query processing with the DBO engine. ACM Transactions on Database Systems, 33(4):

23:1–23:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Libkin:2008:IPS

[713] Leonid Libkin. Introduction to the PODS 2007 special section. ACM Transactions on Database Systems, 33 (4):24:1–24:??, November 2008. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2008:ECX

[714] Wenfei Fan, Floris Geerts, and Frank Neven. Expressiveness and complexity of XML publishing transducers. ACM Transactions on Database Systems, 33(4):25:1–25:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jayram:2008:ESA

[715] T. S. Jayram, Andrew McGregor, S. Muthukrishnan, and Erik Vee. Estimating statistical aggregates on probabilistic data streams. ACM Transactions on Database Systems, 33(4): 26:1–26:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schwentick:2008:IIS

[716] Thomas Schwentick and Dan Suciu. Introduction to ICDT 2007 special section. ACM Transactions on Database Systems, 33(4):27:1–27:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Buneman:2008:EIP

[717] Peter Buneman, James Cheney, and Stijn Vansummeren. On the expressiveness of implicit provenance in query and update languages. ACM Transactions on Database Systems, 33(4): 28:1–28:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ghelli:2008:CAX

[718] Giorgio Ghelli, Kristoffer Rose, and Jérôme Siméon. Commutativity analysis for XML updates. ACM Transactions on Database Systems, 33(4): 29:1–29:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pavlou:2008:FAD

[719] Kyriacos E. Pavlou and Richard T. Snodgrass. Forensic analysis of database tampering. ACM Transactions on Database Systems, 33(4): 30:1–30:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bartolini:2008:ESB

[720] Ilaria Bartolini, Paolo Ciaccia, and Marco Patella. Efficient sort-based skyline evaluation. ACM Transactions on Database Systems, 33(4): 31:1–31:??, November 2008. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mishra:2009:DQM

[721] Chaitanya Mishra and Nick Koudas. The design of a query monitoring system. ACM Transactions on Database Systems, 34(1):1:1–1:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheng:2009:EQP

[722] James Cheng, Yiping Ke, and Wilfred Ng. Efficient query processing on graph databases. ACM Transactions on Database Systems, 34(1):2:1–2:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Spiegel:2009:TSA

[723] Joshua Spiegel and Neoklis Polyzotis. TuG synopses for approximate query answering. ACM Transactions on Database Systems, 34(1):3:1–3:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kramer:2009:SIC

[724] Jürgen Krämer and Bernhard Seeger. Semantics and implementation of continuous sliding window queries over data streams. ACM Transactions on Database Systems, 34(1):4:1–4:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jain:2009:QAO

[725] Alpa Jain and Panagiotis G. Ipeirotis. A quality-aware optimizer for information extraction. ACM Transactions on Database Systems, 34(1):5:1–5:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Considine:2009:RAA

[726] Jeffrey Considine, Marios Hadjieleftheriou, Feifei Li, John Byers, and George Kollios. Robust approximate aggregation in sensor data management systems. ACM Transactions on Database Systems, 34(1):6:1–6:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Angiulli:2009:DOP

[727] Fabrizio Angiulli, Fabio Fassetti, and Luigi Palopoli. Detecting outlying properties of exceptional objects. ACM Transactions on Database Systems, 34 (1):7:1–7:??, April 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wong:2009:ABA

[728] Raymond Chi-Wing Wong, Ada Wai-Chee Fu, Ke Wang, and Jian Pei. Anonymization-based attacks in privacy-preserving data publishing. ACM Transactions on Database Systems, 34(2):8:1–8:??, June 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ghinita:2009:FED

[729] Gabriel Ghinita, Panagiotis Karras, Panos Kalnis, and Nikos Mamoulis. A framework for efficient data anonymization under privacy and accuracy constraints. ACM Transactions on Database Systems, 34(2):9:1–9:??, June 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hartmann:2009:ERA

[730] Sven Hartmann and Sebastian Link. Efficient reasoning about a robust XML key fragment. ACM Transactions on Database Systems, 34(2):10:1–10:??, June 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lin:2009:SII

[731] Yi Lin, Bettina Kemme, Ricardo Jiménez-Peris, Marta Patiño-Martínez, and José Enrique Armendáriz-Iñigo. Snapshot isolation and integrity constraints in replicated databases. ACM Transactions on Database Systems, 34 (2):11:1–11:??, June 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Su:2009:OOA

[732] Weifeng Su, Jiying Wang, and Frederick H. Lochovsky. ODE: Ontologyassisted data extraction. ACM Transactions on Database Systems, 34(2): 12:1–12:??, June 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Agarwal:2009:ISS

[733] Pankaj K. Agarwal, Junyi Xie, Jun Yang, and Hai Yu. Input-sensitive scalable continuous join query processing. ACM Transactions on Database Systems, 34(3):13:1–13:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sharifzadeh:2009:PSS

[734] Mehdi Sharifzadeh, Cyrus Shahabi, and Leyla Kazemi. Processing spatial skyline queries in both vector spaces and spatial network databases. ACM Transactions on Database Systems, 34 (3):14:1–14:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yi:2009:SSG

[735] Ke Yi, Feifei Li, Graham Cormode, Marios Hadjieleftheriou, George Kollios, and Divesh Srivastava. Small synopses for group-by query verification on outsourced data streams. ACM Transactions on Database Systems, 34 (3):15:1–15:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Perez:2009:SCS

[736] Jorge Pérez, Marcelo Arenas, and Claudio Gutierrez. Semantics and complexity of SPARQL. ACM Transactions on Database Systems, 34(3): 16:1–16:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Markowetz:2009:KSR

[737] Alexander Markowetz, Yin Yang, and Dimitris Papadias. Keyword search over relational tables and streams. ACM Transactions on Database Systems, 34(3):17:1–17:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cohen:2009:ICP

[738] Sara Cohen, Benny Kimelfeld, and Yehoshua Sagiv. Incorporating constraints in probabilistic XML. ACM Transactions on Database Systems, 34 (3):18:1–18:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sasha:2009:FTS

[739] Dennis Sasha, Maurizio Lenzerini, and Z. Meral Özsoyoğlu. Foreword to TODS SIGMOD/PODS 2008 special issue. ACM Transactions on Database Systems, 34(4):19:1–19:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cahill:2009:SIS

[740] Michael J. Cahill, Uwe Röhm, and Alan D. Fekete. Serializable isolation for snapshot databases. ACM Transactions on Database Systems, 34(4): 20:1–20:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

He:2009:RQC

[741] Bingsheng He, Mian Lu, Ke Yang, Rui Fang, Naga K. Govindaraju, Qiong Luo, and Pedro V. Sander. Relational query coprocessing on graphics processors. ACM Transactions on Database Systems, 34(4):21:1–21:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2009:RSM

[742] Marcelo Arenas, Jorge Pérez, and Cristian Riveros. The recovery of a schema mapping: Bringing exchanged data back. ACM Transactions on Database Systems, 34(4):22:1–22:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Abiteboul:2009:SAA

[743] Serge Abiteboul, Luc Segoufin, and Victor Vianu. Static analysis of active XML systems. ACM Transactions on Database Systems, 34(4): 23:1–23:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chow:2009:CQP

[744] Chi-Yin Chow, Mohamed F. Mokbel, and Walid G. Aref. Casper*: Query processing for location services without compromising privacy. ACM Transactions on Database Systems, 34(4): 24:1–24:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Benedikt:2009:XRL

[745] Michael Benedikt and Christoph Koch. From XQuery to relational logics. ACM Transactions on Database Systems, 34(4):25:1–25:??, December 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ghanem:2010:SVD

[746] Thanaa M. Ghanem, Ahmed K. Elmagarmid, Per-Åke Larson, and Walid G. Aref. Supporting views in data stream management systems. ACM Transactions on Database Systems, 35(1): 1:1–1:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wu:2010:AML

[747] Kesheng Wu, Arie Shoshani, and Kurt Stockinger. Analyses of multilevel and multi-component compressed bitmap indexes. ACM Transactions on Database Systems, 35(1):2:1–2:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lian:2010:RSS

[748] Xiang Lian and Lei Chen. Reverse skyline search in uncertain databases. ACM Transactions on Database Systems, 35(1):3:1–3:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Augsten:2010:GDB

[749] Nikolaus Augsten, Michael Böhlen, and Johann Gamper. The pq-gram distance between ordered labeled trees. ACM Transactions on Database Systems, 35(1):4:1–4:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kolahi:2010:ITA

[750] Solmaz Kolahi and Leonid Libkin. An information-theoretic analysis of worstcase redundancy in database design. ACM Transactions on Database Systems, 35(1):5:1–5:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schnaitter:2010:OAE

[751] Karl Schnaitter and Neoklis Polyzotis. Optimal algorithms for evaluating rank joins in database systems. ACM Transactions on Database Systems, 35 (1):6:1–6:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Soror:2010:AVM

[752] Ahmed A. Soror, Umar Farooq Minhas, Ashraf Aboulnaga, Kenneth Salem, Peter Kokosielis, and Sunil Kamath. Automatic virtual machine configuration for database workloads. ACM Transactions on Database Systems, 35(1): 7:1–7:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xiao:2010:TAT

[753] Xiaokui Xiao, Yufei Tao, and Nick Koudas. Transparent anonymization: Thwarting adversaries who know the algorithm. ACM Transactions on Database Systems, 35(2):8:1–8:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

U:2010:OMB

[754] Leong Hou U., Kyriakos Mouratidis, Man Lung Yiu, and Nikos Mamoulis. Optimal matching between spatial datasets under capacity constraints. ACM Transactions on Database Systems, 35(2):9:1–9:??, April 2010. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2010:RSI

[755] Ziyang Liu and Yi Chen. Return specification inference and result clustering for keyword search on XML. ACM Transactions on Database Systems, 35 (2):10:1–10:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bex:2010:ICR

[756] Geert Jan Bex, Frank Neven, Thomas Schwentick, and Stijn Vansummeren. Inference of concise regular expressions and DTDs. ACM Transactions on Database Systems, 35(2):11:1–11:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

DeCapitaniDiVimercati:2010:EPR

[757] Sabrina De Capitani Di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, and Pierangela Samarati. Encryption policies for regulating access to outsourced data. ACM Transactions on Database Systems, 35(2): 12:1–12:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Koutrika:2010:PQB

[758] Georgia Koutrika and Yannis Ioannidis. Personalizing queries based on networks of composite preferences. ACM Transactions on Database Systems, 35 (2):13:1–13:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Flesca:2010:QRI

[759] Sergio Flesca, Filippo Furfaro, and Francesco Parisi. Querying and repairing inconsistent numerical databases. ACM Transactions on Database Systems, 35(2):14:1–14:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2010:CIE

[760] Hongyan Liu, Xiaoyu Wang, and Yinghui Yang. Comments on “An integrated efficient solution for computing frequent and top-k elements in data streams”. ACM Transactions on Database Systems, 35(2):15:1–15:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [648].

Graefe:2010:SBT

[761] Goetz Graefe. A survey of B-tree locking techniques. ACM Transactions on Database Systems, 35(3):16:1–16:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chen:2010:COI

[762] Su Chen, Mario A. Nascimento, Beng Chin Ooi, and Kian-Lee Tan.

Continuous online index tuning in moving object databases. ACM Transactions on Database Systems, 35(3):17:1– 17:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hu:2010:PAL

[763] Haibo Hu, Jianliang Xu, Sai Tung On, Jing Du, and Joseph Kee-Yin Ng. Privacy-aware location data publishing. ACM Transactions on Database Systems, 35(3):18:1–18:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2010:IXS

[764] Ziyang Liu, Yu Huang, and Yi Chen. Improving XML search by generating and utilizing informative result snippets. ACM Transactions on Database Systems, 35(3):19:1–19:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tao:2010:EAN

[765] Yufei Tao, Ke Yi, Cheng Sheng, and Panos Kalnis. Efficient and accurate nearest neighbor and closest pair search in high-dimensional space. ACM Transactions on Database Systems, 35 (3):20:1–20:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nash:2010:VQD

[766] Alan Nash, Luc Segoufin, and Victor Vianu. Views and queries: Determinacy and rewriting. ACM Transactions on Database Systems, 35(3):21:1–21:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Denecker:2010:TLR

[767] Marc Denecker, Álvaro Cortés-Calabuig, Maurice Bruynooghes, and Ofer Arieli. Towards a logical reconstruction of a theory for locally closed databases. ACM Transactions on Database Systems, 35(3):22:1–22:??, July 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2010:FTI

[768] Z. Meral Özsoyoglu. Foreword to TODS invited papers issue. ACM Transactions on Database Systems, 35 (4):23:1–23:??, November 2010. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ivanova:2010:ARI

[769] Milena G. Ivanova, Martin L. Kersten, Niels J. Nes, and Romulo A. P. Gonçalves. An architecture for recycling intermediates in a columnstore. ACM Transactions on Database Systems, 35(4):24:1–24:??, November 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ghoting:2010:EAS

[770] Amol Ghoting and Konstantin Makarychev. I/O efficient algorithms for serial and parallel suffix tree construction. ACM Transactions on Database Systems, 35 (4):25:1–25:??, November 2010. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Berinde:2010:SOH

[771] Radu Berinde, Piotr Indyk, Graham Cormode, and Martin J. Strauss. Space-optimal heavy hitters with strong error bounds. ACM Transactions on Database Systems, 35(4): 26:1–26:??, November 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2010:RIC

[772] Wenfei Fan and Floris Geerts. Relative information completeness. ACM Transactions on Database Systems, 35 (4):27:1–27:??, November 2010. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fletcher:2010:TTS

[773] George H. L. Fletcher, Jan Van Den Bussche, Dirk Van Gucht, and Stijn Vansummeren. Towards a theory of search queries. ACM Transactions on Database Systems, 35(4):28:1–28:??, November 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bjorklund:2010:IXE

[774] Henrik Björklund, Wouter Gelade, and Wim Martens. Incremental XPath evaluation. ACM Transactions on Database Systems, 35(4):29:1–29:??, November 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wang:2011:OPD

[775] Ting Wang and Ling Liu. Output privacy in data mining. ACM Transactions on Database Systems, 36 (1):1:1–1:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nergiz:2011:IA

[776] Mehmet Ercan Nergiz, Acar Tamersoy, and Yucel Saygin. Instant anonymization. ACM Transactions on Database Systems, 36(1):2:1–2:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gross-Amblard:2011:QPW

[777] David Gross-Amblard. Querypreserving watermarking of relational databases and XML documents. ACM Transactions on Database Systems, 36 (1):3:1–3:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Termehchy:2011:USI

[778] Arash Termehchy and Marianne Winslett. Using structural information in XML keyword search effectively. ACM Transactions on Database Systems, 36(1):4:1–4:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cautis:2011:QXD

[779] Bogdan Cautis, Alin Deutsch, Nicola Onose, and Vasilis Vassalos. Querying XML data sources that export very large sets of views. ACM Transactions on Database Systems, 36(1):5:1–5:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2011:GSW

[780] Ziyang Liu, Susan B. Davidson, and Yi Chen. Generating sound workflow views for correct provenance analysis. ACM Transactions on Database Systems, 36(1):6:1–6:??, March 2011.

CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jin:2011:PTE

[781] Ruoming Jin, Ning Ruan, Yang Xiang, and Haixun Wang. Path-tree: An efficient reachability indexing scheme for large directed graphs. ACM Transactions on Database Systems, 36 (1):7:1–7:??, March 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Law:2011:RLD

[782] Yan-Nei Law, Haixun Wang, and Carlo Zaniolo. Relational languages and data models for continuous queries on sequences and data streams. ACM Transactions on Database Systems, 36 (2):8:1–8:??, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gao:2011:CNN

[783] Yunjun Gao, Baihua Zheng, Gang Chen, Chun Chen, and Qing Li. Continuous nearest-neighbor search in the presence of obstacles. ACM Transactions on Database Systems, 36(2):9:1– 9:??, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yiu:2011:DAR

[784] Man Lung Yiu, Christian S. Jensen, Jesper Møller, and Hua Lu. Design and analysis of a ranking approach to private location-based services. ACM Transactions on Database Systems, 36 (2):10:1–10:??, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2011:RDE

[785] Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Reverse data exchange: Coping with nulls. ACM Transactions on Database Systems, 36(2):11:1–11:??, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Atallah:2011:AEA

[786] Mikhail J. Atallah, Yinian Qi, and Hao Yuan. Asymptotically efficient algorithms for skyline probabilities of uncertain data. ACM Transactions on Database Systems, 36(2):12:1–12:28, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lee:2011:RAF

[787] Tony T. Lee and Tong Ye. A relational approach to functional decomposition of logic circuits. ACM Transactions on Database Systems, 36 (2):13:1–13:??, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hernich:2011:CWD

[788] André Hernich, Leonid Libkin, and Nicole Schweikardt. Closed world data exchange. ACM Transactions on Database Systems, 36(2):14:1–14:40, May 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xiao:2011:ESJ

[789] Chuan Xiao, Wei Wang, Xuemin Lin, Jeffrey Xu Yu, and Guoren Wang. Efficient similarity joins for near-duplicate detection. ACM Transactions on Database Systems, 36(3): 15:1–15:41, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Song:2011:DDR

[790] Shaoxu Song and Lei Chen. Differential dependencies: Reasoning and discovery. ACM Transactions on Database Systems, 36(3):16:1–16:41, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See comments [911] and response [958].

Papapetrou:2011:EBS

[791] Panagiotis Papapetrou, Vassilis Athitsos, Michalis Potamias, George Kollios, and Dimitrios Gunopulos. Embeddingbased subsequence matching in timeseries databases. ACM Transactions on Database Systems, 36(3):17:1–17:39, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jampani:2011:MCD

[792] Ravi Jampani, Fei Xu, Mingxi Wu, Luis Perez, Chris Jermaine, and Peter J. Haas. The Monte Carlo database system: Stochastic analysis close to the data. ACM Transactions on Database Systems, 36(3): 18:1–18:41, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Stefanidis:2011:SRC

[793] Kostas Stefanidis, Georgia Koutrika, and Evaggelia Pitoura. A survey on representation, composition and application of preferences in database systems. ACM Transactions on Database Systems, 36(3):19:1–19:45, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2011:FTI

[794] Z. Meral Özsoyoglu. Foreword to TODS invited papers issue 2011. ACM Transactions on Database Systems, 36 (4):20:1–20:2, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheng:2011:FMC

[795] James Cheng, Yiping Ke, Ada Wai-Chee Fu, Jeffrey Xu Yu, and Linhong Zhu. Finding maximal cliques in massive networks. ACM Transactions on Database Systems, 36(4): 21:1–21:34, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kim:2011:DFA

[796] Changkyu Kim, Jatin Chhugani, Nadathur Satish, Eric Sedlar, Anthony D. Nguyen, Tim Kaldewey, Victor W. Lee, Scott A. Brandt, and Pradeep Dubey. Designing fast architecturesensitive tree search on modern multicore/many-core processors. ACM Transactions on Database Systems, 36 (4):22:1–22:34, December 2011. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Alexe:2011:CSM

[797] Bogdan Alexe, Balder ten Cate, Phokion G. Kolaitis, and Wang-Chiew Tan. Characterizing schema mappings via data examples. ACM Transactions on Database Systems, 36(4):

23:1–23:48, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cohen:2011:BET

[798] Sara Cohen and Yaacov Y. Weiss. Bag equivalence of tree patterns. ACM Transactions on Database Systems, 36 (4):24:1–24:35, December 2011. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Abiteboul:2011:CCD

[799] Serge Abiteboul, T.-H. Hubert Chan, Evgeny Kharlamov, Werner Nutt, and Pierre Senellart. Capturing continuous data and answering aggregate queries in probabilistic XML. ACM Transactions on Database Systems, 36(4): 25:1–25:45, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bai:2011:CPT

[800] Xiao Bai, Rachid Guerraoui, Anne-Marie Kermarrec, and Vincent Leroy. Collaborative personalized top-k processing. ACM Transactions on Database Systems, 36(4):26:1–26:38, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Goncalves:2011:DCQ

[801] Romulo Goncalves and Martin Kersten. The Data Cyclotron query processing scheme. ACM Transactions on Database Systems, 36(4):27:1–27:35, December 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Graefe:2012:SBT

[802] Goetz Graefe. A survey of B-tree logging and recovery techniques. ACM Transactions on Database Systems, 37 (1):1:1–1:35, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Martinenghi:2012:PMR

[803] Davide Martinenghi and Marco Tagliasacchi. Proximity measures for rank join. ACM Transactions on Database Systems, 37(1):2:1–2:46, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nuray-Turan:2012:AOS

[804] Rabia Nuray-Turan, Dmitri V. Kalashnikov, Sharad Mehrotra, and Yaming Yu. Attribute and object selection queries on objects with probabilistic attributes. ACM Transactions on Database Systems, 37(1): 3:1–3:??, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2012:DSR

[805] Ziyang Liu and Yi Chen. Differentiating search results on structured data. ACM Transactions on Database Systems, 37(1):4:1–4:??, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yang:2012:SES

[806] Di Yang, Elke A. Rundensteiner, and Matthew O. Ward. Shared execution strategy for neighbor-based pattern mining requests over streaming windows. ACM Transactions on Database Systems, 37(1):5:1–5:??, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Re:2012:UCE

[807] Christopher Ré and D. Suciu. Understanding cardinality estimation using entropy maximization. ACM Transactions on Database Systems, 37(1): 6:1–6:??, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nuray-Turan:2012:EWQ

[808] Rabia Nuray-Turan, Dmitri V. Kalashnikov, and Sharad Mehrotra. Exploiting Web querying for Web people search. ACM Transactions on Database Systems, 37(1):7:1–7:??, February 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schneider:2012:CDB

[809] Markus Schneider, Tao Chen, Ganesh Viswanathan, and Wenjie Yuan. Cardinal directions between complex regions. ACM Transactions on Database Systems, 37(2):8:1–8:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wijsen:2012:CCQ

[810] Jef Wijsen. Certain conjunctive query answering in first-order logic. ACM Transactions on Database Systems, 37 (2):9:1–9:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Abiteboul:2012:CWS

[811] Serge Abiteboul, Pierre Bourhis, and Victor Vianu. Comparing workflow specification languages: a matter of views. ACM Transactions on Database Systems, 37(2):10:1–10:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tassa:2012:SDC

[812] Tamir Tassa and Ehud Gudes. Secure distributed computation of anonymized views of shared databases. ACM Transactions on Database Systems, 37 (2):11:1–11:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sheng:2012:EAA

[813] Cheng Sheng, Yufei Tao, and Jianzhong Li. Exact and approximate algorithms for the most connected vertex problem. ACM Transactions on Database Systems, 37(2):12:1–12:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hartmann:2012:IPD

[814] Sven Hartmann and Sebastian Link. The implication problem of data dependencies over SQL table definitions: Axiomatic, algorithmic and logical characterizations. ACM Transactions on Database Systems, 37(2):13:1–13:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2012:SS

[815] Wenjie Zhang, Xuemin Lin, Ying Zhang, Muhammad Aamir Cheema, and Qing Zhang. Stochastic skylines. ACM Transactions on Database Systems, 37(2):14:1–14:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xia:2012:OSS

[816] Tian Xia, Donghui Zhang, Zheng Fang, Cindy Chen, and Jie Wang. Online subspace skyline query processing using the Compressed SkyCube. ACM Transactions on Database Systems, 37 (2):15:1–15:??, May 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2012:Fa

[817] Z. Meral Özsoyoglu. Foreword. ACM Transactions on Database Systems, 37 (3):16:1–16:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Graefe:2012:FBT

[818] Goetz Graefe, Hideaki Kimura, and Harumi Kuno. Foster B-trees. ACM Transactions on Database Systems, 37 (3):17:1–17:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wang:2012:RAT

[819] Junhu Wang and Jeffrey Xu Yu. Revisiting answering tree pattern queries using views. ACM Transactions on Database Systems, 37(3):18:1–18:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

David:2012:ERA

[820] Claire David, Leonid Libkin, and Tony Tan. Efficient reasoning about data trees via integer linear programming. ACM Transactions on Database Systems, 37(3):19:1–19:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lemire:2012:RRB

[821] Daniel Lemire, Owen Kaser, and Eduardo Gutarra. Reordering rows for better compression: Beyond the lexicographic order. ACM Transactions on Database Systems, 37(3):20:1–20:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gupta:2012:EQE

[822] Nitin Gupta, Lucja Kot, Sudip Roy, Gabriel Bender, Johannes Gehrke, and Christoph Koch. Entangled queries: Enabling declarative data-driven coordination. ACM Transactions on Database Systems, 37(3):21:1–21:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Damaggio:2012:ASD

[823] Elio Damaggio, Alin Deutsch, and Victor Vianu. Artifact systems with data dependencies and arithmetic. ACM Transactions on Database Systems, 37 (3):22:1–22:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2012:Fb

[824] Z. Meral Özsoyoglu. Foreword. ACM Transactions on Database Systems, 37 (4):23:1–23:??, December 2012. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kimelfeld:2012:MCV

[825] Benny Kimelfeld, Jan Vondrák, and Ryan Williams. Maximizing conjunctive views in deletion propagation.

ACM Transactions on Database Systems, 37(4):24:1–24:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2012:DCD

[826] Wenfei Fan, Floris Geerts, and Jef Wijsen. Determining the currency of data. ACM Transactions on Database Systems, 37(4):25:1–25:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sheng:2012:WCE

[827] Cheng Sheng and Yufei Tao. Worstcase I/O-efficient skyline algorithms. ACM Transactions on Database Systems, 37(4):26:1–26:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2012:SPS

[828] Boduo Li, Edward Mazur, Yanlei Diao, Andrew McGregor, and Prashant Shenoy. SCALLA: a platform for scalable one-pass analytics using MapReduce. ACM Transactions on Database Systems, 37(4):27:1–27:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2012:WHT

[829] Yinan Li, Jignesh M. Patel, and Allison Terrell. WHAM: a high-throughput sequence alignment method. ACM Transactions on Database Systems, 37 (4):28:1–28:??, December 2012. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xie:2012:FAS

[830] Kexin Xie, Ke Deng, Shuo Shang, Xiaofang Zhou, and Kai Zheng. Finding alternative shortest paths in spatial networks. ACM Transactions on Database Systems, 37(4):29:1–29:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Amsterdamer:2012:PM

[831] Yael Amsterdamer, Daniel Deutch, Tova Milo, and Val Tannen. On provenance minimization. ACM Transactions on Database Systems, 37(4): 30:1–30:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Barcelo:2012:ELP

[832] Pablo Barceló, Leonid Libkin, Anthony W. Lin, and Peter T. Wood. Expressive languages for path queries over graph-structured data. ACM Transactions on Database Systems, 37(4): 31:1–31:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cong:2012:PED

[833] Gao Cong, Wenfei Fan, Anastasios Kementsietsidis, Jianzhong Li, and Xianmin Liu. Partial evaluation for distributed XPath query processing and beyond. ACM Transactions on Database Systems, 37(4):32:1–32:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Papavasileiou:2013:HLC

[834] Vicky Papavasileiou, Giorgos Flouris, Irini Fundulaki, Dimitris Kotzinos, and Vassilis Christophides. High-level change detection in RDF(S) KBs. ACM Transactions on Database Systems, 38 (1):1:1–1:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jaiswal:2013:SME

[835] Anuj Jaiswal, David J. Miller, and Prasenjit Mitra. Schema matching and embedded value mapping for databases with opaque column names and mixed continuous and discrete-valued data fields. ACM Transactions on Database Systems, 38(1):2:1–2:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Grust:2013:OSQ

[836] Torsten Grust and Jan Rittinger. Observing SQL queries in their natural habitat. ACM Transactions on Database Systems, 38(1):3:1–3:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Benzaken:2013:OXQ

[837] Véronique Benzaken, Giuseppe Castagna, Dario Colazzo, and Kim Nguy´ ên. Optimizing XML querying using typebased document projection. ACM Transactions on Database Systems, 38 (1):4:1–4:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Das:2013:EES

[838] Sudipto Das, Divyakant Agrawal, and Amr El Abbadi. ElasTraS: an elastic, scalable, and self-managing transactional database for the cloud. ACM Transactions on Database Systems, 38 (1):5:1–5:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Franceschet:2013:GTA

[839] Massimo Franceschet, Donatella Gubiani, Angelo Montanari, and Carla Piazza. A graph-theoretic approach to map conceptual designs to XML schemas. ACM Transactions on Database Systems, 38(1):6:1–6:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wu:2013:MSK

[840] Dingming Wu, Man Lung Yiu, and Christian S. Jensen. Moving spatial keyword queries: Formulation, methods, and analysis. ACM Transactions on Database Systems, 38(1):7:1–7:??, April 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sadoghi:2013:AOB

[841] Mohammad Sadoghi and Hans-Arno Jacobsen. Analysis and optimization for Boolean expression indexing. ACM Transactions on Database Systems, 38 (2):8:1–8:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2013:PBM

[842] Guoliang Li, Dong Deng, and Jianhua Feng. A partition-based method for string similarity joins with editdistance constraints. ACM Transactions on Database Systems, 38(2):9:1– 9:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Catallo:2013:TKD

[843] Ilio Catallo, Eleonora Ciceri, Piero Fraternali, Davide Martinenghi, and Marco Tagliasacchi. Top-k diversity queries over bounded regions. ACM Transactions on Database Systems, 38 (2):10:1–10:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fazzinga:2013:RDC

[844] Bettina Fazzinga, Sergio Flesca, Filippo Furfaro, and Elio Masciari. RFIDdata compression for supporting aggregate queries. ACM Transactions on Database Systems, 38(2):11:1–11:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pavlou:2013:GDF

[845] Kyriacos E. Pavlou and Richard T. Snodgrass. Generalizing database forensics. ACM Transactions on Database Systems, 38(2):12:1–12:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheney:2013:RFN

[846] James Cheney. Revisiting “forward node-selecting queries over trees”. ACM Transactions on Database Systems, 38(2):13:1–13:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bonifati:2013:AIM

[847] Angela Bonifati, Martin Goodfellow, Ioana Manolescu, and Domenica Sileo. Algebraic incremental maintenance of XML views. ACM Transactions on Database Systems, 38(3):14:1–14:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Colazzo:2013:ALI

[848] Dario Colazzo, Giorgio Ghelli, Luca Pardini, and Carlo Sartiani. Almostlinear inclusion for XML regular expression types. ACM Transactions on Database Systems, 38(3):15:1–15:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Qin:2013:ASS

[849] Jianbin Qin, Wei Wang, Chuan Xiao, Yifei Lu, Xuemin Lin, and Haixun Wang. Asymmetric signature schemes for efficient exact edit similarity query processing. ACM Transactions on Database Systems, 38(3):16:1–16:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Levandoski:2013:FEP

[850] Justin J. Levandoski, Ahmed Eldawy, Mohamed F. Mokbel, and Mohamed E. Khalefa. Flexible and extensible preference evaluation in database systems. ACM Transactions on Database Systems, 38(3):17:1–17:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2013:IGP

[851] Wenfei Fan, Xin Wang, and Yinghui Wu. Incremental graph pattern matching. ACM Transactions on Database Systems, 38(3):18:1–18:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Karvounarakis:2013:CDS

[852] Grigoris Karvounarakis, Todd J. Green, Zachary G. Ives, and Val Tannen. Collaborative data sharing via update exchange and provenance. ACM Transactions on Database Systems, 38 (3):19:1–19:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ozsoyoglu:2013:FIP

[853] Z. Meral Özsoyoglu. Foreword to invited papers issue. ACM Transactions on Database Systems, 38(4): 20:1–20:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mozafari:2013:HPC

[854] Barzan Mozafari, Kai Zeng, Loris D’antoni, and Carlo Zaniolo. Highperformance complex event processing over hierarchical data. ACM Transactions on Database Systems, 38(4): 21:1–21:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sarma:2013:CTL

[855] Anish Das Sarma, Hongrae Lee, Hector Gonzalez, Jayant Madhavan, and Alon Halevy. Consistent thinning of large geographical data for map visualization. ACM Transactions on Database Systems, 38(4):22:1–22:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Teubner:2013:XFB

[856] Jens Teubner, Louis Woods, and Chongling Nie. XLynx — an FPGAbased XML filter for hybrid XQuery processing. ACM Transactions on Database Systems, 38(4):23:1–23:??, November 2013. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic).

Losemann:2013:CRE

[857] Katja Losemann and Wim Martens. The complexity of regular expressions and property paths in SPARQL. ACM Transactions on Database Systems, 38 (4):24:1–24:??, November 2013. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Letelier:2013:SAO

[858] Andrés Letelier, Jorge Pérez, Reinhard Pichler, and Sebastian Skritek. Static analysis and optimization of Semantic Web queries. ACM Transactions on Database Systems, 38(4): 25:1–25:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Agarwal:2013:MS

[859] Pankaj K. Agarwal, Graham Cormode, Zengfeng Huang, Jeff M. Phillips, Zhewei Wei, and Ke Yi. Mergeable summaries. ACM Transactions on Database Systems, 38(4):26:1–26:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Konrad:2013:VXD

[860] Christian Konrad and Frédéric Magniez. Validating XML documents in the streaming model with external memory. ACM Transactions on Database Systems, 38(4):27:1–27:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cate:2013:LSM

[861] Balder Ten Cate, Víctor Dalmau, and Phokion G. Kolaitis. Learning schema mappings. ACM Transactions on Database Systems, 38(4): 28:1–28:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kostylev:2014:CAS

[862] Egor V. Kostylev, Juan L. Reutter, and András Z. Salamon. Classification of annotation semirings over containment of conjunctive queries. ACM Transactions on Database Systems, 39 (1):1:1–1:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yi:2014:ISQ

[863] Ke Yi, Lu Wang, and Zhewei Wei. Indexing for summary queries: Theory and practice. ACM Transactions on Database Systems, 39(1):2:1–2:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kifer:2014:PFM

[864] Daniel Kifer and Ashwin Machanavajjhala. Pufferfish: a framework for mathematical privacy definitions. ACM Transactions on Database Systems, 39 (1):3:1–3:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ma:2014:SSC

[865] Shuai Ma, Yang Cao, Wenfei Fan, Jinpeng Huai, and Tianyu Wo. Strong simulation: Capturing topology in graph pattern matching. ACM Transactions on Database Systems, 39(1): 4:1–4:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gatterbauer:2014:OBP

[866] Wolfgang Gatterbauer and Dan Suciu. Oblivious bounds on the probability of Boolean functions. ACM Transactions on Database Systems, 39 (1):5:1–5:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fang:2014:MOP

[867] Qiong Fang, Wilfred Ng, Jianlin Feng, and Yuliang Li. Mining orderpreserving submatrices from probabilistic matrices. ACM Transactions on Database Systems, 39(1):6:1–6:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wang:2014:ESS

[868] Jiannan Wang, Guoliang Li, and Jianhua Feng. Extending string similarity join to tolerant fuzzy token matching. ACM Transactions on Database Systems, 39(1):7:1–7:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sen:2014:DRM

[869] Siddhartha Sen and Robert E. Tarjan. Deletion without rebalancing in multiway search trees. ACM Transactions on Database Systems, 39(1): 8:1–8:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nekrich:2014:ERS

[870] Yakov Nekrich. Efficient range searching for categorical and plain data. ACM Transactions on Database Systems, 39 (1):9:1–9:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Xu:2014:OCP

[871] Pan Xu and Srikanta Tirthapura. Optimality of clustering properties of space-filling curves. ACM Transactions on Database Systems, 39(2):10:1–10:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Thomson:2014:FDT

[872] Alexander Thomson, Thaddeus Diamond, Shu-Chun Weng, Kun Ren, Philip Shao, and Daniel J. Abadi. Fast distributed transactions and strongly consistent replication for OLTP database systems. ACM Transactions on Database Systems, 39(2): 11:1–11:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nykiel:2014:SAM

[873] Tomasz Nykiel, Michalis Potamias, Chaitanya Mishra, George Kollios, and Nick Koudas. Sharing across multiple MapReduce jobs. ACM Transactions on Database Systems, 39(2):12:1–12:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lu:2014:EAC

[874] Ying Lu, Jiaheng Lu, Gao Cong, Wei Wu, and Cyrus Shahabi. Efficient algorithms and cost models for reverse spatial-keyword k-nearest neighbor search. ACM Transactions on Database Systems, 39(2):13:1–13:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bartolini:2014:DPW

[875] Ilaria Bartolini, Paolo Ciaccia, and Marco Patella. Domination in the probabilistic world: Computing skylines for arbitrary correlations and ranking semantics. ACM Transactions on Database Systems, 39(2):14:1–14:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Deng:2014:CQR

[876] Ting Deng and Wenfei Fan. On the complexity of query result diversification. ACM Transactions on Database Systems, 39(2):15:1–15:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Giatrakos:2014:DGQ

[877] Nikos Giatrakos, Antonios Deligiannakis, Minos Garofalakis, Izchak Sharfman, and Assaf Schuster. Distributed geometric query monitoring using prediction models. ACM Transactions on Database Systems, 39(2):16:1–16:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lei:2014:RDQ

[878] Chuan Lei and Elke A. Rundensteiner. Robust distributed query processing for streaming data. ACM Transactions on Database Systems, 39(2):17:1–17:??, May 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2014:E

[879] Christian S. Jensen. Editorial. ACM Transactions on Database Systems, 39 (3):18:1–18:??, September 2014. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2014:TPI

[880] Rui Zhang, Jianzhong Qi, Martin Stradling, and Jin Huang. Towards a painless index for spatial objects. ACM Transactions on Database Systems, 39 (3):19:1–19:??, September 2014. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tatti:2014:FRI

[881] Nikolaj Tatti, Fabian Moerchen, and Toon Calders. Finding robust itemsets under subsampling. ACM Transactions on Database Systems, 39(3): 20:1–20:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Choi:2014:MRS

[882] Dong-Wan Choi, Chin-Wan Chung, and Yufei Tao. Maximizing range sum in external memory. ACM Transactions on Database Systems, 39(3): 21:1–21:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Karwa:2014:PAG

[883] Vishesh Karwa, Sofya Raskhodnikova, Adam Smith, and Grigory Yaroslavtsev. Private analysis of graph structure. ACM Transactions on Database Systems, 39(3):22:1–22:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pang:2014:PPA

[884] Hweehwa Pang and Xuhua Ding. Privacy-preserving ad-hoc equi-join on outsourced data. ACM Transactions on Database Systems, 39(3):23:1–23:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Malvestuto:2014:JLO

[885] Francesco M. Malvestuto. A joinlike operator to combine data cubes and answer queries from multiple data cubes. ACM Transactions on Database Systems, 39(3):24:1–24:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gottlob:2014:QRO

[886] Georg Gottlob, Giorgio Orsi, and Andreas Pieris. Query rewriting and optimization for ontological databases. ACM Transactions on Database Systems, 39(3):25:1–25:??, September 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2014:FIA

[887] Christian S. Jensen. Foreword to invited articles issue. ACM Transactions on Database Systems, 39(4): 26:1–26:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hu:2014:EAT

[888] Xiaocheng Hu, Yufei Tao, and Chin-Wan Chung. I/O-efficient algorithms on triangle listing and counting. ACM Transactions on Database Systems, 39 (4):27:1–27:??, December 2014. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2014:DXK

[889] Marcelo Arenas, Jonny Daenen, Frank Neven, Martin Ugarte, Jan Van Den Bussche, and Stijn Vansummeren. Discovering XSD keys from XML data. ACM Transactions on Database Systems, 39(4):28:1–28:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jung:2014:SLM

[890] Hyungsoo Jung, Hyuck Han, Alan Fekete, Gernot Heiser, and Heon Y. Yeom. A scalable lock manager for multicores. ACM Transactions on Database Systems, 39(4):29:1–29:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Papadopoulos:2014:LQA

[891] Stavros Papadopoulos, Graham Cormode, Antonios Deligiannakis, and Minos Garofalakis. Lightweight query authentication on streams. ACM Transactions on Database Systems, 39(4): 30:1–30:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gheerbrant:2014:NEQ

[892] Amélie Gheerbrant, Leonid Libkin, and Cristina Sirangelo. Naïve evaluation of queries over incomplete databases. ACM Transactions on Database Systems, 39(4):31:1–31:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kimelfeld:2014:CMM

[893] Benny Kimelfeld and Phokion G. Kolaitis. The complexity of mining maximal frequent subgraphs. ACM Transactions on Database Systems, 39(4): 32:1–32:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bienvenu:2014:OBD

[894] Meghyn Bienvenu, Balder Ten Cate, Carsten Lutz, and Frank Wolter. Ontology-based data access: a study through disjunctive Datalog, CSP, and MMSNP. ACM Transactions on Database Systems, 39(4):33:1–33:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2014:TPP

[895] Chao Li, Daniel Yang Li, Gerome Miklau, and Dan Suciu. A theory of pricing private data. ACM Transactions on Database Systems, 39(4): 34:1–34:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Davidson:2014:TCN

[896] Susan Davidson, Sanjeev Khanna, Tova Milo, and Sudeepa Roy. Top-k and clustering with noisy comparisons. ACM Transactions on Database Systems, 39(4):35:1–35:??, December 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2015:EUE

[897] Christian S. Jensen. Editorial: Updates to the editorial board. ACM Transactions on Database Systems, 40 (1):1:1–1:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pripuzic:2015:TSE

[898] Kresimir Pripuzić, Ivana Podnar Zarko, and Karl Aberer. Time- and space-efficient sliding window top-k query processing. ACM Transactions on Database Systems, 40(1):1:1–1:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Olteanu:2015:SBF

[899] Dan Olteanu and Jakub Závodný. Size bounds for factorised representations of query results. ACM Transactions on Database Systems, 40(1):2:1–2:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pawlik:2015:ECT

[900] Mateusz Pawlik and Nikolaus Augsten. Efficient computation of the tree edit distance. ACM Transactions on Database Systems, 40(1):3:1–3:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Drosou:2015:MRD

[901] Marina Drosou and Evaggelia Pitoura. Multiple radii DisC diversity: Result diversification based on dissimilarity and coverage. ACM Transactions on Database Systems, 40(1):4:1–4:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Benedikt:2015:ASA

[902] Michael Benedikt, Pierre Bourhis, and Clemens Ley. Analysis of schemas with access restrictions. ACM Transactions on Database Systems, 40(1):5:1–5:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Athanassoulis:2015:OUD

[903] Manos Athanassoulis, Shimin Chen, Anastasia Ailamaki, Philip B. Gibbons, and Radu Stoica. Online updates on data warehouses via judicious use of solid-state storage. ACM Transactions on Database Systems, 40 (1):6:1–6:??, March 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2015:EBT

[904] Christian S. Jensen. Editorial: The best of two worlds — present your TODS paper at SIGMOD. ACM Transactions on Database Systems, 40 (2):7:1–7:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Beedkar:2015:CGS

[905] Kaustubh Beedkar, Klaus Berberich, Rainer Gemulla, and Iris Miliaraki. Closing the gap: Sequence mining at scale. ACM Transactions on Database Systems, 40(2):8:1–8:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ameloot:2015:DDF

[906] Tom J. Ameloot. Deciding determinism with fairness for simple transducer networks. ACM Transactions on Database Systems, 40(2):9:1–9:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Nagendra:2015:EPS

[907] Mithila Nagendra and K. Selçuk Candan. Efficient processing of skylinejoin queries over multiple data sources. ACM Transactions on Database Systems, 40(2):10:1–10:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yuan:2015:OBL

[908] Ganzhao Yuan, Zhenjie Zhang, Marianne Winslett, Xiaokui Xiao, Yin Yang, and Zhifeng Hao. Optimizing batch linear queries under exact and approximate differential privacy. ACM Transactions on Database Systems, 40 (2):11:1–11:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Termehchy:2015:CEC

[909] Arash Termehchy, Ali Vakilian, Yodsawalai Chodpathumwan, and Marianne Winslett. Cost-effective conceptual design for information extraction. ACM Transactions on Database Systems, 40(2):12:1–12:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cao:2015:EPS

[910] Xin Cao, Gao Cong, Tao Guo, Christian S. Jensen, and Beng Chin Ooi. Efficient processing of spatial group keyword queries. ACM Transactions on Database Systems, 40(2):13:1–13:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Vincent:2015:TCD

[911] Millist Vincent, Jixue Liu, Hong-Cheu Liu, and Sebastian Link. Technical correspondence: “Differential Dependencies: Reasoning and Discovery” revisited. ACM Transactions on Database Systems, 40(2):14:1–14:??, June 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [790] and response [958].

Lu:2015:BQA

[912] Jiaheng Lu, Chunbin Lin, Wei Wang, Chen Li, and Xiaokui Xiao. Boosting the quality of approximate string matching by synonyms. ACM Transactions on Database Systems, 40(3): 15:1–15:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rusu:2015:WDA

[913] Florin Rusu, Zixuan Zhuang, Mingxi Wu, and Chris Jermaine. Workloaddriven antijoin cardinality estimation. ACM Transactions on Database Systems, 40(3):16:1–16:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chen:2015:OLQ

[914] Zitong Chen, Yubao Liu, Raymond Chi-Wing Wong, Jiamin Xiong, Ganglin Mai, and Cheng Long. Optimal location queries in road networks. ACM Transactions on Database Systems, 40(3):17:1–17:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rietveld:2015:RLD

[915] Kristian F. D. Rietveld and Harry A. G. Wijshoff. Reducing layered database applications to their essence through vertical integration. ACM Transactions on Database Systems, 40 (3):18:1–18:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheng:2015:SDM

[916] Yu Cheng and Florin Rusu. SCAN-RAW: a database meta-operator for parallel in-situ processing and loading. ACM Transactions on Database Systems, 40(3):19:1–19:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Parchas:2015:UGP

[917] Panos Parchas, Francesco Gullo, Dimitris Papadias, and Francesco Bonchi. Uncertain graph processing through representative instances. ACM Transactions on Database Systems, 40(3): 20:1–20:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Ameloot:2016:WFM

[918] Tom J. Ameloot, Bas Ketsman, Frank Neven, and Daniel Zinn. Weaker forms of monotonicity for declarative networking: a more fine-grained answer to the CALM-conjecture. ACM Transactions on Database Systems, 40(4): 21:1–21:??, January 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bonifati:2016:LJQ

[919] Angela Bonifati, Radu Ciucanu, and Sławek Staworko. Learning join queries from user examples. ACM Transactions on Database Systems, 40(4): 24:1–24:??, January 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yang:2016:NFI

[920] Xiaochun Yang, Tao Qiu, Bin Wang, Baihua Zheng, Yaoshu Wang, and Chen Li. Negative factor: Improving regular-expression matching in strings. ACM Transactions on Database Systems, 40(4):25:1–25:46, January 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhou:2016:FRW

[921] Zhuojie Zhou, Nan Zhang, Zhiguo Gong, and Gautam Das. Faster random walks by rewiring online social networks on-the-fly. ACM Transactions on Database Systems, 40(4): 26:1–26:??, February 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2016:EUE

[922] Christian S. Jensen. Editorial: Updates to the editorial board. ACM Transactions on Database Systems, 41 (1):1:1–1:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Libkin:2016:STV

[923] Leonid Libkin. SQL’s three-valued logic and certain answers. ACM Transactions on Database Systems, 41 (1):1:1–1:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2016:MOF

[924] Ce Zhang, Arun Kumar, and Christopher Ré. Materialization optimizations for feature selection workloads. ACM Transactions on Database Systems, 41 (1):2:1–2:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Olteanu:2016:EFP

[925] Dan Olteanu and Sebastiaan J. Van Schaik. ENFrame: a framework for processing probabilistic data. ACM Transactions on Database Systems, 41 (1):3:1–3:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fink:2016:DQN

[926] Robert Fink and Dan Olteanu. Dichotomies for queries with negation in probabilistic databases. ACM Transactions on Database Systems, 41(1):4:1– 4:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhou:2016:BEQ

[927] Xiaoling Zhou, Jianbin Qin, Chuan Xiao, Wei Wang, Xuemin Lin, and Yoshiharu Ishikawa. BEVA: an efficient query processing algorithm for error-tolerant autocompletion. ACM Transactions on Database Systems, 41 (1):5:1–5:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fagin:2016:DCI

[928] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Declarative cleaning of inconsistencies in information extraction. ACM Transactions on Database Systems, 41 (1):6:1–6:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pham:2016:ISS

[929] Huy Pham, Cyrus Shahabi, and Yan Liu. Inferring social strength from spatiotemporal data. ACM Transactions on Database Systems, 41(1):7:1–7:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cheung:2016:SBL

[930] Alvin Cheung, Samuel Madden, and Armando Solar-Lezama. Sloth: Being lazy is a virtue (when issuing database queries). ACM Transactions on Database Systems, 41(2):8:1–8:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2016:EDC

[931] Christian S. Jensen. Editorial: The dark citations of TODS papers and what to do about it — or: Cite the journal paper. ACM Transactions on Database Systems, 41(2):8:1–8:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Caruccio:2016:SQV

[932] Loredana Caruccio, Giuseppe Polese, and Genoveffa Tortora. Synchronization of queries and views upon schema evolutions: a survey. ACM Transactions on Database Systems, 41 (2):9:1–9:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Deng:2016:CMT

[933] Ting Deng, Wenfei Fan, and Floris Geerts. Capturing missing tuples and missing values. ACM Transactions on Database Systems, 41(2):10:1–10:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Dutt:2016:PBF

[934] Anshuman Dutt and Jayant R. Haritsa. Plan bouquets: a fragrant approach to robust query processing. ACM Transactions on Database Systems, 41 (2):11:1–11:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jung:2016:RWR

[935] Jinhong Jung, Kijung Shin, Lee Sael, and U. Kang. Random walk with restart on large graphs using block elimination. ACM Transactions on Database Systems, 41(2):12:1–12:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2016:ECF

[936] Marcelo Arenas and Gonzalo I. Diaz. The exact complexity of the firstorder logic definability problem. ACM Transactions on Database Systems, 41 (2):13:1–13:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cohen:2016:CLT

[937] Sara Cohen and Yaacov Y. Weiss. The complexity of learning tree patterns from example graphs. ACM Transactions on Database Systems, 41 (2):14:1–14:??, June 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bailis:2016:SAV

[938] Peter Bailis, Alan Fekete, Ali Ghodsi, Joseph M. Hellerstein, and Ion Stoica. Scalable atomic visibility with RAMP transactions. ACM Transactions on Database Systems, 41(3): 15:1–15:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

To:2016:PSE

[939] Quoc-Cuong To, Benjamin Nguyen, and Philippe Pucheral. Private and scalable execution of SQL aggregates on a secure decentralized architecture. ACM Transactions on Database Systems, 41(3):16:1–16:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Burdick:2016:DFL

[940] Douglas Burdick, Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. A declarative framework for linking entities. ACM Transactions on Database Systems, 41(3): 17:1–17:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bourhis:2016:BRR

[941] Pierre Bourhis, Gabriele Puppis, Cristian Riveros, and Slawek Staworko. Bounded repairability for regular tree languages. ACM Transactions on Database Systems, 41(3):18:1–18:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bender:2016:BTC

[942] Michael A. Bender, Roozbeh Ebrahimi, Haodong Hu, and Bradley C. Kuszmaul. B-trees and cache-oblivious Btrees with different-sized atomic keys. ACM Transactions on Database Systems, 41(3):19:1–19:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Mazowiecki:2016:MDR

[943] Filip Mazowiecki, Filip Murlak, and Adam Witkowski. Monadic Datalog and regular tree pattern queries. ACM Transactions on Database Systems, 41 (3):20:1–20:??, August 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tian:2016:BHW

[944] Yuanyuan Tian, Fatma Özcan, Tao Zou, Romulo Goncalves, and Hamid Pirahesh. Building a hybrid warehouse: Efficient joins between data stored in HDFS and enterprise warehouse. ACM Transactions on Database Systems, 41(4):21:1–21:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Khamis:2016:JGR

[945] Mahmoud Abo Khamis, Hung Q. Ngo, Christopher Ré, and Atri Rudra. Joins via geometric resolutions: Worst case and beyond. ACM Transactions on Database Systems, 41(4): 22:1–22:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Papadimitriou:2016:GBA

[946] Dimitra Papadimitriou, Georgia Koutrika, John Mylopoulos, and Yannis Velegrakis. The goal behind the action: Toward goal-aware systems and applications. ACM Transactions on Database Systems, 41(4):23:1–23:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fazzinga:2016:EIC

[947] Bettina Fazzinga, Sergio Flesca, Filippo Furfaro, and Francesco Parisi. Exploiting integrity constraints for cleaning trajectories of RFID-monitored objects. ACM Transactions on Database Systems, 41(4):24:1–24:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Maabout:2016:SMU

[948] Sofian Maabout, Carlos Ordonez, Patrick Kamnang Wanko, and Nicolas Hanusse. Skycube materialization using the topmost skyline or functional dependencies. ACM Transactions on Database Systems, 41(4): 25:1–25:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Dignos:2016:EKR

[949] Anton Dignös, Michael H. Böhlen, Johann Gamper, and Christian S. Jensen. Extending the kernel of a relational DBMS with comprehensive support for sequenced temporal queries. ACM Transactions on Database Systems, 41(4):26:1–26:??, December 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bourhis:2016:GBD

[950] Pierre Bourhis, Marco Manna, Michael Morak, and Andreas Pieris. Guardedbased disjunctive tuple-generating dependencies. ACM Transactions on Database Systems, 41(4):27:1–27:??, December 2016. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic).

Beame:2017:EMC

[951] Paul Beame, Jerry Li, Sudeepa Roy, and Dan Suciu. Exact model counting of query expressions: Limitations of propositional methods. ACM Transactions on Database Systems, 42 (1):1:1–1:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2017:EUE

[952] Christian S. Jensen. Editorial: Updates to the editorial board. ACM Transactions on Database Systems, 42 (1):1:1–1:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Liu:2017:SMD

[953] Xiufeng Liu, Lukasz Golab, Wojciech Golab, Ihab F. Ilyas, and Shichao Jin. Smart meter data analytics: Systems, algorithms, and benchmarking. ACM Transactions on Database Systems, 42 (1):2:1–2:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Currim:2017:DMM

[954] Sabah Currim, Richard T. Snodgrass, Young-Kyoon Suh, and Rui Zhang. DBMS metrology: Measuring query time. ACM Transactions on Database Systems, 42(1):3:1–3:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wu:2017:CFC

[955] You Wu, Pankaj K. Agarwal, Chengkai Li, Jun Yang, and Cong Yu. Computational fact checking through query perturbations. ACM Transactions on Database Systems, 42(1):4:1–4:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Teflioudi:2017:EAM

[956] Christina Teflioudi and Rainer Gemulla. Exact and approximate maximum inner product search with LEMP. ACM Transactions on Database Systems, 42 (1):5:1–5:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Shi:2017:UUA

[957] Xiaogang Shi, Bin Cui, Gillian Dobbie, and Beng Chin Ooi. UniAD: a unified ad hoc data processing system. ACM Transactions on Database Systems, 42(1):6:1–6:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Song:2017:RDD

[958] Shaoxu Song and Lei Chen. Response to “Differential Dependencies Revisited”. ACM Transactions on Database Systems, 42(1):7:1–7:??, March 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [790, 911].

Paparrizos:2017:FAT

[959] John Paparrizos and Luis Gravano. Fast and accurate time-series clustering. ACM Transactions on Database Systems, 42(2):8:1–8:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Koutris:2017:CQA

[960] Paraschos Koutris and Jef Wijsen. Consistent query answering for selfjoin-free conjunctive queries under primary key constraints. ACM Transactions on Database Systems, 42(2):9:1– 9:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yu:2017:ODM

[961] Yanwei Yu, Lei Cao, Elke A. Rundensteiner, and Qin Wang. Outlier detection over massive-scale trajectory streams. ACM Transactions on Database Systems, 42(2):10:1–10:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Han:2017:CCF

[962] Yunheng Han, Weiwei Sun, and Baihua Zheng. COMPRESS: a comprehensive framework of trajectory compression in road networks. ACM Transactions on Database Systems, 42 (2):11:1–11:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

TenCate:2017:AAS

[963] Balder Ten Cate, Phokion G. Kolaitis, Kun Qian, and Wang-Chiew Tan. Approximation algorithms for schemamapping discovery from data examples. ACM Transactions on Database Systems, 42(2):12:1–12:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Maniu:2017:IFQ

[964] Silviu Maniu, Reynold Cheng, and Pierre Senellart. An indexing framework for queries on probabilistic graphs. ACM Transactions on Database Systems, 42(2):13:1–13:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Gan:2017:HAE

[965] Junhao Gan and Yufei Tao. On the hardness and approximation of Euclidean DBSCAN. ACM Transactions on Database Systems, 42(3): 14:1–14:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Martens:2017:BCS

[966] Wim Martens, Frank Neven, Matthias Niewerth, and Thomas Schwentick. BonXai: Combining the simplicity of DTD with the expressiveness of XML schema. ACM Transactions on Database Systems, 42(3):15:1–15:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zheng:2017:ESB

[967] Weiguo Zheng, Lei Zou, Lei Chen, and Dongyan Zhao. Efficient SimRankbased similarity join. ACM Transactions on Database Systems, 42(3): 16:1–16:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kaminski:2017:QNA

[968] Mark Kaminski, Egor V. Kostylev, and Bernardo Cuenca Grau. Query nesting, assignment, and aggregation in SPARQL 1.1. ACM Transactions on Database Systems, 42(3):17:1–17:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Tschirschnitz:2017:DID

[969] Fabian Tschirschnitz, Thorsten Papenbrock, and Felix Naumann. Detecting inclusion dependencies on very many tables. ACM Transactions on Database Systems, 42(3):18:1–18:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schubert:2017:DRR

[970] Erich Schubert, Jörg Sander, Martin Ester, Hans Peter Kriegel, and Xiaowei Xu. DBSCAN revisited, revisited: Why and how you should (still) use DBSCAN. ACM Transactions on Database Systems, 42(3): 19:1–19:??, August 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Aberger:2017:ERE

[971] Christopher R. Aberger, Andrew Lamb, Susan Tu, Andres Nötzli, Kunle Olukotun, and Christopher Ré. EmptyHeaded: a relational engine for graph processing. ACM Transactions on Database Systems, 42(4):20:1–20:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2017:DQL

[972] Marcelo Arenas and Martin Ugarte. Designing a query language for RDF: Marrying open and closed worlds. ACM Transactions on Database Systems, 42(4):21:1–21:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Barany:2017:DPP

[973] Vince Bárány, Balder Ten Cate, Benny Kimelfeld, Dan Olteanu, and Zografoula Vagena. Declarative probabilistic programming with Datalog. ACM Transactions on Database Systems, 42(4):22:1–22:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Alvaro:2017:BCA

[974] Peter Alvaro, Neil Conway, Joseph M. Hellerstein, and David Maier. Blazes: Coordination analysis and placement for distributed programs. ACM Transactions on Database Systems, 42(4): 23:1–23:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Colazzo:2017:LTM

[975] Dario Colazzo, Giorgio Ghelli, and Carlo Sartiani. Linear time membership in a class of regular expressions with counting, interleaving, and unordered concatenation. ACM Transactions on Database Systems, 42(4): 24:1–24:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Zhang:2017:PPD

[976] Jun Zhang, Graham Cormode, Cecilia M. Procopiuc, Divesh Srivastava, and Xiaokui Xiao. PrivBayes: Private data release via Bayesian networks. ACM Transactions on Database Systems, 42(4):25:1–25:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Jensen:2018:EUE

[977] Christian S. Jensen. Editorial: Updates to the Editorial Board. ACM Transactions on Database Systems, 43 (1):1:1–1:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Demertzis:2018:PPR

[978] Ioannis Demertzis, Stavros Papadopoulos, Odysseas Papapetrou, Antonios Deligiannakis, Minos Garofalakis, and Charalampos Papamanthou. Practical private range search in depth. ACM Transactions on Database Systems, 43 (1):2:1–2:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chung:2018:EIU

[979] Yeounoh Chung, Michael Lind Mortensen, Carsten Binnig, and Tim Kraska. Estimating the impact of unknown unknowns on aggregate query results. ACM Transactions on Database Systems, 43(1):3:1–3:??, April 2018. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Shaikhha:2018:BEQ

[980] Amir Shaikhha, Yannis Klonatos, and Christoph Koch. Building efficient query engines in a high-level language. ACM Transactions on Database Systems, 43(1):4:1–4:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Libkin:2018:TNA

[981] Leonid Libkin, Juan L. Reutter, Adrián Soto, and Domagoj Vrgoc. TriAL: a navigational algebra for RDF triplestores. ACM Transactions on Database Systems, 43(1):5:1–5:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Cao:2018:BQR

[982] Yang Cao, Wenfei Fan, Floris Geerts, and Ping Lu. Bounded query rewriting using views. ACM Transactions on Database Systems, 43(1):6:1–6:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Berkholz:2018:AFQ

[983] Christoph Berkholz, Jens Keppeler, and Nicole Schweikardt. Answering FO+MOD queries under updates on bounded degree databases. ACM Transactions on Database Systems, 43 (2):7:1–7:??, September 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Barcelo:2018:EES

[984] Pablo Barceló, Markus Kröll, Reinhard Pichler, and Sebastian Skritek. Efficient evaluation and static analysis for well-designed pattern trees with projection. ACM Transactions on Database Systems, 43(2):8:1–8:??, September 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lazerson:2018:LMD

[985] Arnon Lazerson, Daniel Keren, and Assaf Schuster. Lightweight monitoring of distributed streams. ACM Transactions on Database Systems, 43(2): 9:1–9:??, September 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Qi:2018:RQU

[986] Jianzhong Qi, Fei Zuo, Hanan Samet, and Jia Cheng Yao. K-regret queries using multiplicative utility functions. ACM Transactions on Database Systems, 43(2):10:1–10:??, September 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kimelfeld:2018:RFC

[987] Benny Kimelfeld and Christopher Ré. A relational framework for classifier engineering. ACM Transactions on Database Systems, 43(3):11:1–11:??, November 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Schwentick:2018:DCU

[988] Thomas Schwentick, Nils Vortmeier, and Thomas Zeume. Dynamic complexity under definable changes. ACM Transactions on Database Systems, 43 (3):12:1–12:??, November 2018. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Arenas:2018:ELQ

[989] Marcelo Arenas, Georg Gottlob, and Andreas Pieris. Expressive languages for querying the Semantic Web. ACM Transactions on Database Systems, 43 (3):13:1–13:??, November 2018. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Polychroniou:2018:DJD

[990] Orestis Polychroniou, Wangda Zhang, and Kenneth A. Ross. Distributed joins and data placement for minimal network traffic. ACM Transactions on Database Systems, 43(3): 14:1–14:??, November 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Abuzaid:2018:MPA

[991] Firas Abuzaid, Peter Bailis, Jialin Ding, Edward Gan, Samuel Madden, Deepak Narayanan, Kexin Rong, and Sahaana Suri. MacroBase: Prioritizing attention in fast data. ACM Transactions on Database Systems, 43(4): 15:1–15:??, December 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Dayan:2018:OBF

[992] Niv Dayan, Manos Athanassoulis, and Stratos Idreos. Optimal Bloom filters and adaptive merging for LSM-trees. ACM Transactions on Database Systems, 43(4):16:1–16:??, December 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Meneghetti:2018:LQA

[993] Niccolò Meneghetti, Oliver Kennedy, and Wolfgang Gatterbauer. Learning from query-answers: a scalable approach to belief updating and parameter learning. ACM Transactions on Database Systems, 43(4):17:1–17:??, December 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2018:PSG

[994] Wenfei Fan, Wenyuan Yu, Jingbo Xu, Jingren Zhou, Xiaojian Luo, Qiang Yin, Ping Lu, Yang Cao, and Ruiqi Xu. Parallelizing sequential graph computations. ACM Transactions on Database Systems, 43(4):18:1–18:??, December 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Kipf:2019:SAF

[995] Andreas Kipf, Varun Pandey, Jan Böttcher, Lucas Braun, Thomas Neumann, and Alfons Kemper. Scalable analytics on fast data. ACM Transactions on Database Systems, 44 (1):1:1–1:??, January 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2019:WJX

[996] Feifei Li, Bin Wu, Ke Yi, and Zhuoyue Zhao. Wander join and XDB: Online aggregation via random walks. ACM Transactions on Database Systems, 44 (1):2:1–2:??, January 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bai:2019:HMD

[997] Ran Bai, Wing Kai Hon, Eric Lo, Zhian He, and Kenny Zhu. Historic moments discovery in sequence data. ACM Transactions on Database Systems, 44 (1):3:1–3:??, January 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Alexandrov:2019:ROE

[998] Alexander Alexandrov, Georgi Krastev, and Volker Markl. Representations and optimizations for embedded parallel dataflow languages. ACM Transactions on Database Systems, 44(1):4:1–4:??, January 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fan:2019:DG

[999] Wenfei Fan and Ping Lu. Dependencies for graphs. ACM Transactions on Database Systems, 44(2):5:1–5:??, April 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3287285.

Hu:2019:OOM

[1000] Xiao Hu, Ke Yi, and Yufei Tao. Output-optimal massively parallel algorithms for similarity joins. ACM Transactions on Database Systems, 44 (2):6:1–6:??, April 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.cfm?id=
                                        3311967.

Dautrich:2019:IIT

[1001] Jonathan L. Dautrich, Jr. and Chinya V. Ravishankar. Inferring insertion times and optimizing error penalties in time-decaying Bloom filters. ACM Transactions on Database Systems, 44(2):7:1–7:??, April 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3284552.

Gummidi:2019:SSC

[1002] Srinivasa Raghavendra Bhuvan Gummidi, Xike Xie, and Torben Bach Pedersen. A survey of spatial crowdsourcing. ACM Transactions on Database Systems, 44(2):8:1–8:??, April 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3291933.

Damme:2019:CES

[1003] Patrick Damme, Annett Ungethüm, Juliana Hildebrandt, Dirk Habich, and Wolfgang Lehner. From a comprehensive experimental survey to a cost-based selection strategy for lightweight integer compression algorithms. ACM Transactions on Database Systems, 44(3):9:1–9:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
ft_gateway.cfm?id=3323991.

Bonifati:2019:IMS

[1004] Angela Bonifati, Ugo Comignani, Emmanuel Coquery, and Romuald Thion. Interactive mapping specification with exemplar tuples. ACM Transactions on Database Systems, 44(3):10:1–10:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
ft_gateway.cfm?id=3321485.

Beedkar:2019:UFF

[1005] Kaustubh Beedkar, Rainer Gemulla, and Wim Martens. A unified framework for frequent sequence mining with subsequence constraints. ACM Transactions on Database Systems, 44 (3):11:1–11:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=
3321486.

Deutsch:2019:VHA

[1006] Alin Deutsch, Yuliang Li, and Victor Vianu. Verification of hierarchical artifact systems. ACM Transactions on Database Systems, 44(3):12:1–12:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3321487.

Seidemann:2019:CHP

[1007] Marc Seidemann, Nikolaus Glombiewski, Michael Körber, and Bernhard Seeger. ChronicleDB: a highperformance event store. ACM Transactions on Database Systems, 44(4): 13:1–13:??, October 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.cfm?id=
                                        3342357.

Brijder:2019:EPQ

[1008] Robert Brijder, Floris Geerts, Jan Van Den Bussche, and Timmy Weerwag. On the expressive power of query languages for matrices. ACM Transactions on Database Systems, 44(4):15:1–15:??, October 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3331445.

Martens:2019:DES

[1009] Wim Martens and Tina Trautner. Dichotomies for evaluating simple regular path queries. ACM Transactions on Database Systems, 44(4):16:1–16:??, October 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3331446.

Wang:2019:EAA

[1010] Sibo Wang, Renchi Yang, Runhui Wang, Xiaokui Xiao, Zhewei Wei, Wenqing Lin, Yin Yang, and Nan Tang. Efficient algorithms for approximate single-source personalized PageRank queries. ACM Transactions on Database Systems, 44(4):18:1–18:??, October 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
ft_gateway.cfm?id=3360902.

Jensen:2020:EUE

[1011] Christian S. Jensen. Editorial: Updates to the Editorial Board. ACM Transactions on Database Systems, 45(1):1e:1, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3381020.

McCamish:2020:GTA

[1012] Ben McCamish, Vahid Ghadakchi, Arash Termehchy, Behrouz Touri, Eduardo Cotilla-Sanchez, Liang Huang, and Soravit Changpinyo. A gametheoretic approach to data interaction. ACM Transactions on Database Systems, 45(1):1:1–1:44, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3351450.

Zhang:2020:FDD

[1013] Dan Zhang, Ryan McKenna, Ios Kotsogiannis, George Bissias, Michael Hay, Ashwin Machanavajjhala, and Gerome Miklau. ϵKTELO: a framework for defining differentially private computations. ACM Transactions on Database Systems, 45(1):2:1–2:44, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3362032.

Florenzano:2020:EEA

[1014] Fernando Florenzano, Cristian Riveros, Martín Ugarte, Stijn Vansummeren, and Domagoj Vrgoc. Efficient enumeration algorithms for regular document spanners. ACM Transactions on Database Systems, 45(1):3:1–3:42, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3351451.

Livshits:2020:COR

[1015] Ester Livshits, Benny Kimelfeld, and Sudeepa Roy. Computing optimal repairs for functional dependencies. ACM Transactions on Database Systems, 45(1):4:1–4:46, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3360904.

Zhang:2020:SRF

[1016] Huanchen Zhang, Hyeontaek Lim, Viktor Leis, David G. Andersen, Michael Kaminsky, Kimberly Keeton, and Andrew Pavlo. Succinct range filters. ACM Transactions on Database Systems, 45(2):5:1–5:31, July 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3375660.

Fan:2020:AAP

[1017] Wenfei Fan, Ping Lu, Wenyuan Yu, Jingbo Xu, Qiang Yin, Xiaojian Luo, Jingren Zhou, and Ruochun Jin. Adaptive asynchronous parallelization of graph algorithms. ACM Transactions on Database Systems, 45(2):6:1–6:45, July 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3397491.

Khamis:2020:LMR

[1018] Mahmoud Abo Khamis, Hung Q.

Ngo, Xuanlong Nguyen, Dan Olteanu, and Maximilian Schleich. Learning models over relational data using sparse tensors and functional dependencies. ACM Transactions on Database Systems, 45(2):7:1–7:66, July 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3375661.

Kolaitis:2020:LNT

[1019] Phokion G. Kolaitis, Reinhard Pichler, Emanuel Sallinger, and Vadim Savenkov. On the language of nested tuple generating dependencies. ACM Transactions on Database Systems, 45(2):8:1–8:59, July 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3369554.

Fan:2020:CNI

[1020] Wenfei Fan, Xueli Liu, Ping Lu, and Chao Tian. Catching numeric inconsistencies in graphs. ACM Transactions on Database Systems, 45(2):9:1–9:47, July 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3385031.

Jermaine:2020:E

[1021] Chris Jermaine. Editorial. ACM Transactions on Database Systems, 45 (3):10:1, September 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3417730.

Kara:2020:MTQ

[1022] Ahmet Kara, Hung Q. Ngo, Milos Nikolic, Dan Olteanu, and Haozhe Zhang. Maintaining triangle queries under updates. ACM Transactions on Database Systems, 45(3):11:1–11:46, September 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3396375.

Shaikhha:2020:SIL

[1023] Amir Shaikhha, Mohammed Elseidy, Stephan Mihaila, Daniel Espino, and Christoph Koch. Synthesis of incremental linear algebra programs. ACM Transactions on Database Systems, 45 (3):12:1–12:44, September 2020. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3385398.

Schirmer:2020:EDM

[1024] Philipp Schirmer, Thorsten Papenbrock, Ioannis Koumarelas, and Felix Naumann. Efficient discovery of matching dependencies. ACM Transactions on Database Systems, 45(3): 13:1–13:33, September 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3392778.

Qi:2020:PTS

[1025] Jianzhong Qi, Yufei Tao, Yanchuan Chang, and Rui Zhang. Packing R-trees with space-filling curves: Theoretical optimality, empirical efficiency, and bulk-loading parallelizability. ACM Transactions on Database Systems, 45(3):14:1–14:47, September 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3397506.

Fan:2020:DGF

[1026] Wenfei Fan, Chunming Hu, Xueli Liu, and Ping Lu. Discovering graph functional dependencies. ACM Transactions on Database Systems, 45(3): 15:1–15:42, September 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3397198.

Nakandala:2020:IAC

[1027] Supun Nakandala, Kabir Nagrecha, Arun Kumar, and Yannis Papakonstantinou. Incremental and approximate computations for accelerating deep CNN inference. ACM Transactions on Database Systems, 45(4): 16:1–16:42, December 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3397461.

Khamis:2020:FAQ

[1028] Mahmoud Abo Khamis, Ryan R.

Curtin, Benjamin Moseley, Hung Q. Ngo, Xuanlong Nguyen, Dan Olteanu, and Maximilian Schleich. Functional aggregate queries with additive inequalities. ACM Transactions on Database Systems, 45(4):17:1–17:41, December 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

org/doi/10.1145/3426865.

Ciaccia:2020:FSD

[1029] Paolo Ciaccia and Davide Martinenghi. Flexible skylines: Dominance for arbitrary sets of monotone functions. ACM Transactions on Database Systems, 45(4):18:1–18:45, December 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3406113.

Zimanyi:2020:MMD

[1030] Esteban Zimányi, Mahmoud Sakr, and Arthur Lesuisse. MobilityDB: a mobility database based on PostgreSQL and PostGIS. ACM Transactions on Database Systems, 45(4):19:1–19:42, December 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3406534.

Traub:2021:SGE

[1031] Jonas Traub, Philipp Marian Grulich, Alejandro Rodríguez Cuéllar, Sebastian Breß, Asterios Katsifodimos, Tilmann Rabl, and Volker Markl. Scotty: General and efficient opensource window aggregation for stream processing systems. ACM Transactions on Database Systems, 46(1):1:1–1:46, April 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3433675.

Amarilli:2021:CDE

[1032] Antoine Amarilli, Pierre Bourhis, Stefan Mengel, and Matthias Niewerth. Constant-delay enumeration for nondeterministic document spanners. ACM Transactions on Database Systems, 46 (1):2:1–2:30, April 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3436487.

Mitchell:2021:ESM

[1033] Rory Mitchell, Eibe Frank, and Geoffrey Holmes. An empirical study of moment estimators for quantile approximation. ACM Transactions on Database Systems, 46(1):3:1–3:21, April 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442337.

Meduri:2021:EML

[1034] Venkata Vamsikrishna Meduri, Kanchan Chowdhury, and Mohamed Sarwat. Evaluation of machine learning algorithms in predicting the next SQL query from the future. ACM Transactions on Database Systems, 46 (1):4:1–4:46, April 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3442338.

Carmeli:2021:ECU

[1035] Nofar Carmeli and Markus Kröll. On the enumeration complexity of unions of conjunctive queries. ACM Transactions on Database Systems, 46(2):5:1– 5:41, June 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

org/doi/10.1145/3450263.

Mhedhbi:2021:OOT

[1036] Amine Mhedhbi, Chathura Kankanamge, and Semih Salihoglu. Optimizing onetime and continuous subgraph queries using worst-case optimal joins. ACM Transactions on Database Systems, 46 (2):6:1–6:45, June 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3446980.

Wei:2021:EFD

[1037] Ziheng Wei and Sebastian Link. Embedded functional dependencies and data-completeness tailored database design. ACM Transactions on Database Systems, 46(2):7:1–7:46, June 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3450518.

Chen:2021:GIE

[1038] Yangjun Chen and Gagandeep Singh. Graph indexing for efficient evaluation of label-constrained reachability queries. ACM Transactions on Database Systems, 46(2):8:1–8:50, June 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3451159.

Trummer:2021:SRB

[1039] Immanuel Trummer, Junxiong Wang, Ziyun Wei, Deepak Maram, Samuel Moseley, Saehan Jo, Joseph Antonakakis, and Ankush Rayabhari. SkinnerDB: Regret-bounded query evaluation via reinforcement learning. ACM Transactions on Database Systems, 46 (3):9:1–9:45, September 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3464389.

Song:2021:SDC

[1040] Shaoxu Song, Fei Gao, Aoqian Zhang, Jianmin Wang, and Philip S. Yu.

Stream data cleaning under speed and acceleration constraints. ACM Transactions on Database Systems, 46(3): 10:1–10:44, September 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3465740.

Lin:2021:EBL

[1041] Xuelian Lin, Shuai Ma, Jiahao Jiang, Yanchen Hou, and Tianyu Wo. Error bounded line simplification algorithms for trajectory compression: an experimental evaluation. ACM Transactions on Database Systems, 46(3): 11:1–11:44, September 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3474373.

Khamis:2021:BQC

[1042] Mahmoud Abo Khamis, Phokion G.

Kolaitis, Hung Q. Ngo, and Dan Suciu. Bag query containment and information theory. ACM Transactions on Database Systems, 46(3):12:1– 12:39, September 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3472391.

Ma:2021:DDS

[1043] Chenhao Ma, Yixiang Fang, Reynold Cheng, Laks V. S. Lakshmanan, Wenjie Zhang, and Xuemin Lin. On directed densest subgraph discovery. ACM Transactions on Database Systems, 46 (4):13:1–13:45, December 2021. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

https://dl.acm.org/doi/10.1145/
3483940.

Singh:2021:TRH

[1044] Shikha Singh, Prashant Pandey, Michael A. Bender, Jonathan W. Berry, Martín Farach-Colton, Rob Johnson, Thomas M. Kroeger, and Cynthia A. Phillips. Timely reporting of heavy hitters using external memory. ACM Transactions on Database Systems, 46(4):14:1–14:35, December 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3472392.

Benedikt:2021:BEI

[1045] Michael Benedikt, Pierre Bourhis, Louis Jachiet, and Efthymia Tsamoura. Balancing expressiveness and inexpressiveness in view design. ACM Transactions on Database Systems, 46(4): 15:1–15:40, December 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3488370.

Grez:2021:FFC

[1046] Alejandro Grez, Cristian Riveros, Martín Ugarte, and Stijn Vansummeren. A formal framework for complex event recognition. ACM Transactions on Database Systems, 46(4): 16:1–16:49, December 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3485463.

Berger:2022:SEC

[1047] Gerald Berger, Georg Gottlob, Andreas Pieris, and Emanuel Sallinger. The space-efficient core of Vadalog. ACM Transactions on Database Systems, 47 (1):1:1–1:46, March 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3488720.

Tong:2022:URP

[1048] Yongxin Tong, Yuxiang Zeng, Zimu Zhou, Lei Chen, and Ke Xu. Unified route planning for shared mobility: an insertion-based framework. ACM Transactions on Database Systems, 47 (1):2:1–2:48, March 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3488723.

Binna:2022:HOT

[1049] Robert Binna, Eva Zangerle, Martin Pichl, Günther Specht, and Viktor Leis. Height optimized tries. ACM Transactions on Database Systems, 47 (1):3:1–3:46, March 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3506692.

Aumuller:2022:SNN

[1050] Martin Aumüller, Sariel Har-Peled, Sepideh Mahabadi, Rasmus Pagh, and Francesco Silvestri. Sampling a near neighbor in high dimensions — who is the fairest of them all? ACM Transactions on Database Systems, 47 (1):4:1–4:40, March 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3502867.

Schmid:2022:CRP

[1051] Markus L. Schmid. Conjunctive regular path queries with capture groups. ACM Transactions on Database Systems, 47 (2):5:1–5:52, June 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3514230.

Fan:2022:IGC

[1052] Wenfei Fan and Chao Tian. Incremental graph computations: Doable and undoable. ACM Transactions on Database Systems, 47(2):6:1–6:44, June 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3500930.

Cheng:2022:MOP

[1053] Ji Cheng, Da Yan, Wenwen Qu, Xiaotian Hao, Cheng Long, Wilfred Ng, and Xiaoling Wang. Mining order-preserving submatrices under data uncertainty: a possibleworld approach and efficient approximation methods. ACM Transactions on Database Systems, 47(2):7:1–7:57, June 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524915.

Arroyuelo:2022:OJU

[1054] Diego Arroyuelo, Gonzalo Navarro, Juan L. Reutter, and Javiel Rojas-Ledesma. Optimal joins using compressed quadtrees. ACM Transactions on Database Systems, 47(2):8:1–8:53, June 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3514231.

Carmeli:2022:AUC

[1055] Nofar Carmeli, Shai Zeevi, Christoph Berkholz, Alessio Conte, Benny Kimelfeld, and Nicole Schweikardt. Answering (unions of) conjunctive queries using random access and randomorder enumeration. ACM Transactions on Database Systems, 47(3):9:1–9:49, September 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

org/doi/10.1145/3531055.

Asudeh:2022:FRR

[1056] Abolfazl Asudeh, Gautam Das, H. V.

Jagadish, Shangqi Lu, Azade Nazi, Yufei Tao, Nan Zhang, and Jianwen Zhao. On finding rank regret representatives. ACM Transactions on Database Systems, 47(3): 10:1–10:37, September 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3531054.

Zeng:2022:PS

[1057] Tianjing Zeng, Zhewei Wei, Ge Luo, Ke Yi, Xiaoyong Du, and Ji-Rong Wen. Persistent summaries. ACM Transactions on Database Systems, 47(3): 11:1–11:42, September 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3531053.

Guo:2022:IMR

[1058] Qintian Guo, Sibo Wang, Zhewei Wei, Wenqing Lin, and Jing Tang. Influence maximization revisited: Efficient sampling with bound tightened. ACM Transactions on Database Systems, 47(3):12:1–12:45, September 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3533817.

Ketsman:2022:DRL

[1059] Bas Ketsman, Christoph Koch, Frank Neven, and Brecht Vandevoort. Deciding robustness for lower SQL isolation levels. ACM Transactions on Database Systems, 47(4):13:1–13:??, December 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3561049.

TenCate:2022:CQU

[1060] Balder Ten Cate and Victor Dalmau. Conjunctive queries: Unique characterizations and exact learnability. ACM Transactions on Database Systems, 47(4):14:1–14:??, December 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3559756.

Wei:2022:PQT

[1061] Victor Junqiu Wei, Raymond Chi-Wing Wong, Cheng Long, David Mount, and Hanan Samet. Proximity queries on terrain surface. ACM Transactions on Database Systems, 47(4): 15:1–15:??, December 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3563773.

Do:2022:ESD

[1062] Thanh Do, Goetz Graefe, and Jeffrey Naughton. Efficient sorting, duplicate removal, grouping, and aggregation. ACM Transactions on Database Systems, 47(4):16:1–16:??, December 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3568027.

Carmeli:2023:TOD

[1063] Nofar Carmeli, Nikolaos Tziavelis, Wolfgang Gatterbauer, Benny Kimelfeld, and Mirek Riedewald. Tractable orders for direct access to ranked answers of conjunctive queries. ACM Transactions on Database Systems, 48 (1):1:1–1:??, March 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://

dl.acm.org/doi/10.1145/3578517.

Do:2023:RES

[1064] Thanh Do and Goetz Graefe. Robust and efficient sorting with offsetvalue coding. ACM Transactions on Database Systems, 48(1):2:1–2:??, March 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3570956.

Huang:2023:ECS

[1065] Ruihong Huang, Jianmin Wang, Shaoxu Song, Xuemin Lin, Xiaochen Zhu, and Jian Pei. Efficiently cleaning structured event logs: a graph repair approach. ACM Transactions on Database Systems, 48(1):3:1–3:??, March 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3571281.

Fakas:2023:PSD

[1066] Georgios J. Fakas and Georgios Kalamatianos. Proportionality on spatial data with context. ACM Transactions on Database Systems, 48(2):4:1–4:??, June 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3588434.

Hu:2023:RDW

[1067] Donghui Hu, Qing Wang, Song Yan, Xiaojun Liu, Meng Li, and Shuli Zheng. Reversible database watermarking based on order-preserving encryption for data sharing. ACM Transactions on Database Systems, 48 (2):5:1–5:??, June 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3589761.

Chen:2023:EBO

[1068] Yaxing Chen, Qinghua Zheng, and Zheng Yan. Efficient bi-objective SQL optimization for enclaved cloud databases with differentially private padding. ACM Transactions on Database Systems, 48(2):6:1–6:??, June 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3597021.

Pavan:2023:MCM

[1069] A. Pavan, N. V. Vinodchandran, Arnab Bhattacharyya, and Kuldeep S. Meel. Model counting meets F_{0} estimation. ACM Transactions on Database Systems, 48(3):7:1–7:??, September 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3603496.

Sarkar:2023:ETP

[1070] Subhadeep Sarkar, Tarikul Islam Papon, Dimitris Staratzis, Zichen Zhu, and Manos Athanassoulis. Enabling timely and persistent deletion in LSM-engines. ACM Transactions on Database Systems, 48(3):8:1–8:??, September 2023. CODEN ATDSD3.

ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

org/doi/10.1145/3599724.

Leventidis:2023:DHD

[1071] Aristotelis Leventidis, Laura Di Rocco, Wolfgang Gatterbauer, Renée J. Miller, and Mirek Riedewald. DomainNet: Homograph detection and understanding in data lake disambiguation. ACM Transactions on Database Systems, 48 (3):9:1–9:??, September 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3612919.

Shangqi:2023:POM

[1072] Lu Shangqi, Wim Martens, Matthias Niewerth, and Yufei Tao. Partial order multiway search. ACM Transactions on Database Systems, 48(4): 10:1–10:??, December 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3626956.

Herodotou:2023:CBD

[1073] Herodotos Herodotou and Elena Kakoulli. Cost-based data prefetching and scheduling in big data platforms over tiered storage systems. ACM Transactions on Database Systems, 48(4): 11:1–11:??, December 2023. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3625389.

Gottlob:2024:FPH

[1074] Georg Gottlob, Matthias Lanzinger, Cem Okulmus, and Reinhard Pichler. Fast parallel hypertree decompositions in logarithmic recursion depth. ACM Transactions on Database Systems, 49 (1):1:1–1:??, March 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3638758.

Fan:2024:LEA

[1075] Wenfei Fan, Ping Lu, Kehan Pang, Ruochun Jin, and Wenyuan Yu. Linking entities across relations and graphs. ACM Transactions on Database Systems, 49(1):2:1–2:??, March 2024. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

https://dl.acm.org/doi/10.1145/
3639363.

Wang:2024:AHT

[1076] Zhaoguo Wang, Chuzhe Tang, Xiaodong Zhang, Qianmian Yu, Binyu Zang, Haibing Guan, and Haibo Chen. Ad hoc transactions through the looking glass: an empirical study of application-level transactions in Web applications. ACM Transactions on Database Systems, 49(1):3:1–3:??, March 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638553.

Currim:2024:IRC

[1077] Sabah Currim, Richard T. Snodgrass, and Young-Kyoon Suh. Identifying the root causes of DBMS suboptimality. ACM Transactions on Database Systems, 49(1):4:1–4:??, March 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3636425.

Arroyuelo:2024:RWC

[1078] Diego Arroyuelo, Adrián Gómez-Brandón, Aidan Hogan, Gonzalo Navarro, Juan Reutter, Javiel Rojas-Ledesma, and Adrián Soto. The ring: Worst-case optimal joins in graph databases using (almost) no extra space. ACM Transactions on Database Systems, 49(2):5:1–5:??, June 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/doi/10.
1145/3644824.

Chapman:2024:SBI

[1079] Adriane Chapman, Luca Lauro, Paolo Missier, and Riccardo Torlone. Supporting better insights of data science pipelines with fine-grained provenance. ACM Transactions on Database Systems, 49(2):6:1–6:??, June 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

https://dl.acm.org/doi/10.1145/
3644385.

Zhang:2024:SQN

[1080] Chao Zhang and Toumani Farouk.

Sharing queries with nonequivalent user-defined aggregate functions. ACM Transactions on Database Systems, 49 (2):7:1–7:??, June 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649133.

Carmeli:2024:DRS

[1081] Nofar Carmeli, Martin Grohe, Benny Kimelfeld, Ester Livshits, and Muhammad Tibi. Database repairing with soft functional dependencies. ACM Transactions on Database Systems, 49 (2):8:1–8:??, June 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3651156.

Tench:2024:GHF

[1082] David Tench, Evan West, Victor Zhang, Michael A. Bender, Abiyaz Chowdhury, Daniel Delayo, J. Ahmed Dellas, Martín Farach-Colton, Tyler Seip, and Kenny Zhang. GraphZeppelin: How to find connected components (even when graphs are dense, dynamic, and massive). ACM Transactions on Database Systems, 49(3): 9:1–9:??, September 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3643846.

Salas:2024:FCM

[1083] Jorge Salas, Cristian Riveros, and Sebastián Bugedo. A family of centrality measures for graph data based on subgraphs. ACM Transactions on Database Systems, 49(3):10:1–10:??, September 2024. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3649134.

Gershtein:2024:ACT

[1084] Shay Gershtein, Uri Avron, Ido Guy, Tova Milo, and Slava Novgorodov. Automated category tree construction: Hardness bounds and algorithms. ACM Transactions on Database Systems, 49 (3):11:1–11:??, September 2024. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3664283.

Hu:2024:ASB

[1085] Pan Hu and Boris Motik. Accurate sampling-based cardinality estimation for complex graph queries. ACM Transactions on Database Systems, 49 (3):12:1–12:??, September 2024. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

https://dl.acm.org/doi/10.1145/
3689209.

Litwin:1980:LHN

[1086] W. Litwin. Linear hashing: a new tool for file and table addressing. In Lochovsky and Taylor [1089], pages 212– 223. Reprinted in [1093, p. 570–581].

Regnier:1985:AGF

[1087] Mireille Regnier. Analysis of grid file algorithms. BIT, 25(2):335– 357, June 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=25&issue=2&spage=335.

Kerr:1975:PIC

[1088] Douglas S. Kerr, editor. Proceedings of the International Conference on Very Large Data Bases, Framingham, MA, USA, September 22–24, 1975. ACM Press, New York, NY 10036, USA, 1975. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. ACM SIGMOD v. 1, no. 1, September 1975.

Lochovsky:1980:SIC

[1089] Frederick H. Lochovsky and ?. Taylor, editors. Sixth International Conference on Very Large Data Bases: reprinted from Very large data bases. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Schkolnick:1983:ICV

[1090] Mario Schkolnick and C. (Costantino) Thanos, editors. 9th International Conference on Very Large Data Bases (Florence, Italy, October 31–November 2, 1983). VLDB Endowment, P.O. Box 2245, Saratoga, CA, USA, 1983. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

ACM:1985:PFA

[1091] ACM, editor. Proceedings of the Fourth ACM SIGACT-SIGMOD Symposium on Principles of Database Systems, March 25–27, 1985, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1985. ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Kambayashi:1986:TIC

[1092] Yahiko Kambayashi, Wesley Chu, Georges Gardarin, and Setsuo Ohsuga, editors. Twelfth International Conference on Very Large Data Bases, Proceedings (VLDB ’86). Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1986. ISBN 0-934613-18-4. xiii + 512 pp. LCCN QA 76.9 D3 I61 1986.

Stonebraker:1988:RDS

[1093] Michael Stonebraker, editor. Readings in Database Systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1988. ISBN 0-934613-65-6. xii + 644 pp. LCCN QA76.9.D3 R4 1988. US$29.95. Contains reprint of [1086].

Garcia-Molina:1990:PAS

[1094] Héctor García-Molina and H. V. Jagadish, editors. Proceedings of the 1990 ACM SIGMOD International Conference on Management of Data, May 23– 25, 1990, Atlantic City, NJ, volume 19(2) of SIGMOD Record. ACM Press, New York, NY 10036, USA, June 1990. ISBN 0-89791-365-5. LCCN QA 76.9 D3 S53 v.19 no.2 1990.

Zdonik:1990:ROO

[1095] B. Zdonik and David Maier, editors.

Readings in Object-Oriented Database Systems. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1990. ISBN 1-55860-000-0. x + 629 pp. LCCN QA76.9.D3 R42 1990. US$29.95.